### v1

# Policymaker meetings as heteroscedasticity shifters:

Abstract
We propose a novel approach to identification in structural vector autoregressions (SVARs) that uses external instruments for heteroscedasticiy of a structural shock of interest. This approach does not require lead/lag exogeneity for identification, does not require heteroskedasticity to be persistent, and facilitates interpretation of the structural shocks. To implement this identification approach in applications, we develop a new method for simultaneous inference of structural impulse responses and other parameters, employing a dependent wild-bootstrap of local projection estimators. This method is robust to an arbitrary number of unit roots and cointegration relationships (including seasonal unit roots), time-varying local means and drifts, and conditional heteroskedasticity of unknown form and can be used with other identification schemes, including Cholesky and the conventional external IV.
We show how to construct pointwise and simultaneous confidence bounds for structural impulse responses and how to compute smoothed local projections with the corresponding confidence bounds. Using simulated data from a standard log-linearized DSGE model, we show that the method can reliably recover the true impulse responses in realistic datasets. As an empirical application, we adopt the proposed method in order to identify monetary policy shock using the dates of Federal Open Market Committee (FOMC) meetings in a standard six-variable VAR. The robustness of our identification and inference methods allows us to construct an instrumental variable for monetary policy shock that dates back to 1965. The resulting impulse response functions (IRFs) for all variables align with the classical Cholesky identification scheme and are different from the narative sign restricted Bayesian VAR estimates. In particular, the response to inflation manifests a price puzzle that is indicative of the cost channel of the interest rates.
Keywords: SVAR, conditional heteroscedasticity, unit roots, cointegration, identification, simultaneous inference, smoothed local projections, narative sign restrictions, minimum distance estimator, monetary policy shocks.
1 Introduction
Unexpected actions by policymakers provide a valuable source of variation that helps to identify how the economy would react to future policy actions (Sims, 1980). Unfortunately, it is often difficult to detect the unexpected component of the actions using the publicly available data. The true plan of future actions and commitments of the policy makers can be kept secret from the public, and the pure unexpected part of the available policy announcements is always a matter of modeler’s assumptions. We propose an identification approach that remains agnostic about the actual size and directions of change in the policy variable (for example, interest rate) due to policymakers’ discretion. Instead, we only assume that the conditional variance of the unexpected change in the policy variable is higher during policymakers’ meetings than at other times. We show that the information contained in the meetings dates is sufficient to uniquely identify both the size and direction of impulse responses to the structural policy shock in a vector autoregression model (VAR).
Our paper applies this novel approach to identification of the monetary policy shocks using the number of the US Federal Reserve Open Market Committee (FOMC) meetings per month as an instrument for heteroskedasticity of a monetary policy shock in a standard six-variable VAR model. Information available on the meetings dates goes back to at least 1965, which gives us instruments with sample sizes that were unprecedented in the monetary SVAR literature. Such long samples come with an additional set of complications. In particular, the inference procedure needs to be robust to the potential presence of time trends and unit roots, which makes the existing popular methods like residual-based bootstrap (Gonçalves and Kilian, 2004) inapplicable. Moreover, the use of the external instruments for the conditional heteroscedasticity requires a novel procedure that would allow joint inference on all the relevant parameters in the model.
We address this challenge by proposing a novel inference procedure for the structural impulse responses (IRF) based on local projection (LP) estimators of Jordà (2005). Specifically, we propose both a formula and a proof of joint asymptotic normality of the local projection estimator in the potential presence of trends, cointegration and any number of unit roots. Next, we extend the proof of consistency of the joint HAC covariance estimator of Andrews (1991) to accomodate nonstationary regressors. Finally, we propose a version of the dependent wild bootstrap procedure of Shao (2010) that can mimic the joint distribution of the local projection estimators across multiple horizons and the estimates of structural parameters. The only regularity conditions that we impose on the non-explosive VAR model is that the one-step ahead forecast errors, jointly with the instrumental variable, follow a stationary vector process with a mixing property and certain higher-order moments necessary for consistency of the HAC estimator. To our knowledge, this provides the first proof of validity of inference on the IRF estimators that allows the general case of unstable but non-explosive VAR models with a general form of stationary conditional heteroscedasticity without pretests. [^2] Our results build upon the long literature on the non-stationary time series including Sims et al. (1990); Tsay and Tiao (1990); De Jong and Davidson (2000); Inoue and Kilian (2002); Montiel Olea and Plagborg-Møller (2021).
The proposed procedure consists of three main steps. First, we estimate reduced-form impulse response functions (i.e. responses to one-step ahead linear forecast errors $\hat{\eta}_{it}$ for each time series $i$). This step is computed using local projection regressions. Second, we estimate contemporaneous impact of the structural shock of interest on a time series $i$ as the sample covariance between pairwise products of one-step ahead forecast errors $\hat{\eta}_{1t}\hat{\eta}_{it}$ and an instrumental variable correlated with the heteroscedasticity of the structural shock of interest. (The time series with index $1$ plays a special role since it is assumed to have non-zero correlation of squared forecast errors with the instrumental variable.) Third, we compute structural IRF for other horizons as products of the reduced form IRF and vector the contemporaneous structural responses after appropriate normalization. The dependent wild-bootstrap is then applied to the relevant score functions.
The novel simultaneous inference on local projections has multiple uses. First, as emphasized in Brüggemann et al. (2016), it is important to make joint draws of the vector of the IRF coefficients and the structural impact vector to obtain valid confidence limits. Second, the procedure allows for simultaneous confidence bounds that cover the entire IRF across all horizons and, for example, allows for inference on the largest and smallest values of IRF. Third, one can compute smoothed local projection estimators and the corresponding confidence bounds using the minimum distance approach to improve efficiency of the estimates.
We illustrate the proposed method in two applications. First, we show that the identification strategy can recover the true IRF of a monetary policy shock in simulated data based on a medium-sized DSGE model of US economy of Smets and Wouters (2007). Then we apply the method to study the impact of monetary policy shocks in six-variable VAR model using the data set of Antolín-Díaz and Rubio-Ramírez (2018). The variables include Real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and Federal funds rate. The sample period is from January 1965 to November 2007. Our primary instrumental variable for conditional heteroscedasticity of the monetary policy shocks is the number of FOMC meetings per month, including telephone conferences and unscheduled meetings. For reduced form estimates, we use local projection with polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. We report various confidence bands: point-wise, simultaneous sup-t, smoothed point-wise and smoothed simultaneous bands.
Our main empirical findings can be summarized as follows: (i) the identification through the heteroscedasticity IV results in IRF that are identical to the classical Cholesky identification a la Christiano et al. (1999);(ii) smoothed LP-IRF are noticeably different from the conventional recursive VAR estimates at the long horizons; (iii) effects on real GDP are comparable with BVAR identified using the narrataive sign restriction effects (Antolín-Díaz and Rubio-Ramírez, 2018), while the effects on the monetary shocks on inflation are much weaker in the novel scheme.
The literature on identification through hetersocedasticity has a long history (for example, Feenstra (1994); Rigobon (2003); Lewis (2021)). To our knowledge, we are the first to propose an IV-type procedure that assumes that the conditional heteroscedasticity is not-persistent (c.f. Rigobon, 2003; Lewis, 2021), or structural shocks are fully independent and non-Gaussian (c.f. Lütkepohl and Netšunajev, 2017). Moreover, we do not assume any particular parametric model of the innovations and the instruments. In contrast to statistical identification procedures, our approach does not require full identification of all shocks and gives an intuitive interpretation of the identified shock of interest.
In this paper, we apply our identification method to monetary policy shocks, but the external heterscedasticity IV can be proposed in other applications. For example, Känzig (2021) shows that OPEC meetings have high-frequency impacts on oil prices that can be used to identify oil supply shocks; Janzen et al. (2018) show that cotton storage levels change the heteroskedasticity of precautionary demand shocks. So we believe that our approach can be used more broadly.
Our inference procedure is based on the insight that both the reduced-form impulse response and covariance estimators involved in the structural identification can be represented as a mean of vector-valued stationary time series of cross-products of the $1$-step and $h$-step ahead forecast errors. This representation allows us to employ the dependent wild bootstrap of Shao (2010) to make draws from the joint asymptotic distribution of the estimators without explicitly computing the long-run variance-covariance matrix estimator. This novel inference procedure can be used for many other identification approaches that include Cholesky (Christiano et al., 1999), sign restrictions (Gafarov et al., 2016), and external IV (Stock, 2008; Mertens and Ravn, 2013).
The paper is structured as follows. Section 2 discusses the setup and the novel identification idea. Section 3 provides the estimation and inference algorithms with the corresponding theoretical proofs. Section 4 discusses the DSGE experiment. Section 5 has the empirical application to US monetary policy. Section 6 concludes. All the proofs and additional figures are provided in Appendix.
2 Model and the identification
2.1 Motivating example
Consider a simultaneous equations model
$\eta_{t}=B\varepsilon_{t},$ (1)
where observations $\eta_{t}\in\mathbb{R}^{n}$ and $\varepsilon_{t}$ is an i.i.d. zero-mean process with $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\prime}=I_{n}$ that is refereed to as structural shocks. [^3] The simultaneous equations model with lags were first proposed and studied in Haavelmo (1943); Mann and Wald (1943b). Suppose that we are interested in measuring the impact of the structural shock $\varepsilon_{1t}$ on the vector of observables $\eta_{t}$, which is given by the first column of $B=(b_{\cdot 1},\dots,b_{\cdot n})$. We can estimate $\Sigma_{\eta}\triangleq\mathbb{E}\eta_{t}\eta_{t}^{\prime}$ using the sample analog and the solve the equation $\Sigma_{\eta}=BB^{\prime}$ for the first column of $B$. However, there is a continuum of solutions to this equation. Therefore, additional assumptions or data are required to uniquely select $B$. The purpose of the paper is to propose a novel way of identifying $b_{\cdot 1}$ using external instruments for the conditional heteroscedasticity process of $\varepsilon_{t}$ and develop the inference procedure. This identification problem occurs in the more general SVAR models, which we will study later in the paper.
Before we state the novel identification proposal in the general form, let us consider a simple motivating example.
Example 1.
Suppose $n=2$ and that structural shocks have a representation
$\varepsilon_{it}=\xi_{it}\sigma_{it},$ (2)
such that $\xi_{it}$ is i.i.d. sequence with $\mathbb{E}\xi_{t}=0$ and $\mathbb{E}\xi_{t}\xi^{\prime}_{t}=I_{2}$; the process $\sigma_{it}$ is a stationary ergodic process that is independent of $\xi_{it}$. The volatility process $\sigma^{2}_{1t}$ for the shock of interest may be correlated with some observable time series of $Z_{t}$. Suppose that this time series $Z_{t}$ only increases volatility of the first structural shock, but does not change volatility for other shocks. We can then use it as an instrument for identification of $b_{\cdot 1}$. In contrast to the conventional proxy-VAR identification, the instrument $Z_{t}$ may be independent of $\xi_{1t}$. Indeed, suppose that
$\mathbb{E}(Z_{t}-\mathbb{E}Z_{t})(\sigma^{2}_{it}-\mathbb{E}\sigma^{2}_{it})=\rho_{i}$ (3)
such that $\rho_{1}\neq 0$ and $\rho_{2}=0$ and $\xi_{t}$ is independent of $Z_{t}$. Then since
$\eta^{2}_{1t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t}).$ (4)
$\eta_{1t}\eta_{2t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{21}\varepsilon_{1t}+b_{22}\varepsilon_{2t}),$ (5)
we get
$\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}^{2}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}^{2}\rho_{1}$ (6)
$\mathbb{E}\eta_{1t}\eta_{2t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}b_{21}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}b_{21}\rho_{1}.$ (7)
Assuming that $b_{11}$ is not zero, we can take the ratio of these covariances to get
$\frac{\mathbb{E}\eta_{1t}\eta_{2t}(Z_{t}-\mathbb{E}Z_{t})}{\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})}=\frac{b_{21}}{b_{11}}.$ (8)
Once the ratio is calculated, we can recover the full vector $b_{\cdot 1}$ of impacts to a shock of size equal to one standard deviation or using another normalization choice. Thus, we can recover structural impulse responses to the first structural shock $\varepsilon_{1t}$.
The same logic applies if $Z_{t}$ is correlated with $\xi_{t}$ as long as $\rho_{1}\neq 0$, $\rho_{2}=0$ and $\mathbb{E}\xi_{1t}\xi_{2t}(Z_{t}-\mathbb{E}Z_{t})=0$. ∎
The identification conditions in Example 1 can be generalized to
Assumption 1.
Instruments and shocks satisfy
$\gamma\triangleq$ $\mathbb{E}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})\neq 0,$ (9)
$\mathbb{E}\varepsilon_{it}\varepsilon_{jt}(Z_{t}-\mathbb{E}Z_{t})=\rho_{1}\mathds{1}\{i=j=1\}$ (10)
for some $\rho_{1}\neq 0$.
This assumption implies that the heteroscedasticity of the innovations for this time series must be correlated with the instrument $Z_{t}$. In the monetary shock example, one can by definition assume that the impact $b_{11}$ is non-zero for federal funds rate ($i=1$). The instrument can only change the conditional variance of the shock of interest and cannot result in conditional correlation of the other structural shocks.
Theorem 1.
Suppose that Assumption 1 holds and $\Sigma_{\eta}$ is a full-rank matrix. Then the impact vector for the first structural shock $\varepsilon_{1t}$ on $\eta_{t}$ is given by
$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma.$ (11)
Proof.
See Appendix B.1. ∎
A recent work Lewis (2021) showed that in the case when heterscedasticity of a structural shock is persistent, one can use internal instruments $Z_{t}=\eta^{2}_{t-1}$. However, the approach can be applied more generally if appropriate external instruments are available, as in our empirical application.
In most practical applications, data $y_{t}$ have a time dependence and innovations $\eta_{t}$ have to be estimated, which is the topic of the next section.
2.2 Structural impulse responses in the general VAR model.
Suppose that we observe a $n-$vector time series $y_{t}$ that follows a VAR(p) model,
$y_{t}=V\mu_{t}+A_{1}y_{t-1}+\cdots+A_{p}y_{t-p}+\eta_{t},$ (12)
where $y_{t}$ is $n$-dimensional time series, $V\mu_{t}$ is a deterministic polynomial time trend (or time-varying drift for integrated processes) with $\mu_{t}=(1,t,t^{2},\dots,t^{k})$ for some $k$; $A_{1},\dots,A_{p},V$ are $(n\times n)$ parameter matrices, $\eta_{t}$ are serially uncorrelated forecast errors (innovations). We will assume that $p$ is finite and known.
Assumption 2.
The innovations $\eta_{t}$ satisfy
  1. 1.
    $\eta_{t}$ is a martingale difference sequence (MDS), i.e. $\forall$ $t$ $\mathbb{E}[\eta_{t}|\mathcal{F}_{t-1}]=0$
  2. 2.
    $(\eta_{t},Z_{t})$ is a stationary sequence with $\alpha$-mixing coefficients satisfying $\alpha(j)=O(j^{-3(1+\epsilon)/\epsilon})$ for some $\epsilon>0$.
  3. 3.
    $\mathbb{E}||\eta_{t}||^{8+2\epsilon}<\infty$ and $\mathbb{E}|Z|^{4+\epsilon}<\infty$ for some $\epsilon>0$.
  4. 4.
    $\Sigma_{\eta}=\mathbb{E}(\eta_{t}\eta^{\prime}_{t})$ is a full-rank matrix.
  5. 5.
    $\mathbb{E}Z_{t}\eta_{t}=0$ and $\mathbb{E}Z_{t}\eta_{t}\eta^{\prime}_{t-\ell}=0$ for all $\ell\geq 1$.
Assumption 2.1 implies that innovations are white noise, which is required for the consistency of LS estimators in stationary VAR models. The martingale difference requirement with respect to the sigma algebra $\mathcal{F}_{t-1}=\sigma(Y_{t-1},Y_{t-2},\dots)$ is only necessary for interpretation of the impulse response coefficients that will be introduced later in this section. Assumption 2.2 allows for many stationary models of conditional heteroscedasticity. [^4] See, for example, Assumption 2.1.(i-iii) in Brüggemann et al. (2016). This assumption, for example, allows for asymmetric conditional heteroscedasticity models like the ones considered in Rabemananjara and Zakoian (1993). Such models are ruled out by popular residual wild bootsrap methods (Gonçalves and Kilian, 2004). We do not consider unconditional heteroscedasticity in innovations, since it would eliminate the need for structural identification. The 8-th moment condition imposed in our paper is sufficient for consistency of the HAC estimators, but a weaker 4-th moment condition would be sufficient for asymptotic normality of the LP estimators. Assumption 2.4 is standard in papers on structural identification of the shocks. Assumption 2.5 is required for asymptotic normality of $\hat{\gamma}$ and would be satisfied, for example, if $\eta_{t}$ is MDS with respect to a sigma field $\mathcal{F}_{t-1}$ that also includes $Z_{t}$ and past values of $\eta_{t}$ (for example, it is possible that the number of meetings in month $t$ is fully predictable based on the past economic conditions and the past number of meetings). This assumption is not required for consistency of $\hat{\gamma}$ and the structural IRF.
Using the VAR model, one can define impulse responses for individual innovations $\eta_{t}$ since
$C_{h,i,j}=\mathbb{E}(y_{t+h}|\eta_{t}=e_{j},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\eta_{t}=0,\mathcal{F}_{t-1})$ (13)
where $e_{j}\in R^{n}$ is vector of zeros with $1$ in $j-$th position. These coefficients appear as regression coefficients in $H$-step ahead autoregression representation of (12),
$y_{t+H}$ $=V(H)\mu_{t}+\sum_{i=1}^{p}A_{i}{(H)}y_{t-i}+\eta_{t+H}(H),$ (14)
$\eta_{t}(H)$ $\triangleq\sum_{h=0}^{H}C_{h}\eta_{t-h},\text{ for all }H,t\in\mathbb{N},$ (15)
where $C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ (assuming $A_{\ell}=0$ for $\ell>p$), matrices $(V{(H)},A_{1}{(H)},\dots,A_{p}{(H)})$ are functions of $(V,A_{1},\dots,A_{p})$ that can also be computed recursively.
In the absence of deterministic trend and under conditions $det(I_{n}-A_{1}z-\dots-A_{p}z^{p})\neq 0$ for $|z|\leq 1$ and $V=0$, the VAR model can be inverted to obtain the familiar causal $MA(\infty)$ representation (Wold, 1938),
$y_{t}=c+\sum_{h=0}^{\infty}C_{h}\eta_{t-h}.$ (16)
In this case, coefficients $C_{h}$ coincide with the coefficients in the Wold representation. We do not need to rely on this stability assumption throughout the paper. Instead, we allow for an arbitrary number of unit roots and cointegrating vectors.
Assumption 3.
$det(I_{n}-A_{1}z-\dots-A_{p}z^{p})\neq 0$ for $|z|<1$.
Assumption 3 allows for arbitrary number of unit roots, including complex unit roots, and cointegrating relations between $y_{t}$. It only rules out explosive roots. In practical terms, it means that one can apply our inference procedure to local projections estimated with or without trend in levels without seasonal adjustment. It can be done without any pretests for trend stationarity, abnsency of cointegration, or unit roots. We allow both a polynomial trend as in Sims et al. (1990) and complex unit roots as in Tsay and Tiao (1990). Under this assumption, the least squares estimator for the coefficients in equations (12) and (14) is consistent and has well-defined distribution limits as shown in Sims et al. (1990); Tsay and Tiao (1990) under conditionally homosecdastic MDS innovations $\eta_{t}$. In contrast, we allow for more general white noise innovations with mixing by using strong laws of large numbers and functional central limit theorems for strong mixing sequences (De Jong and Davidson, 2000; Rio et al., 2017).
Using the coefficients $C_{0},\dots,C_{H}$ under Assumption 2.1, we can compute structural impulse response function of $y_{t+h}$ to $\varepsilon_{1t}$, which is given by
$\psi_{h}\triangleq\mathbb{E}(y_{t+h}|\varepsilon_{t}=e_{1},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\varepsilon_{t}=0,\mathcal{F}_{t-1})=C_{h}b_{\cdot 1}.$ (17)
The coefficients $C_{1},\dots,C_{H}$ can be estimated using local projection estimators of Jordà (2005). However, our empirical strategy uses conditional heteroskedasticity in the structural shocks in the dataset that has unit root time series with potential cointegration and time trends. Furthermore, structural impulse responses depend on both estimators of $\Sigma_{\eta}$ and $C_{1},\dots,C_{H}$ and on covariances with external instruments $Z_{t}$. Existing inference methods cannot take into account all of these features together. Specifically, there are no proposals for joint inference of $\Sigma_{\eta}$ and $C_{0},\dots,C_{H}$ without pretests for unit roots.  [^5] Brüggemann et al. (2016) provide an example where commonly used residual wild-bootstrap is invalid for joint inference on $C_{i}$ and $\Sigma_{\eta}$ even in the stationary case. Their proposed procedure, residual block bootstrap, solves the issue for stationary VARs. Jentsch and Lunsford (2022) then extends this residual block bootstrap approach to allow for external instrumental variables, also in the stationary case. Inoue and Kilian (2002); Mikusheva (2012) allow for joint inference on individual $C_{h}$ and $\Sigma_{\eta}$, but impose an i.i.d. assumption on the innovations.
Consider the following example that illustrates the distinction between the single coefficient and joint inference on impulse responses in non-stationary case.
Example 2.
Consider a univariate AR(1) model
$y_{t}=\phi_{1}y_{t-1}+\eta_{t},$ (18)
with $\eta_{t}$ being i.i.d. Gaussian sequence with zero mean and unit variance. The impulse response sequence $C_{h}$ has explicit form $C_{h}=\phi_{1}^{h}$. It is well known that the OLS estimator of $\phi_{1}$ has a non-Gaussian Dickey-Fuller (DF) distribution when $\phi_{1}=1$ (Dickey and Fuller, 1979). Common bootstrap procedures, such as residual bootsrap, are not valid for inference on $\phi_{1}$ and therefore functions of it when $\phi_{1}=1$. However, as proposed in Inoue and Kilian (2002), one can add a time trend or “lag-augment” the model (18),
$y_{t}=\phi_{1}y_{t-1}+\phi_{2}y_{t-2}+\eta_{t},$ (19)
where the restriction $\phi_{2}=0$ is not imposed. The addition of such an irrelevant regressor results in OLS estimates that have an asymptotic Gaussian marginal distribution of individual components but are not jointly Gaussian. Sims et al. (1990) showed that there exist “forbidden” linear combinations of OLS estimates, corresponding to the so-called generalized cointegrating vectors. Indeed, one can rewrite the model in ADF form
$y_{t}=(\phi_{1}+\phi_{2})y_{t-1}+\phi_{2}\Delta y_{t-2}+\eta_{t}.$ (20)
Then, the OLS estimator of $\gamma=(\phi_{1}+\phi_{2})$ will be super-consistent and will have a Dickey-Fuller (DF) asymptotic distribution, while $\hat{\phi}_{2}$ will be asymptotically Gaussian. By the linearity property of the OLS estimators, $\hat{\phi}_{1}=\hat{\gamma}-\hat{\phi}_{2}$. Due to the superconsitency of $\hat{\gamma}$ for 1, the asymptotic distribution of $\hat{\phi}_{1}$ is dominated by the asymptotic Gaussian distribution of $\hat{\phi}_{2}$. As a result, both $\hat{\phi}_{1}$ and $\hat{\phi}_{2}$ have marginal distributions for which the residual bootstrap is valid. Similarly, IRF estimators based on the lag-augmented model (19)
$\hat{C}_{1}-C_{1}$ $=\hat{\phi}_{1}-1=O_{P}(\frac{1}{\sqrt{T}}),$ (21)
$\hat{C}_{2}-C_{2}$ $=\hat{\phi}_{2}+\hat{\phi}^{2}_{1}-1=(\hat{\phi}_{1}-1)+(\hat{\gamma}-1)+(\hat{\phi}_{1}-1)^{2}=\hat{\phi}_{1}-1+O_{P}(\frac{1}{{T}})=O_{P}(\frac{1}{\sqrt{T}}),$ (22)
have marginal asymptotic Gaussian distributions with the standard rate of consistency (since asymptotic distribution of $\hat{C}_{2}$ is dominated by the asymptotically Gaussian component $\hat{\phi}_{1}-1$). However the IRF vector is not jointly Gaussian, since the following difference has non-Gaussian distribution
$\hat{C}_{2}-\hat{C}_{1}=(\hat{\gamma}-1)+(\hat{\phi}_{1}-1)^{2}=O_{P}(\frac{1}{T}).$ (23)
This difference is asymptotically equal to a linear combination of correlated DF random variables and $\chi^{2}(1)$ and is super-consistent for $C_{2}-C_{1}=0$. As a result, the residual bootstrap will not approximate the distribution of the difference $\hat{C}_{2}-\hat{C}_{1}$ even in the point-wise sense.
To summarize, the residual bootstrap procedures (without a unit root pre-test and differencing transformation) cannot be directly used for simultaneous inference on IRF in a potential presence of unit root. Similarly, the avaliable proofs for local projection estimators of IRF study asymptotic Gaussian distribution at individual horizons, but do not prove jointly Gaussian distribution (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023). Both approaches require a conservative Bonferroni correction for the construction of simultaneous confidence bounds on multiple horizons (Inoue et al., 2023). [^6] Montiel Olea and Plagborg-Møller (2019) show that sup-t bounds are less conservative than Bonferroni. Other articles concerned with joint inference on IRF in Bayesian VAR framework include Inoue and Kilian (2022); Arias et al. (2023).
In the following sections, we propose an estimation algorithm based on LP, establish consistency and joint asymptotic normality of LP estimators under Assumptions 2-3. Finally, we propose and prove the validity of a novel dependent wild boostrap procedure under the maintained assumptions. This inference procedure provides a way to conduct simultaneous inference on structural IRF for all horizons without pre-tests.
3 Estimation and inference
3.1 Estimation of IRF using LP
Our goal is to estimate
$\theta(H_{1})\triangleq vec(\Sigma_{\eta},C_{1},\dots,C_{H_{1}},\gamma)$ (24)
and then compute functions of this parameter.
The first step for our analysis is to estimate $C_{h}$ for $h=1,\dots,H$ and $\Sigma_{\eta}$ using the following hybrid local projection estimator.
Algorithm 1
  1. Step 0.
    Pick $H_{1}\geq p$.
  2. Step 1.
    For each $h=1,\dots,H_{1}$ estimate (14) by least squares and store the residuals
    $\hat{\eta}_{t+h}(h)=y_{t+h}-\hat{V}(h)\mu_{t}+\sum_{i=1}^{p}\hat{A}_{i}{(H)}y_{t-i}\text{ for }t=p,\dots,T-h.$ (25)
  3. Step 2.
    Compute
    $\hat{\Sigma}_{\eta}=\frac{1}{T-p}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}.$ (26)
  4. Step 3.
    For each $h=1,\dots,H_{1}$ compute
    $\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h).$ (27)
  5. Step 4.
    Compute the non-normalized vector of structural impacts
    $\hat{\gamma}$ $=\frac{1}{T-p}\sum_{t=p}^{T}(\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{T}),$ (28)
    where $\bar{Z}_{T}=\frac{1}{T}\sum_{t=p}^{T}Z_{t}$.
  6. Step 5.
    For each $i=1,\dots,p$ use backward recursion (BR) to re-estimate
    $\hat{A}^{BR}_{i}=\hat{A}^{BR}_{i}(\hat{\theta}(H_{1}))=\hat{C}_{i}-\sum_{\ell=1}^{i}\hat{A}^{BR}_{\ell}\hat{C}_{i-\ell}$ (29)
    assuming $\hat{C}_{0}=I$ and $\hat{A}^{BR}_{\ell}=0$ for $\ell\leq 0$.
  7. Step 6.
    For each $h=H_{1}+1,\dots,H_{2}$ compute recursively
    $\hat{C}^{BR}_{h}=\hat{C}^{BR}_{h}(\hat{\theta}(H_{1}))=\sum_{\ell=1}^{i}\hat{A}^{BR}_{\ell}\hat{C}^{*}_{i-\ell},$ (30)
    where $\hat{C}^{*}_{h}=\hat{C}_{h}$ for $h=1,\dots,H_{1}$ and $\hat{C}^{*}_{h}=\hat{C}^{BR}_{h}$ for $h\geq H_{1}$.
Algorithm 1 gives estimates of $C_{h}$ for the first $H_{1}$ horizons that are identical to the local projection estimator of Jordà (2005) as Theorem 2 below shows. Since we are interested in simultaneous inference on all relevant parameters of an SVAR, this algorithm imposes the VAR(p) structure on impulse responses of longer horizons. In doing so, we reduce the problem of inference about IRF at long horizons to the problem of simultaneous inference on a finite set of parameters $C_{1},\dots,C_{H_{1}}$. This is a much easier problem and can be addressed by our bootstrap procedure, which will be outlined later.
Note that, in general, the backward recursion estimators of $A_{i}$ are different from the corresponding LS estimators. The LS estimates may be super-consitent for some linear combinations of $A$, but it comes at the expense of non-standard asymptotic distributions that are hard to use for confidence sets (Mikusheva, 2012). On the contrary, $\hat{A}^{BR}_{i}$ are polynomial functions of asymptotically Gaussian estimators $\hat{C}_{h}$. Therefore, a bootstrap procedure can be used to estimate the bias and variance of $\hat{A}^{BR}_{i}$. This approach allows us to bypass any need to pre-test data for presence of deterministic trends, unit roots, and cointegration relations and avoid unnecessary data transformations (which may reduce efficiency and introduce pretesting biases).
Theorem 2.
Suppose Assumptions 2 and 3 hold. Then local projection estimators of $C_{h}$ using regressions
$y_{t+h}=C_{h}y_{t}+\tilde{V}(h)\mu_{t-1}+\sum_{i=1}^{p}\tilde{A}_{i}(h)y_{t-i}+\xi_{t+h}(h),$ (31)
have representation
$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)$ (32)
where $\hat{\eta}_{t}=\hat{\eta}_{t}(0)$ and $\hat{\eta}_{t+h}(h)$ are regression residuals from $h$-step ahead autoregressions defined in (25). Moreover, $\hat{C}_{h}\stackrel{{\scriptstyle p}}{{\rightarrow}}C_{h}$ for all $h$ and $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$.
Proof.
See Appendix B.2
Using estimates $\hat{C}_{h}$ and $\hat{\gamma}$, we can apply the following algorithm to calculate the structural IRF.
Algorithm 2
  1. Step 1.
    Compute the structural impacts of a one-standard deviation shock
    $\hat{b}_{\cdot 1}=\frac{\hat{\gamma}}{\sqrt{\hat{\gamma}^{\prime}\hat{\Sigma}^{-1}_{\eta}\hat{\gamma}}}$ (33)
  2. Step 2.
    For each horizon $h=0,\dots,H$, compute structural IRF using $\hat{\psi}_{h}=\hat{C}_{h}\hat{b}_{\cdot 1}$.
  3. Step 3.
    Compute forecast error variance decomposition for each $r=1,\dots,n$ and $H$ using
    $\widehat{FEVD}_{r}(H)=\frac{\sum_{h=0}^{H}(e^{\prime}_{r}\hat{\psi}_{h})^{2}}{\sum_{h=0}^{H}e^{\prime}_{r}\hat{C}_{h}\hat{\Sigma}_{\eta}\hat{C}^{\prime}_{h}\iota}.$ (34)
Algorithm 2 is consistent for estimation of structural IRF and FEVD for the shock of interest as follows from
Theorem 3.
Suppose that Assumptions 1-3 hold. Then $\hat{b}_{\cdot 1}\stackrel{{\scriptstyle p}}{{\rightarrow}}b_{\cdot 1}$.
Proof.
See Appendix Section B.3. ∎
3.2 HAC-inference
In the previous section, we obtained the asymptotic Gaussian distribution and the corresponding covariance matrix for $\hat{\theta}(H_{1})$. This covariance matrix can be estimated using heteroscedasticity and autocorrealtion consistent (HAC) covariance estimator as the following theorem shows.
Theorem 4.
Suppose Assumptions 1-3 holds. Then
[…]
with $\Omega(H)\triangleq\sum^{\infty}_{\ell=-\infty}\mathbb{E}\Xi_{t}(H)\Xi^{\prime}_{t+\ell}(H)$, where
$\Xi_{t}(H)\triangleq\begin{pmatrix}vec(\eta_{t}\eta_{t}^{\prime}-\Sigma_{\eta})\\ vec(\Sigma_{\eta}^{-1}\eta_{t}\eta_{t+1}^{\prime})\\ \cdots\\ vec(\Sigma_{\eta}^{-1}\sum_{s=0}^{H-1}\eta_{t}\eta_{t+H-s}^{\prime}C_{s}^{\prime})\\ (\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})-\gamma\end{pmatrix}.$ (36)
Proof.
See Appendix Section B.4
The variance-covariance matrix $\Omega$ can be estimated using HAC estimator along the lines in Andrews (1991). We introduce
$\hat{\Xi}_{t}(H)\triangleq\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec(\hat{\Sigma}_{\eta}^{-1}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec(\hat{\Sigma}_{\eta}^{-1}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}.$ (37)
and
$\hat{\Omega}(H)=\sum^{B_{T}}_{\ell=-B_{T}}K(\frac{\ell}{T})\overline{\hat{\Xi}_{t}(H)\hat{\Xi}^{\prime}_{t+\ell}(H)}$ (38)
for some of kernel function $K(\cdot)$ and bandwidth sequence $B_{T}$.
Theorem 5.
Suppose Assumptions 1-3 hold. Then for any bandwidth sequence satisfying $B_{T}^{2}/T\to 0$, we get $\hat{\Omega}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Omega$.
Proof.
See Appendix Section B.5
Even though we had to modify the original proofs of Andrews (1991) to account for non-stationarity of $y_{t}$, the asymptotic theory for optimal selection of kernels and bandwidths still holds. This is due to the fact that the infeasible analog of $\hat{\Omega}(H)$ that uses the true $H$-step innovations $\eta_{t}(H)$ has the same properties regardless of the presence of unit roots and deterministic trends in the VAR model. In particular, one can use the conventional rule $B_{T}=\{0.75\sqrt[3]{T}$ that is appropriate for Bartlett kernel. 777 Recent literature emphasized the fact that when data is very persistent, HAC estimators do not perform well (see, fore example Lazarus et al., 2021). In particular, in the near-to-unity asymptotics, one should use a very large bandwidth $B_{T}$ to control the size of $t$ tests. However, if the innovations are two-sided MDS, then Montiel Olea and Plagborg-Møller (2021) showed that one can use $B_{T}=0$ for inference on individual coefficients $\hat{C}_{h}$. Even under these stronger assumptions, joint inference on the entire vector $\theta(H_{1})$ still requires $B_{T}$ to grow with the sample size to account for potential presence of conditional hetersedasticity of unknown form.
3.3 Dependent wild bootstrap
Using the Delta method, an asymptotic normal distribution can be derived for other estimators, including $\hat{A}^{BR}_{j}$, $\hat{C}^{BR}_{h}$ at long horizons $h$, and $\widehat{FEVD}(h)$. The wild bootstrap approach of Shao (2010) can be used as a convincing numerical implementation of the delta method. A particular way to perform bootstrap is given in the following algorithm.
Algorithm 3
  1. Step 1.
    Select bandwidth $B_{T}$. For example, $B_{T}=0.75\sqrt[3]{T}$
  2. Step 2.
    Generate i.i.d. $\zeta^{s}_{i}\sim N(0,\frac{1}{B_{T}})$ for $i=-B_{T}+1,-B_{T}+2,...,0,1,...T$
  3. Step 3.
    Recursively compute $u^{s}_{t}=\zeta^{s}_{t}+\zeta^{s}_{t-1}+...+\zeta^{s}_{t-B_{T}+1}$ , $t=1,...T$
  4. Step 4.
    Compute
    $\hat{\theta}^{s}(H_{1})=\hat{\theta}(H_{1})+\frac{1}{T-H-p}\sum_{t=p}^{T-H_{1}}\hat{\Xi}_{t}(H_{1})u^{s}_{t}.$ (39)
  5. Step 5.
    Compute $\hat{A}^{BR,s}_{i}=A^{BR}_{i}(\hat{\theta}^{s}(H_{1}))$, $\hat{C}^{BR,s}_{h}=C^{BR}_{h}(\hat{\theta}^{s}(H_{1}))$, $\hat{b}^{s}_{\cdot 1}(\hat{\theta}^{s}(H_{1}))$ using Algorithms 1 and 2.
This algorithm is consistent for estimating the asymptotic distribution of $\hat{\theta}$ as the following theorem shows.
Theorem 6.
Suppose Assumption 1-3 holds and $B_{T}^{2}/T\to 0$. Then
$\lim_{T\to\infty}\sup_{x\in R,\|a\|=1}|P(\sqrt{T}a^{\prime}(\hat{\theta}^{s}(H_{1})-\hat{\theta}(H_{1}))<x|y_{1},...,y_{T})-P(\sqrt{T}a^{\prime}(\hat{\theta}(H_{1})-\theta(H_{1}))<x)|=0.$ (40)
Proof.
See Appendix Section B.6. ∎
To our knowledge, this is the first joint bootstrap resampling procedure for LP estimators. The validity of the dependent wild bootstrap is based on the fact that it mimics the asymptotic Gaussian distribution of $\hat{\theta}(H_{1})$ with a covariance matrix that coincides with the HAC estimator given in Theorem 5. This argument does not impose any parametric model on $\eta_{t}$ and $Z_{t}$ and is generally applicable under the maintained weak dependence assumption (Assumption 2.2). Note that the algorithm uses $\hat{\Xi}_{t}$ that depend on possibly non-stationary processes with strong dependence $Y_{t}$. As a result, conventional block bootstrap procedures are not directly applicable to $\hat{\Xi}_{t}$. Unlike the popular recursive residual-based bootstrap procedures of Gonçalves and Kilian (2004), we do not assume stability (or stationarity) of $Y_{t}$ and allow for asymmetric conditional heteroscedasticity in $\eta_{t}$.
3.4 Simultaneous confidence bounds
[…]
Algorithm 4
  1. 1.
    Use Algoritm 3 to compute $\{\hat{\theta}^{s}\}_{s=1}^{S}$.
  2. 2.
    Estimate standard error for each $h=1,\dots,H$
    $\hat{\sigma}_{h}=(q(f_{h}(\hat{\theta}^{s}),\Phi(1))-q(f_{h}(\hat{\theta}^{s}),1-\Phi(1)))/2$ (41)
    where $q(f_{h}(\hat{\theta}^{s}),\alpha)$ is the order statistics of $\{f_{h}(\hat{\theta}^{s})\}_{s=1}^{S}$ corresponding to the quantile $\alpha$ and $\Phi(\cdot)$ is the standard normal CDF.
  3. 3.
    For each bootstrap draw $s=1,\dots,S$ compute
    $\xi^{s}=\sup_{h=1,\dots,H}\frac{|f_{h}(\hat{\theta}^{s})-f_{h}(\hat{\theta})|}{\hat{\sigma}_{h}}.$ (42)
  4. 4.
    Compute $CV_{\alpha}=q(\xi^{s},1-\alpha)$.
  5. 5.
    Compute
    $[f_{j}(\hat{\theta})-\hat{\sigma}_{h}CV_{\alpha},f_{h}(\hat{\theta})+\hat{\sigma}_{h}CV_{\alpha}].$ (43)
In large samples, by the delta method, the boostrap distribution of $f_{h}(\hat{\theta}^{s})$ is approximately Gaussian. As a result, the difference between the $q(f_{h}(\hat{\theta}^{s}),\Phi(1))$ and $q(f_{h}(\hat{\theta}^{s}),1-\Phi(1))$ quantiles is converging in probablity to 2 standard deviations of $f_{h}(\hat{\theta})$ as $T,S\to\infty$. Step 1 can be replaced with a conventional bootstrap standard error estimator if the bootstrap distribution tails are not fat. Another advantage of using the s.e. estimator based on the quantiles is that the simultaneous bounds coincide with the $68\%$ pointwise bounds if we set the critical value $CV_{\alpha}=1$.
An alternative use of Algorithm 4 is a test of misspecification. Suppose that we have two competing identification assumptions, for example, heteroscedasticity IV and Cholesky. We can construct a joint confidence set for the vector of differences of impacts under these two schemes, $\hat{b}^{Cholesky}_{.1}-\hat{b}^{HetIV}_{.1}$. Alogrithm 4 can then be used to compute $\sup t$ critical values for a test of the hypothesis $b^{Cholesky}_{.1}-b^{HetIV}_{.1}=0$.
We illustrate both uses of Algorithm 4 in Sections 4 and 5.
3.5 Smoothed local projections using minimum distance approach.
Throughout the paper, we assume that the underlying process that generates the data follows a finite order VAR($p$) model. This model imposes particular relations on the matrices $C_{h}$, $h=1,\dots,H$ ($C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ with $A_{\ell}=0$ for $\ell>p$) that can be used to improve efficiency. The backward recursion estimator $\hat{A}^{BR}_{i}$, introduced in Algorithm 1, imposes the VAR structure on $C_{h}$ with $h=1,\dots,p$ but does not use the overidentifying information about $C_{h}$ for $h>p$. One can improve efficincy of the LP estimates using the following miminum distance estimator. Take any positive definite matrix $\hat{W}$ that converges to a positive definite matrix $W$ of dimension $n^{2}\times H_{1}$ and minimize
$(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})=\underset{(A_{1},\dots,A_{p})\in\mathbb{R}^{pn^{2}}}{\operatorname{argmin}}(g(A_{1},\dots,A_{p})-\hat{g}^{LP})^{\prime}\hat{W}(g(A_{1},\dots,A_{p})-\hat{g}^{LP}),$ (44)
where the vector of local projection is denoted as $\hat{g}^{LP}=vec(\hat{C}_{1},\dots,\hat{C}_{H_{1}})$ and $g(A_{1},\dots,A_{p})=vec(C_{1}(A_{1},\dots,A_{p}),\dots,C_{H_{1}}(A_{1},\dots,A_{p}))$ is the corresponding recursively computed IRF coefficients for any values of matrices $(A_{1},\dots,A_{p})$. From the theory of the minimum distance estimators (see, for example, Hansen (2022)) the optimal choice of $\hat{W}$ correspond to the inverse of the long run covariance matrix, given in Theorem 4 earlier in this paper. In small samples, however, the HAC estimator given in Theorem 5 may result in non-invertible matrix that needs to be regularized. In this paper, we instead use a diagonal matrix $\hat{W}$ with vector of inverse variances of the components of $\hat{g}^{LP}$ on the diagonal. These variances are estimated using dependent wild bootstrap.
Based on the estimates of minimum distance $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$, we can compute smoothed impulse responses for any horizon $H$, $(\hat{C}^{MD}_{1},\dots,\hat{C}^{MD}_{H})\triangleq C_{1}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),\dots,C_{H}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$, using the usual recursive formulas.
Under the correct VAR($p$) specification, the limit of (44) coincides with the true parameters $A_{1},\dots,A_{p}$. Moreover, the estimator $((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$ is a smooth transformation of $\hat{\theta}$. So, by the delta method, the dependent wild bootstrap from Theorem 6 remains valid for $\hat{C}^{MD}_{i}$. By implication, we can use it both for point-wise inference and simultaneous inference on $C_{h}$.
It will be evident in Section 5, that the smoothed LP estimates of the IRF are substantially different from the conventional recursive VAR estimates at long horizons (see also Appendix Figure 12). It should also be noticed that conventional recursive VAR estimates of IRF based on the least squares approach are only asymptotically efficient under homoscedastic and Gaussian innovations $\eta_{t}$ (see, for example, Hamilton, 2008). In the presence of conditional heteroscedasticity, the smoothed LP estimators may be more efficient than the conventional recursive VAR estimators of $C_{h}$.
The idea of smoothing local projection is not new. Barnichon and Brownlees (2019) imposed a particular flexible form on the LP estimates within a ridge-regression framework without providing the corresponding inference procedure. In contrast to the earlier studies, we use the VAR structure of the IRF. This allows us to improve efficiency under correct specification and use the dependent wild bootstrap as a robust inference procedure.
4 Simulation experiment based on a DSGE model of US economy
In this section, we would like to illustrate how hetersocedasticity IV identification works together with the dependent wild bootstrap using data created by a dynamic general equilibrium model.
For the data generation process, we will use the well-known model Smets and Wouters (2007) (SW later on), which is a medium-scale New Keynesian model with a wide range of nominal and real rigidities. In particular, the model introduces nominal price and wage rigidity, consumption habits, investment adjustment, and capital utilization costs. At the same time, the Taylor rule for the interest rate assumes that the monetary authority reacts to the deviation of output from the level that would prevail in an economy with flexible prices and wages, which further increases the dimension of the problem, since this requires additionally describing the dynamics of a hypothetical economy with flexible prices.
The state vector in the model has a dimension of about 20 variables. As in the original paper, we introduced 7 structural macroeconomic shocks into the model, including the monetary policy shock. Typically, the model is estimated using the Kalman filter with the observation vector consisting of 7 variables (the number of variables cannot be less than the number of shocks): the log difference of real GDP, real consumption, real investment and the real wage, log hours worked, the log difference of the GDP deflator, and the federal funds rate. The filter is used to estimate the linear state space model that corresponds to the solution of the log-linearized DSGE models. The unobserved state vector follows an SVAR model. Since the observation vector has a smaller dimension than the state vector (7 vs 20), the observation vector is not generally a finite-order vector autoregression. However, in a fairly general case, the vector of observed variables can be represented as a VARMA process, and, assuming invertibility of the MA part, a VAR of finite order can act as an approximation for the corresponding VAR($\infty$).  [^8] The quality of approximation by a finite-order VAR depends on the specific structure of the model, the macroeconomics variables used as observables, and on the persistence of shocks (Ravenna, 2007; Fernandez-Villaverde et al., 2007).
In the experiment of this section, we generate samples of various lengths from the log-linearized SW DSGE model (equivalent to a VARMA model) calibrated using the posterior mean estimates obtained in the (Smets and Wouters, 2007). The only exception was the standard deviation parameter of the monetary policy shock, for which we introduced two regimes described by a Markov chain. We consider a symmetric case where the probability of remaining in one of the regimes is 70%, and the switch to another regime is 30%. Then we introduce an auxiliary instrumental variable $Z_{t}$, which takes the value 0 in the first regime and 1 in the second. The value of the standard deviation of the monetary policy shock, $\sigma_{r,t}$, was set according to the following equation:
$\sigma_{r,t}=\sigma_{r,0}(1+3Z_{t})$ (45)
where $\sigma_{r,0}$ takes a value of posterior mean estimate for the standard deviation of the monetary policy shock in the Smets and Wouters (2007) paper. This specification also assumes that the standard deviation of the monetary shock increases by a factor of 4 in the second regime.
Figure 1: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 160 periods.LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line)
Figure 2: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 1600 periods. LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
Next, using the simulated dataset of seven observable variables and the instrument $Z_{t}$ we apply our new identification method to extract the structural monetary policy shock and compute the corresponding IRFs. We included in the vector of observed variables in local projection model the following variables: log-deviations from the steady state level for real GDP, household consumption, investment, labor, wage, GDP deflator growth rate, and the gross nominal interest rate. To emphasize that the model can handle both stationary and nonstationary variables, we used GDP deflator index rather than growth rates in the econometric exercise. Thus, the model will use 6 stationary variables and 1 non-stationary variable. The original calibration of the model is based on the quarterly frequency of observation. The model includes 4 lags, which corresponds to a lag of 1 year. $Z_{t}$ is used as an instrument for heteroscedasticity of monetary policy shock.
Finally, we compare the IRFs generated from the SW model (VARMA, theoretical) to those obtained from the estimated VAR model to validate our method’s accuracy and effectiveness in capturing the dynamics prescribed by the SW framework. Figure 1 shows the estimation results for 160 periods (40 years), and Figure 2 shows the results for 1600 periods. In Figure 1, it can be seen that the method can reconstruct the IRF for monetary policy shocks within the approximate VAR(4) model using a typical sample size. The true IRF are well within the simultaneous 68% confidence bounds. As we increase the sample size to 1600 periods (Figure 2), the LP estimates are getting even closer to the true values. This shows that LP estimators based on finite VAR(4) can reasonably well approximate the underlying VARMA model. Moreover, the simultaneous confidence bounds successfully covered the true IRF in all the figures, including the non-stationary GDP deflator index.
5 Empirical Application: identificaiton of monetary policy shock in US
Identification of a monetary policy shock is a classical application of structural VAR models that goes back to Sims (1980). This application has been extensively studied with many identification schemes being proposed, including the classical Cholesky decomposition (Sims, 1980; Christiano et al., 1999), instrumental variables (for example, Romer and Romer, 1989, 2004; Gertler and Karadi, 2015), sign restrictions (for example, Uhlig, 2005; Gafarov et al., 2018; Antolín-Díaz and Rubio-Ramírez, 2018), high-frequency identification (for example, Faust et al., 2004), and regime-switching models (for example, Sims and Zha, 2006; Lütkepohl and Netšunajev, 2017) to name a few. [^9] For foundational context and a detailed examination of prior empirical efforts we refer reader to seminal works by Christiano et al. (1999), Boivin et al. (2010), Ramey (2016), which provide comprehensive reviews on the subject. Such an abundance of benchmarks provides an opportunity to test the novel identification approach and validate or reject selected existing alternative methods.
5.1 Data and Model
To compare our identification strategy to those commonly used in the literature, we focus on the reduced-form model studied in seminal papers including, but not limited to Christiano et al. (1999), Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018). Our specification includes six variables for the U.S.: real GDP, GDP deflator, commodity price index, total reserves, nonborrowed reserves, and federal funds rate, described in Table 1. The sample covers the period from January 1965 to November 2007, with monthly data similar to Antolín-Díaz and Rubio-Ramírez (2018). [^10] The monthly series for real GDP and the GDP deflator are interpolated from quarterly data using Chow and Lin (1971) methodology. We use the industrial production index for real GDP interpolation and the consumer price index for the GDP deflator. Our specification includes a constant and a polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. [^11] When we compare results to those from the literature, we mimic number of lags and trends. For example, Christiano et al. (1999) uses constant but no time trend, number of lags 6 and 12. Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018) use 12 lags and no constant or deterministic trends.
Table 1: Description of Variables and Data Sources
Variable Source Seasonally Mnemonics/Link Adjusted Real GDP FRED Yes https://fred.stlouisfed.org/series/GDPC1 GDP Deflator FRED Yes https://fred.stlouisfed.org/series/GDPDEF) Total Reserves FRED Yes https://fred.stlouisfed.org/series/TRARR Nonborrowed Reserves FRED Yes https://fred.stlouisfed.org/series/BOGNONBR Fed Funds Rate FRED No https://fred.stlouisfed.org/series/FEDFUNDS Commodity Price Index GFD No Global Financial Data Consumer Price Index FRED Yes https://fred.stlouisfed.org/series/CPIAUCSL Industrial Production Index FRED Yes https://fred.stlouisfed.org/series/INDPRO
5.2 FOMC meetings and heteroscedasticity of monetary policy shocks
It is well documented in the high-frequency identification literaturethat announcements made after FOMC meetings immediately impact financial variables. This change in financial indicators is associated with unforeseen changes in monetary policy, also referred to as monetary policy shocks.121212 See, for example, Cochrane and Piazzesi (2002), Faust et al. (2004), Gürkaynak et al. (2005), Gertler and Karadi (2015), Ramey (2016), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022)
We focus exclusively on the dates of FOMC meeting announcements including meetings, telephone and video conferences, unscheduled meetings, and sequential day meetings. Our baseline instrumental variable indicates the number of meetings occurring in a given month, including telephone conferences and unscheduled meetings. [^13] An alternative instrumental variable is a dummy with one’s in the months when the meetings occur and zero otherwise. Results are available upon request. Our instrument remains agnostic about the sign and size of the policy change. What matters is the event itself. [^14] We construct series using the FRASER St. Louis Fed website https://fraser.stlouisfed.org/title/federal-open-market-committee-meeting-minutes-transcripts-documents-677?browse=1980s Figure 3 plots the monthly time series of our baseline instrument starting from 1965 till 2007. As one can see the number of the meetings in a given month varies from zero to four.
Figure 3: FOMC meetings
Notes: The blue bars indicate the number of FOMC meetings per month, including telephone conferences and unscheduled meetings
Most of the papers that use FOMC meetings for identification purposes focus their analysis on the sample starting from 1994 onwards. This is done due to the fact the FED releases no public announcements about monetary policy decisions prior to 1994. Some researches, like, Jayawickrema and Swanson (2021) extend the sample backward to 1988. As our identification assumption remains agnostic about the direction of the change, size, we can use the sample back to 1965.
5.3 Results
5.3.1 Benchmark specification
Figure 4 displays the impulse response functions of six endogenous variables to the structural monetary policy shock. We estimate those responses using Heteroskedssticity-IV LP estimator in our benchmark specification with $p=12$ without a trend. The dotted lines represent 68% point-wise confidence intervals, while the dot-dashed lines are simultaneous sup-t bands. Although simultaneous confidence bands are wider relative to point-wise, they are extremely useful, when one, for example, is interested in comparing short-run versus medium-run versus long-run effects. [^15] See section 5.3.4 below for comparison with an alternative simultaneous bounds based on Bonferroni inequality. Recent contribution regarding simultaneous confidence intervals are Jordà (2009), Inoue and Kilian (2016), Montiel Olea and Plagborg-Møller (2019), Inoue and Kilian (2022),Arias et al. (2023).
Figure 4: Identification of monetary policy shock in the US. IRF estimates with point-wise and simultaneous bands. Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
As observed, monetary policy tightening leads to a significant drop in output, which remains significant from 10 to 60 months using point-wise confidence bands. When considering joint significance, this decline is significant from approximately 15 to 35 months, a duration notably longer than previously reported results by Inoue and Kilian (2016). The results on inflation are similar to Wald-joint confidence intervals considered in that paper under a stationary VAR specification. In both cases, the price puzzle disappears once we consider simultaneous confidence intervals. It is important to keep in mind that our procedure of simultaneous confidence intervals construction allows for unit roots without pre-tests, while Inoue and Kilian (2016) requires transforming the model into a stationary one which may cause a pretesting bias.
5.3.2 Sensitivity to number of lags and trends
Our inference methodology allows for a flexible trend specification. Appendix Figures 7-9 compare the benchmark model results (blue lines) with a more flexible models (golden lines): $(k=4,p=12)$, $(k=0,p=24)$, and $(k=4,p=24)$. The polynomial trend can account for slow and gradual changes in the mean values of stationary variables and the growth rates of unit root variables.
Either including a trend or additional lags (or both) results in shorter-lived responses of the real GDP and weaker effects of the monetary shock on inflation and commodity prices. These results show that long-run effects of the monetary policy are sensitive to model specification. One possible explanation is that the trend (or additional lags) absorbs the long-run variation in the real GDP which is otherwise be attributed to monetary policy shocks. As a result, the more flexible specifications only detect the impact of monetary policy shocks that are neutral in the long run. This result is consistent with the panel regression findings at 10 year horizons in Jordà et al. (2020): the monetary tightening shocks can be non-neutral in the long run, while monetary loosening shocks are typically short lived.
In the current literature, it is a convention to exclude the time trend from the monetary SVARs computed in levels (see, for example, Uhlig, 2005; Antolín-Díaz and Rubio-Ramírez, 2018). To facilitate the comparison, we will report only the results for the benchmark specification $(k=0,p=12)$ in the subsequent sections.
5.3.3 Smoothed LP
Appendix Figure 10 presents similar impulse response functions with smoothed point-wise and simultaneous bands for the benchmark specification. Compared to the raw LP estimates, Figure 4, the smoothed IRF results in a stronger impact of the shock on GDP at the long horizons, but a weaker impact on inflation in the long horizons. Notice that the confidence bounds got tighter as a result of the smoothing. This tightening occurs as a result of efficiency gains from imposing restrictions on the LP coefficient under the assumption of a correct specification of the VAR(12) model.
The smoothed LP estimators are based on a minimum distance estimator that imposes the VAR(12) model on the coefficients $C_{h}$. The distance between the VAR(12) model and the raw estimates was weighed with the inverse of the asymptotic variance of the individual LP estimators. Such weights are not fully optimal and could be further improved by using the inverse variance-covariance matrix of LP estimators as in Theorem 5. In practice, however, the optimal weights require some regularization in small samples. We leave this for future research.
5.3.4 Comparison with the union (Bonferroni) bounds
There are alternative existing methods that allow simultaneous inference on LP estimates without knowledge of the joint covariance matrix using so-called union or Bonferroni bounds (Inoue et al., 2023). Such an approach assumes the worst-case dependence between the individual components of the IRF estimates. As discussed in Montiel Olea and Plagborg-Møller (2019), the simultaneous sup-t bounds will be tighter than the union bounds in most cases. Appendix Figure 11 shows IRFs with sup-t versus union bands. Similar to Montiel Olea and Plagborg-Møller (2019) we find that Bonferroni confidence bands are noticeably wider relative to sup-t confidence bands.
5.3.5 Comparison with Cholesky
Figure 5 compares the IRFs to a monetary policy shock identified Heteroscedasticity-IV LP to IRFs from Christiano et al. (1999). Both identification schemes use the smoothed LP IRF. The only difference is the estimation of the impact vector. As can be seen, the results of the new identification approach are very close to those of Christiano et al. (1999), which is a standard Cholesky identification scheme with a priori zeros impacts on real GDP, GDP deflator, and commodity price index. Specifically, the $\sup t$ test for the equality of the impact vectors for the two identification schemes given in Subsection 3.4 has $\sup t=1.2637$ with p-value $0.53$ (critical value at 32% is 1.8083). Without imposing restrictions, we obtain a similar conclusion.
Notice that the confidence bounds for Heteroscedasticity-IV estimators account for uncertainty in the structural rotation matrix (which depends on both $\hat{\gamma}$ and $\hat{\Sigma}_{\eta}$), while the Cholesky approach only accounts for uncertainty in $\hat{\Sigma}_{\eta}$. As a result, the Cholesky IRF has misleadingly tighter confidence bounds that are only correct under the a priori zero restrictions on real GDP, GDP deflator and commodity prices implied by the identificaiton scheme.
The classical paper Christiano et al. (1999) used recursive formulas for the computation of IRF. Appendix Figure 12 compares the results of smoothed LP Cholesky with the conventional VAR estimations under homoscedastic residuals. One can see that smoothed LP point estimates are only similar to the conventional VAR estimates at the short horizons. At the long horizons, the plots for real GDP, GDP deflator, and the commodity price index are substantially different. What is more surprising is that the smoothed LP point-wise confidence bounds are comparable and even tighter than the conventional VAR bootstrap bounds. It highlights the fact that the additional robustness of the proposed procedure does not automatically imply wider confidence bounds than its conventional alternative.
Figure 5: Identification of monetary policy shock in the US. Comparison with Christiano et al. (1999). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and Cholesky-LP estimator (golden dot-dash) with corresponding 68% point-wise bands (blue dots and golden dashes).
5.3.6 Comparison with narrative sign restrictions
In the previous subsection, we found out that the zero restrictions on impacts on real GDP, GDP deflator, and the commodity price index cannot be rejected simultaneously. However, these restrictions were questioned by the literature. Uhlig (2005) replaced these zero restrictions with sign restrictions on impulse responses on short horizons. Such an approach by construction eliminates the price puzzle. In a more recent study by Antolín-Díaz and Rubio-Ramírez (2018) the authors further imposed restrictions on signs of monetary policy shock on particular dates to tighten the bounds. Both papers use Bayesian Gaussian homoscedastic SVAR models with uninformative priors over the structural rotation matrix for inference. [^16] Alternatively, Gafarov et al. (2018) considered frequentist bounds on partially identified sets of IRF using moment inequality methods. This approach results in wider confidence bounds since it is designed to cover the worst-case values of the unidentified structural rotations as opposed to the prior distribution weighted average.
Figure 6: Identification of monetary policy shock in the US. Comparison with narrative sign restrictions (Antolín-Díaz and Rubio-Ramírez, 2018). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and narrative sign restricted BVAR estimator (golden dot-dash) with corresponding 68% point-wise confidence and credibility bands (blue dots and golden dashes)
[…]
There is an interesting discussion in the literature and some researchers argue that there is no price puzzle and this is a reality. For example, Mishkin (2007) argues that an increase in short-term interest rate increases the cost of building new houses and reduces housing activity. Through the supply cost channel that may cause an increase in prices in the short run. This example is from the construction sector, although similar discussions are shared in the agricultural sector. With monetary policy tightening, the cost of mortgages increases, which impacts production costs in the affected sectors.
6 Conclusion
This paper makes several contributions to the literature on SVARs with potential unit roots and co-integration. First, we propose a novel heteroscedasticity IV for monetary policy shocks that requires weaker assumtions than the conventional external IV approaches. As a result, we are able to construct much longer time series of IV. Second, we show how one can perform point-wise and simultaneous inference on IRF with such extenral IV using LP. We specifically generalize the HAC estimator of LP account for the non-statioanry regressors and show equivalence of dependent wild bootstrap with the analytical HAC inference. Third, we use the joint draws of the LP to construct minimum distance smoothed LP IRF estimator and the corresponding confidence bounds.
We then illustrate the usefullness of these procedures in two applications: a simulated data based on a realistic DSGE model and a monetary SVAR for the US. We find that the heterscedasticiy IV based on the number of FOMC meetings in a given month results in IRF that are very close to the classical Christiano et al. (1999) bounds and are different from a comparable BVAR with (narrative) sign restrictions that rule out price puzzle. In particular, our results suggest a much weaker long-run impact on inflation than the BVAR would imply.
Appendix
Appendix A Additional figures
A.1 Sensitivity to model specification
Figure 7: Identification of monetary policy shock in the US. IRF estimates with a flexible trend ($k=4$ and $p=12$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=12$ (golden dot-dash line) with corresponding 68% point-wise bands
Figure 8: Identification of monetary policy shock in the US. IRF estimates with a additional lags ($k=0$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=0,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands
Figure 9: Identification of monetary policy shock in the US. IRF estimates with both flexible trend and additional lags ($k=4$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands
A.2 Smoothed LP
Figure 10: Identification of monetary policy shock in the US. Smoothed IRF estimates with point-wise and simultaneous bands. Heteroscedasticity-IV smoothed LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
A.3 Bonferroni bounds
Figure 11: Identification of monetary policy shock in the US. Sup-t vs Bonferroni simultaneous bands. Heteroscedasticity-IV LP with $k=0,p=12$ estimator (blue solid line), Bonferroni bounds (gold dots) and simultaneous sup-t bounds (dot-dash line)
A.4 Comparison with conventional VAR approch
Figure 12: Identification of monetary policy shock in the US. LP vs VAR using Christiano et al. (1999). Cholesky identification with LP estimator with $k=0,p=12$ (blue line) and homoscedastic VAR estimator (violet solid) with corresponding 68% point-wise bands ( gold dashes and violet dots)
Appendix B Proofs of the main results
B.1 Proof of Theorem 1
Proof.
Step 1. Since
$\Sigma^{-1}_{\eta}=(BB^{\prime})^{-1}=(B^{\prime})^{-1}B^{-1},$ (46)
we have
$B^{\prime}\Sigma^{-1}_{\eta}B=I.$ (47)
In particular, it means that the first column of $B$, $b_{\cdot 1}$, satisfies
$b_{\cdot 1}^{\prime}\Sigma^{-1}_{\eta}b_{\cdot 1}=1$ (48)
Step 2. The reduced form shocks have the following representation for any $k=1,\dots,n$
$\eta_{1t}\eta_{kt}$ $=b_{k1}b_{11}\varepsilon^{2}_{1t}+\sum_{i\neq j}b_{ki}b_{1j}\varepsilon_{it}\varepsilon_{jt}+\sum_{i\neq 1}b_{ki}b_{1i}\varepsilon^{2}_{it}$ (49)
Then by Assumption 1,
$\gamma$ $=b_{\cdot 1}b_{11}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})+\sum_{i\neq j}b_{\cdot i}b_{1j}\mathbb{E}\varepsilon_{it}\varepsilon_{jt}(Z_{t}-\mathbb{E}Z_{t})+\sum_{i\neq 1}b_{\cdot i}b_{1i}\mathbb{E}\varepsilon^{2}_{it}(Z_{t}-\mathbb{E}Z_{t})$ (51)
$=b_{\cdot 1}b_{11}\mathbb{E}(\varepsilon^{2}_{1t}-1)(Z_{t}-\mathbb{E}Z_{t})=b_{\cdot 1}b_{11}\rho$ (52)
By Step 1,
$\sqrt{\gamma^{\prime}\Sigma^{-1}_{\eta}\gamma}=|b_{11}\rho|$ (53)
and
$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma\operatorname{sign}(b_{11}\rho).$ (54)
The statement of the Theorem follows once we assume specific sign normalization for one of the components of $b_{\cdot 1}$. ∎
B.2 Proof of LP consistency
Lemma B.1.
Suppose that Assumptions 2 and 3 hold. Consider LS residuals $\hat{\eta}_{t}(H)$ from $H-$step forward autorgression,
$y_{t+H}$ $=V(H)\mu_{t}+\sum_{i=1}^{p}A_{i}{(H)}y_{t-i}+\eta_{t+H}(H),$ (55)
where $\eta_{t}(H)=\sum_{s=0}^{H}C_{s}\eta_{t-s}$. Then for each $h=0,1,2,\dots,H,$
$\sum_{t=p}^{T-h}\hat{\eta}_{t}(\hat{\eta}_{t+h}(h))^{\prime}-\sum_{t=p}^{T-h}\eta_{t}(\eta_{t+h}(h))^{\prime}=O_{p}(1).$ (56)
Proof.
The proof proceeds in 3 steps. First, we show that (12) can be decomposed into linearly independent univariate AR processes using Jordan decomposition. Then, we show that the LS estimators for individual processes have well-defined asymptotic distributions. Finally, we use invariance of LS estiates with respect to linear transformations to estimate the residual in (56).
Step 1. Notice that $\mu_{t}=(1,t,\dots,t^{k})^{\prime}$ satisfies a first order deterministic difference equation
$\mu_{t}=Q_{\mu}\mu_{t-1}$ (57)
where $Q_{\mu}$ is a lower triangular matrix
$Q_{\mu}\triangleq\begin{bmatrix}1&0&0&0&\cdots&0&0\\ 1&1&0&0&\cdots&0&0\\ 1&2&1&0&\cdots&0&0\\ \vdots&\vdots&&&&\vdots&\vdots\\ 1&{k\choose 2}&\cdots&\cdots&\cdots&{k\choose k-1}&1\end{bmatrix}.$ (58)
Take any solution $W$ to the following linear matrix equation
$W=\sum_{i=1}^{p}A_{i}WQ^{i}_{\mu}+V.$ (59)
Then $y^{\circ}_{t}=y_{t}-W\mu_{t}$ satisfies
$y^{\circ}_{t}=\sum_{i=1}^{p}A_{i}y^{\circ}_{t-1}+\eta_{t},$ (60)
We can now introduce $Y^{\circ}_{t}=vec(y^{\circ}_{t},\dots,y^{\circ}_{t-p+1})$. Then
$\begin{pmatrix}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}=\begin{bmatrix}Q_{y}&0_{np\times k}\\ 0_{k\times np}&Q_{\mu}\end{bmatrix}\begin{pmatrix}Y^{\circ}_{t-1}\\ \mu_{t-1}\end{pmatrix}+\begin{pmatrix}U_{t}\\ 0\end{pmatrix}$ (61)
with the companion matrix
$Q_{y}\triangleq\begin{bmatrix}A_{1}&A_{2}&\dots&A_{p}\\ I&0&\cdots&0\\ \dots&\dots&\dots&\dots\\ 0&0&\cdots&I\\ \end{bmatrix}$ (62)
and $U_{t}=vec(\eta_{t},0,\dots,0)$.
Matrix $Q_{y}$ can be represented in Jordan normal form  [^18] See, for example, Gantmakher (2000).
$Q_{y}=P_{y}J_{y}P_{y}^{-1}.$ (63)
Then, following ideas in Tsay and Tiao (1990) and Sims et al. (1990), we get VAR(1) representation
$\begin{pmatrix}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}=\begin{bmatrix}J_{y}&0_{np\times k}\\ 0_{k\times np}&Q_{\mu}\end{bmatrix}\begin{pmatrix}P_{y}^{-1}Y^{\circ}_{t-1}\\ \mu_{t-1}\end{pmatrix}+\begin{pmatrix}\tilde{U}_{t}\\ 0\end{pmatrix},$ (64)
with $\tilde{U}_{t}=P_{y}^{-1}U_{t}$. Matrix $J_{y}$ has a block-diagonal form $J_{y}=diag\{J_{y,1},\dots,J_{y,g}\}$, where each block $J_{y,i}$ is a $n_{i}\times n_{i}$ matrix
$J_{y,i}=\begin{bmatrix}\lambda_{i}&1&0&\cdots&0\\ 0&\lambda_{i}&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\lambda_{i}&1\\ 0&0&0&0&\lambda_{i}\end{bmatrix}.$ (65)
with $\lambda_{i}$ being $i$-th eigenvalue of $Q_{y}$ (not necessarily distinct) with geometric multiplicity $n_{i}$. Suppose, without loss of generality, that the blocks $J_{y,i}$ are arranged in decreasing order according to the absolute values of the corresponding $\lambda_{i}$. If unit root is present, then start with $\lambda_{1}=1$. By Assumption 3, there are no roots larger than 1 in absolute value.
Now let us find a row permutation matrix $P_{x}$ that rearranges rows of $P_{y}^{-1}Y^{\circ}_{t-1}$ to get
$\tilde{X}_{t}\triangleq\begin{pmatrix}P_{x}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}\triangleq F(L)\nu_{t}$ (66)
where
$F(L)=\begin{bmatrix}F_{11}(L)&0&0&0&\cdots&0&0\\ 0&F_{22}&0&0&\cdots&0&0\\ 0&F_{32}&F_{33}&0&\cdots&0&0\\ 0&F_{42}&F_{43}&F_{44}&\cdots&0&0\\ \vdots&\vdots&&&&\vdots&\vdots\\ 0&F_{G,2}&\cdots&\cdots&\cdots&F_{G,G}&0\\ 0&0&\cdots&\cdots&\cdots&0&I_{k}\end{bmatrix}$ (67)
and $\nu_{t}=(\eta^{\prime}_{t},(\xi^{1}_{t}(\lambda_{1}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{1}))^{\prime},(\xi^{1}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{1}_{t}(\lambda_{b}))^{\prime},\dots,(\xi^{g_{b}}_{t}(\lambda_{1}))^{\prime},1,t,\dots,t^{k})^{\prime}$ with $\xi^{j}_{t}(\lambda)\triangleq\lambda\sum_{s=1}^{t}\xi^{j-1}_{t}(\lambda)$, $\xi^{0}_{t}(\lambda)\triangleq\eta_{t}\lambda^{t}$, $i=1,\dots,b$ $\lambda_{i}$ are unit roots of (12) and $g_{i}$ are their corresponding maximum geometric multiplicities. The blocks $F_{i,j}$ corresponding to two different unit roots $\lambda_{i}$ and $\lambda_{j}$ are 0. The lag matrix $F_{11}(L)$ is obtained by inversion of the stationary rows of (64) and therefore $\sum^{\infty}_{j=0}j^{g}|F_{11j}|<\infty$ and $\sum_{j=0}^{\infty}F_{11j}F_{11j}^{\prime}$ is not singular.
Following Sims et al. (1990); Tsay and Tiao (1990), one can arrange the properly chosen powers of $T^{1/2}$ in a diagonal matrix $\Upsilon_{T}$ such that $\sum_{t=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{t-1}$ have a bounded variance (for example, for stationary components it is $T^{1/2}$, for components containing $\xi^{j}_{t}(\lambda_{2}))$ it is $T^{(1+j)/2}$ etc.).
Step 2. We would like to show now that for all $h=0,1,2,\dots,H$, there exist a diagonal matrix $\Upsilon_{T}$ such that
[…]
where $\phi^{s}$ are tight random vectors and $\tilde{V}$ is a tight random matrix.
Consider $h=0$ first. The proof for deterministic, stationary and ordinary unit root components follows the proofs of Lemmas 1 and 2 in Sims et al. (1990) with a modification required to account for conditionally heteroscedastic shocks $\eta_{t}$. Namely, proof of Lemma 1 Sims et al. (1990) requires replacement of Functional CLT for sums of $\xi^{1}_{t}$ in part (a) and Theorem 2.4(ii) from Chan and Wei (1988) in part (c) for case $p=1$. Under Assumption 2, Theorem 3.1 (De Jong and Davidson, 2000) applied to $\xi^{1}_{t}$ gives the required FCLT under mixing processes assumption. (De Jong and Davidson, 2000, Theorem 4.1) for any pair of components of $\xi^{1}_{t}$ replaces Theorem 2.4(ii) of Chan and Wei (1988) for mixing case (white noise assumption on $\eta_{t}$ ensures that we do not need recenter the sums). For complex unit roots, the limiting distribution is given in Theorem 4.4 of Tsay and Tiao (1990), which also requires replacing FCLT for martingale differences with mixing white noise (De Jong and Davidson, 2000, Theorem 4.1) and stong law of large numbers used in from Chan and Wei (1988) for MDS with Corollary 3.2 in Rio et al. (2017). For our purposes, we do not need explicit form of $\phi$ and $\tilde{V}$.
For $h>0$, the existence of the limiting distributions follows the proof of Theorem 4.5 of Tsay and Tiao (1990).
Step 3. The regression model (55) can be written as
$y_{t+h}=\Theta(h)X_{t-1}+\eta_{t+h}(h),$ (70)
where $\Theta(h)\triangleq(A_{1}(h),A_{2}(h),\dots,A_{p}(h),V(h))$ and $X_{t-1}\triangleq vec(y_{t-1},y_{t-2},\dots,y_{t-p},1,(t-1),\dots(t-1)^{k})$.
Each period $t=p,\dots,(T-h)$ we have
$\hat{\eta}_{t+h}(h)-\eta_{t+h}(h)=(\Theta(h)-\hat{\Theta}(h))X_{t-1}=-X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h).$ (71)
By Step 1, there exist a linear transformation $P$ such that $X_{t}=P^{-1}\tilde{X}_{t}$. So we can get a representation
$X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h)$ $=\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime}(P^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime})^{-1}\sum_{r=p}^{T-h}(P^{-1})\tilde{X}_{r-1}\eta_{r+h}(h),$ (72)
$=\tilde{X}_{t-1}^{\prime}(\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\eta_{r+h}(h).$ (73)
Taking weighted sums over $t$ using matrix $\Upsilon_{T}$ from Step 2, we get representations
$(\sum_{t=p}^{T-h}(\hat{\eta}_{t+h}(h)-\eta_{t+h}(h))\eta_{t}^{\prime})^{\prime}$ $=\sum_{t=p}^{T-h}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T-h}\Upsilon_{T}^{-1}\tilde{X}_{r-1}(\eta_{r+h}(h))^{\prime},$ (74)
$(\sum_{t=p}^{T-h}(\hat{\eta}_{t}-\eta_{t})(\eta_{t+h}(h))^{\prime})^{\prime}$ $=\sum_{t=p}^{T-h}\eta_{t+h}(h)\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta_{r}^{\prime},$ (75)
$\sum_{t=p}^{T-h}(\hat{\eta}_{t}-\eta_{t})(\hat{\eta}_{t+h}(h)-\eta_{t+h}(h))^{\prime}$ $=\sum_{t=p}^{T-h}\eta_{t+h}(h)\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta_{r}^{\prime}.$ (76)
By Step 2, these sums converge in distribution to some random matrices. Convergence in distribution implies that these sums are tight random sequences (Kallenberg, 1997, Lemma 4.8). This means that the sums are $O_{p}(1)$.
To complete the proof, notice
$\sum_{t=p}^{T-h}\hat{\eta}_{t}(\hat{\eta}_{t+h}(h))^{\prime}-\sum_{t=p}^{T-h}\eta_{t}(\eta_{t+h}(h))^{\prime}$ (77)
$=\sum_{t=p}^{T-h}\eta_{t}(\hat{\eta}_{t+h}(h)-\eta_{t+h}(h))^{\prime}+\sum_{t=p}^{T-h}(\hat{\eta}_{t}-\eta_{t})(\eta_{t+h}(h))^{\prime}+\sum_{t=p}^{T-h}(\hat{\eta}_{t}-\eta_{t})(\hat{\eta}_{t+h}(h)-\eta_{t+h}(h))^{\prime}$ (78)
$=O_{p}(1).$ (79)
Proof of Theorem 2.
Step 1. Use Frisch-Waugh-Lovell theorem (Theorem 4.2 in Lovell (1963)) to get a representation
$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}(\hat{\eta}_{t+h}(h))^{\prime}.$ (80)
Step 2. By Lemma B.1
$\hat{C}_{h}^{\prime}$ $=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}(\hat{\eta}_{t+h}(h))^{\prime}=\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\sum_{s=0}^{h}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))$ (81)
$=C_{h}^{\prime}+\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\frac{1}{T}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))+O_{p}(T^{-1})$ (82)
Since $\eta_{t}$ is stationary and mixing and thus sequence $\eta_{t}\eta_{t+h}$ is ergodic with finite fourth moments, so we can apply the ergodic theorem to get the desired result.
The same proof of consistency applies to $\hat{\Sigma}_{\eta}$. ∎
B.3 Proof of consistency of IV estimtor
Proof of Theorem 3.
Consider an individual term in $\hat{\gamma}$,
$\hat{\eta}_{it}\hat{\eta}_{1t}(Z_{t}-\bar{Z}_{T})=$ $\eta_{it}\eta_{1t}(Z_{t}-\bar{Z}_{T})+(\hat{\eta}_{it}-\eta_{it})\eta_{1t}(Z_{t}-\bar{Z}_{T})$ (83)
$+\eta_{it}(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})+(\hat{\eta}_{it}-\eta_{it})(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})$ (84)
First note that by Step 3 in proof of Lemma B.1,
$\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})\eta_{jt}=O_{p}(1)$ (85)
Next, by Assumption 2.5, $Z_{t}\eta_{t}$ is zero meand and is uncorrelated with $\tilde{X}_{t-1}^{\prime}$. Using the same argument as in Step 2 of the proof of Lemma B.1, $\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}$ is a tight random sequence by De Jong and Davidson (2000, Theorem 4.1). So
$(\sum_{t=p}^{T}(\hat{\eta}_{t}-\eta_{t})\eta^{\prime}_{t}Z_{t})^{\prime}=\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r}=O_{p}(1)$ (86)
The last term in (84) can be bounded by Cauchy-Schwarz theorem,
$|\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})|\leq\sqrt{\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})^{2}\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}(Z_{t}-\bar{Z}_{T})^{2}}=O_{p}(T^{1/4}),$ (87)
since
$\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}(Z_{t}-\bar{Z}_{T})^{2}=\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}Z^{2}_{t}-2\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}Z_{t}+\bar{Z}_{T}^{2}\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}.$ (88)
and by Assumptions 2.2 and 2.3 and Lemma B.1
$\max_{t=p,\dots,T}|Z_{t}|$ =$4]{\max_{t=p,\dots,T}|Z_{t}|^{4}}\leq T^{1/4}\sqrt[4]{\sum_{t=p}^{T}\frac{|Z_{t}|^{4}}{T}}=O_{p}(T^{1/4}),$ (89)
$\max_{t=p,\dots,T}|Z_{t}|^{2}$ $=O_{p}(T^{1/2}),$ (90)
$\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}$ $=O_{p}(1).$ (91)
By Assumption 2, the time series $\eta_{t}\eta_{1t}$ and $Z_{t}$ satisfy CLT for mixing processes. Therefore,
$\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\bar{Z}_{T})$ $=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(Z_{t}-\mathbb{E}Z_{t})\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})$ (92)
$=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(1)$ (93)
To summarize,
$\hat{\gamma}=\frac{1}{T-p}\sum_{t=p}^{T}(\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{T})=$ $\frac{1}{T-p}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(T^{-\frac{3}{4}}).$ (94)
By ergodicity of $(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})$, we get $\hat{\gamma}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$. The statement of the theorem then follows from Theorems 1 and 2.
B.4 Proof of asymptotic normality of the LP estimators
Proof of Theorem 4.
By Assumption 2, $\Xi^{H}_{t}$ is a strong mixing process of size at least $-2(1+\epsilon)/\epsilon$ with zero mean and finite $2+\epsilon$ moments (condition DMR is satisfied, see p. 12 in Rio et al. (2017)). By Theorem 4.2 in Rio et al. (2017) and the Cramer-Wold theorem, there exist the long run variance matrix $\Omega$ for $\Xi^{H}_{t}$ and
$\frac{1}{\sqrt{T-p-H}}\sum_{t=p}^{T-H}\Xi_{t}(H)\to N(0,\Omega).$ (95)
By Theorems 2 and 3,
$\sqrt{T-p}(\hat{\Sigma}_{\eta}-\Sigma_{\eta})$ $=\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-\frac{1}{2}}),$ (96)
$\sqrt{T-p-h}(\hat{C}_{h}^{\prime}-C_{h}^{\prime})$ $=\big{(}\frac{1}{T-p-h}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}\big{)}^{-1}(\frac{1}{\sqrt{T-p-h}}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime})+O_{p}(T^{-\frac{1}{2}}),$ (97)
$\sqrt{T-p}(\hat{\gamma}-\gamma)$ $=\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\mathbb{E}Z_{t})+O_{p}(T^{-1/4}).$ (98)
Since l.h.s. of (95) differs from (35) by some $O_{p}(T^{-\frac{1}{4}})$ terms, the statement of the theorem follows from the continuous mapping theorem (Mann and Wald, 1943a).
B.5 Proof of consistency of HAC standard errors
Proof of Theorem 5.
Step 1.To show consitency of the whole matrix $\hat{\Omega}=\hat{\Omega}(H_{1})$, we need to establish consistency of HAC estimators for all combinations $a^{\prime}\Xi_{t}$. Then the HAC estimator for particular $a$ is defined as
$\hat{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\hat{\Xi}_{t})(\hat{\Xi}_{t+\ell}^{\prime}a)$ (99)
This estimator has an infeasible analog
$\tilde{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\Xi_{t})(\Xi_{t+\ell}^{\prime}a)$ (100)
Since $\Xi_{t}=\Xi_{t}(H_{1})$ is strong mixing with $4+\epsilon$ moment, Theorem 1(a) in Andrews (1991) implies that $\tilde{\Omega}(a)$ is consistent for $\Omega(a)=a^{\prime}\Omega a$.
Step 2. What remains to show is $\hat{\Omega}(a)-\tilde{\Omega}(a)\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. Since by Theorem 2, $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$, we just need to establish HAC consitency for vectors
$\Psi_{t}=\begin{pmatrix}\hat{\psi}^{0}_{t}\\ \hat{\psi}^{1}_{t}\\ \cdots\\ \hat{\psi}^{H_{1}}_{t}\\ \hat{\psi}^{H_{1}+1}_{t}\end{pmatrix}=\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}$ (101)
The last component, $\hat{\xi}^{H_{1}+1}_{t}$, has a different structure from the other components since it is multiplied by $(Z_{t}-\bar{Z}_{t})$. From the proof of Theorem 3, we know that $\max_{t=1,\dots T}|Z_{t}-\bar{Z}_{t}|=O_{p}(T^{1/4})$. So, the conistency of the HAC components, including the linear combinations of $\hat{\psi}^{H_{1}+1}_{t}$ can be proven in a similar fashion as the other components. So, to save space, we only consider vector $a=vec(a_{0},a_{1},\dots,a_{H_{1}},0_{n})$ that puts 0 on the last $n$ components and all $a_{i}\in R^{n^{2}}$.
$\hat{\Psi}_{t}^{\prime}aa^{\prime}\hat{\Psi}_{t+\ell}-\Psi_{t}^{\prime}aa^{\prime}\Psi_{t+\ell}$ $=\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}\hat{\psi}^{s}_{t+\ell}$ (102)
$+\sum_{h,s=0}^{H}(\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})+\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})$ (103)
The first term after summation over $t$ becomes
$\sum_{t=p}^{T-\ell}\sum_{h,s=0}^{H_{1}}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a_{s}\hat{\psi}^{s}_{t+\ell}=$ (104)
$=\sum_{t=p}^{T-\ell}\sum_{i,j=1}^{k}\sum_{h,s=0}^{H_{1}}e^{\prime}_{i}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)}-\eta_{t}\eta_{t+h}(h)+\mathbb{E}\eta_{t}\eta^{\prime}_{t+h}(h))^{\prime}a_{ih}$ (105)
×$\times a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s))e_{j}$ (106)
$=\sum_{i,j=1}^{k}\sum_{h,s=0}^{H_{1}}\sum_{t=p}^{T-\ell}e^{\prime}_{i}(\hat{\eta}_{t}\hat{\eta}_{t+h}(h)-\eta_{t}\eta_{t+h}(h))^{\prime}a_{ih}a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s))e_{j}$ (107)
$-\sum_{i,j=1}^{k}\sum_{h,s=0}^{H_{1}}e^{\prime}_{i}(\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)}-\mathbb{E}\eta_{t}\eta^{\prime}_{t+h}(h))a_{ih}\sum_{t=p}^{T-\ell}a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t}\eta^{\prime}_{t+\ell+s}(s))e_{j}$ (108)
The last term is $O_{p}(1)$ by Lemma B.1 and FCLT for mixing processes (De Jong and Davidson, 2000) applied to $\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t}\eta^{\prime}_{t+\ell+s}(s)$.
Next, consider one element of the sum in the first term. By (74)-(76) we get
$\sum_{t=p}^{T-\ell}e^{\prime}_{i}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)-\eta_{t}\eta^{\prime}_{t+h}(h))a_{ih}$ (109)
$=\sum_{t=p}^{T-\ell}a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)e_{j}e^{\prime}_{i}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}\Lambda_{T}^{h}a_{ih}$ (110)
$+2\sum_{t=p}^{T-\ell}a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)e_{j}e^{\prime}_{i}\eta_{t+h}(h)\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}\Lambda_{T}^{0}a_{ih}$ (111)
where $\Lambda_{T}^{h}\triangleq(\Upsilon_{T}^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}(\sum_{r=p}^{T-h}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r+h}(h))=O_{p}(1)$ by the argument in the proof of Lemma B.1. Both terms in the r.h.s of the displayed formula above are $O_{P}(\sqrt{T-B_{T}})$ by Cauchy-Schwarz inequality since, for example, for any constant vectors
$|\sum_{t=p}^{T-\ell}a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)e_{j}e^{\prime}_{i}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}b|$ (112)
$\leq\sqrt{b^{\prime}(\sum_{t=p}^{T-\ell}\Upsilon_{T}^{-1}\tilde{X}_{t-1}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1})b}\sqrt{\sum_{t=p}^{T-\ell}(a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)e_{j}e^{\prime}_{i}\eta_{t})^{2}}.$ (113)
The r.h.s. is $O_{p}(\sqrt{T})$ by proof of Step 2 of Lemma B.1 and Markov inequality implied by existence of $6+\epsilon$ moments of the stationary mixing sequence $\eta_{t}$.
The sums of remaining two terms in (103) are also $O_{p}(\sqrt{T})$, which can be shown using the same techniques.
Step 3. Putting all bounds together gives us
$\hat{\Omega}-\tilde{\Omega}=\frac{B_{T}}{T-B_{T}}O_{p}(\sqrt{T-B_{t}})=O_{p}(\frac{B_{T}}{\sqrt{T}}).$ (114)
Since by assumption of the theorem $B_{T}^{2}/T\to 0$, we get the desired result.
B.6 Proof of Wild bootstrap
Proof of Theorem 6.
By construction, $\sqrt{T}(\hat{\theta}^{s}-\hat{\theta})\sim N(0,\hat{Q}_{T})$ conditional on the data $y_{1},...,y_{T}$. Let us find the covariance matrix $\hat{Q}_{T}$.
$\hat{Q}_{T}$ $=\frac{1}{(T-p-H)}\sum_{t,l=p}^{T-H}\hat{\Xi}_{t}(H_{1})\hat{\Xi}^{\prime}_{l}(H_{1})\mathbb{E}u^{s}_{t}u^{s}_{l}=\frac{1}{(T-p-H)^{2}}\sum_{t,l=p}^{T-H}\hat{\Xi}_{t}(H_{1})\hat{\Xi}^{\prime}_{|t-l|}(H_{1})K(|t-l|/B_{T})$ (115)
$=\sum_{\ell=-B_{T}}^{B_{T}}\frac{K(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})-\sum_{\ell=-B_{T}}^{B_{T}}\frac{\ell K(\ell/B_{T})}{T(T-\ell)}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})=\hat{J}_{T}+O_{P}(\frac{B_{T}}{T})$ (116)
Since $\frac{B_{T}}{T}\to 0$ , by Theorem 5 we get $\hat{Q}_{T}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Omega$. So by Theorem 4,
$\lim_{T\to\infty}\sup_{x\in R,\|a\|=1}|P(\sqrt{T}a^{\prime}(\hat{\theta}^{s}(H_{1})-\hat{\theta}(H_{1}))<x|y_{1},...,y_{T})-P(\sqrt{T}a^{\prime}(\hat{\theta}(H_{1})-\theta(H_{1}))<x)|$ (117)
$\leq$ $\lim_{T\to\infty}\sup_{x\in R,\|a\|=1}|\Phi(\frac{x}{\sqrt{a^{\prime}\hat{Q}_{T}a}})-\Phi(\frac{x}{\sqrt{a^{\prime}\Omega a}})|$ (118)
$+\lim_{T\to\infty}\sup_{x\in R,\|a\|=1}|P(\frac{\sqrt{T}a^{\prime}(\hat{\theta}(H_{1})-\theta(H_{1}))}{\sqrt{a^{\prime}\Omega a}}<x)-\Phi(\frac{x}{\sqrt{a^{\prime}\Omega a}})|=0.$ (119)

### v2

@@
-Policymaker meetings as heteroscedasticity shifters:
+
 Abstract
-We propose a novel approach to identification in structural vector autoregressions (SVARs) that uses external instruments for heteroscedasticiy of a structural shock of interest.
-This approach does not require lead/lag exogeneity for identification, does not require heteroskedasticity to be persistent, and facilitates interpretation of the structural shocks.
-To implement this identification approach in applications, we develop a new method for simultaneous inference of structural impulse responses and other parameters, employing a dependent wild-bootstrap of local projection estimators.
-This method is robust to an arbitrary number of unit roots and cointegration relationships (including seasonal unit roots), time-varying local means and drifts, and conditional heteroskedasticity of unknown form and can be used with other identification schemes, including Cholesky and the conventional external IV.
-We show how to construct pointwise and simultaneous confidence bounds for structural impulse responses and how to compute smoothed local projections with the corresponding confidence bounds.
-Using simulated data from a standard log-linearized DSGE model, we show that the method can reliably recover the true impulse responses in realistic datasets.
-As an empirical application, we adopt the proposed method in order to identify monetary policy shock using the dates of Federal Open Market Committee (FOMC) meetings in a standard six-variable VAR.
-The robustness of our identification and inference methods allows us to construct an instrumental variable for monetary policy shock that dates back to 1965.
-The resulting impulse response functions (IRFs) for all variables align with the classical Cholesky identification scheme and are different from the narative sign restricted Bayesian VAR estimates.
-In particular, the response to inflation manifests a price puzzle that is indicative of the cost channel of the interest rates.
-Keywords: SVAR, conditional heteroscedasticity, unit roots, cointegration, identification, simultaneous inference, smoothed local projections, narative sign restrictions, minimum distance estimator, monetary policy shocks.
+Structural vector autoregressions are used to compute impulse response functions (IRF) for persistent data.
+Existing multiple-parameter inference requires cumbersome pretesting for unit roots, cointegration, and trends with subsequent stationarization.
+To avoid pretesting, we propose a novel dependent wild bootstrap procedure for simultaneous inference on IRF using local projections (LP) estimated in levels in possibly nonstationary and heteroscedastic SVARs.
+The bootstrap also allows efficient smoothing of LP estimates.
+We study IRF to US monetary policy identified using FOMC meetings count as an instrument for heteroscedasticity of monetary shocks.
+We validate our method using DSGE model simulations and alternative SVAR methods.
+Keywords: SVAR, external instruments, conditional heteroscedasticity, unit roots, cointegration, identification, simultaneous inference, smoothed local projections, narrative sign restrictions, minimum distance estimator, monetary policy shocks.
 1 Introduction
 Unexpected actions by policymakers provide a valuable source of variation that helps to identify how the economy would react to future policy actions (Sims, 1980).
-Unfortunately, it is often difficult to detect the unexpected component of the actions using the publicly available data.
-The true plan of future actions and commitments of the policy makers can be kept secret from the public, and the pure unexpected part of the available policy announcements is always a matter of modeler’s assumptions.
-We propose an identification approach that remains agnostic about the actual size and directions of change in the policy variable (for example, interest rate) due to policymakers’ discretion.
+Unfortunately, it is often difficult to detect the unexpected component of the actions using publicly available data.
+The true plan of future actions and commitments of the policymakers can be kept secret from the public, and the pure unexpected part of the available policy announcements is always a matter of modelers’ assumptions.
+We propose an identification approach that remains agnostic about the actual size and direction of change in the policy variable (for example, interest rate) due to policymakers’ discretion.
 Instead, we only assume that the conditional variance of the unexpected change in the policy variable is higher during policymakers’ meetings than at other times.
-We show that the information contained in the meetings dates is sufficient to uniquely identify both the size and direction of impulse responses to the structural policy shock in a vector autoregression model (VAR).
-Our paper applies this novel approach to identification of the monetary policy shocks using the number of the US Federal Reserve Open Market Committee (FOMC) meetings per month as an instrument for heteroskedasticity of a monetary policy shock in a standard six-variable VAR model.
-Information available on the meetings dates goes back to at least 1965, which gives us instruments with sample sizes that were unprecedented in the monetary SVAR literature.
+We show that the information contained in the meeting dates is sufficient to uniquely identify both the size and direction of impulse responses to the structural policy shock in a vector autoregression model (VAR).
+Our paper applies this novel approach to the identification of monetary policy shocks using the number of US Federal Reserve Open Market Committee (FOMC) meetings per month as an instrument for heteroscedasticity of a monetary policy shock in a standard six-variable VAR model.
+The information available on the meeting dates goes back to at least 1965, which gives us instruments with sample sizes that were unprecedented in the monetary SVAR literature.
 Such long samples come with an additional set of complications.
-In particular, the inference procedure needs to be robust to the potential presence of time trends and unit roots, which makes the existing popular methods like residual-based bootstrap (Gonçalves and Kilian, 2004) inapplicable.
-Moreover, the use of the external instruments for the conditional heteroscedasticity requires a novel procedure that would allow joint inference on all the relevant parameters in the model.
-We address this challenge by proposing a novel inference procedure for the structural impulse responses (IRF) based on local projection (LP) estimators of Jordà (2005).
-Specifically, we propose both a formula and a proof of joint asymptotic normality of the local projection estimator in the potential presence of trends, cointegration and any number of unit roots.
-Next, we extend the proof of consistency of the joint HAC covariance estimator of Andrews (1991) to accomodate nonstationary regressors.
+[^2] According to (Ramey, 2016), the longer sample period, on the one hand, includes more variation in interest rate and thus advantageous for monetary policy shock identification; on the other hand, it is accompanied by extra challenges such as structural breaks, regime shifts, state-dependence, and sign-dependence.
+In particular, the inference procedure needs to be robust to the potential presence of time trends and unit roots, which makes the existing popular methods like the residual-based bootstrap (Gonçalves and Kilian, 2004) inapplicable.
+Moreover, the use of external instruments for conditional heteroscedasticity requires a novel procedure that allows joint inference on all the relevant parameters in the model.
+We address this challenge by proposing a novel inference procedure for structural impulse responses (IRF) based on local projection (LP) estimators of Jordà (2005).
+Specifically, we propose both a formula and a proof of joint asymptotic normality of the local projection estimator in the potential presence of trends, cointegration, any number of unit roots, and any weakly-dependent stationary conditionally heteroscedastic innovations.
+Next, we extend the proof of the consistency of the joint HAC covariance estimator of Andrews (1991) to accommodate non-stationary regressors.
 Finally, we propose a version of the dependent wild bootstrap procedure of Shao (2010) that can mimic the joint distribution of the local projection estimators across multiple horizons and the estimates of structural parameters.
-The only regularity conditions that we impose on the non-explosive VAR model is that the one-step ahead forecast errors, jointly with the instrumental variable, follow a stationary vector process with a mixing property and certain higher-order moments necessary for consistency of the HAC estimator.
-To our knowledge, this provides the first proof of validity of inference on the IRF estimators that allows the general case of unstable but non-explosive VAR models with a general form of stationary conditional heteroscedasticity without pretests.
-[^2] Our results build upon the long literature on the non-stationary time series including Sims et al.
+The only regularity conditions that we impose on the non-explosive VAR model is that the one-step-ahead forecast errors, jointly with the instrumental variable, follow a stationary vector process with a mixing property and existence of certain higher-order moments, which are necessary for consistency of the HAC estimator.
+To our knowledge, this provides the first proof of the validity of inference on the IRF estimators that allows for the general case of unstable but nonexplosive VAR models with a general form of stationary conditional heteroscedasticity without pretests.
+[^3] Our results are built upon the long literature on non-stationary time series including Sims et al.
 (1990); Tsay and Tiao (1990); De Jong and Davidson (2000); Inoue and Kilian (2002); Montiel Olea and Plagborg-Møller (2021).
@@
 First, we estimate reduced-form impulse response functions (i.e.
-responses to one-step ahead linear forecast errors $\hat{\eta}_{it}$ for each time series $i$).
+responses to one-step-ahead linear forecast errors $\hat{\eta}_{it}$ for each time series $i$).
 This step is computed using local projection regressions.
-Second, we estimate contemporaneous impact of the structural shock of interest on a time series $i$ as the sample covariance between pairwise products of one-step ahead forecast errors $\hat{\eta}_{1t}\hat{\eta}_{it}$ and an instrumental variable correlated with the heteroscedasticity of the structural shock of interest.
-(The time series with index $1$ plays a special role since it is assumed to have non-zero correlation of squared forecast errors with the instrumental variable.) Third, we compute structural IRF for other horizons as products of the reduced form IRF and vector the contemporaneous structural responses after appropriate normalization.
-The dependent wild-bootstrap is then applied to the relevant score functions.
+Second, we estimate the contemporaneous impact of the structural shock of interest on time series $i$ as the sample covariance between pairwise products of one-step-ahead forecast errors $\hat{\eta}_{1t}\hat{\eta}_{it}$ and an instrumental variable that is correlated with the heteroscedasticity of the structural shock of interest.
+(The time series with index $1$ plays a special role since it is assumed to have a non-zero correlation of squared forecast errors with the instrumental variable.) Third, we compute structural IRF for other horizons as products of the reduced-form IRF and vector of contemporaneous structural responses after appropriate normalization.
+The dependent wild bootstrap is then applied to the relevant score functions.
 The novel simultaneous inference on local projections has multiple uses.
@@
 (2016), it is important to make joint draws of the vector of the IRF coefficients and the structural impact vector to obtain valid confidence limits.
-Second, the procedure allows for simultaneous confidence bounds that cover the entire IRF across all horizons and, for example, allows for inference on the largest and smallest values of IRF.
-Third, one can compute smoothed local projection estimators and the corresponding confidence bounds using the minimum distance approach to improve efficiency of the estimates.
+Second, the procedure allows simultaneous confidence bounds that cover the entire IRF across all horizons and, for example, allows for inference on the largest and smallest values of IRF.
+Third, one can compute smoothed local projection estimators and the corresponding confidence bounds using the minimum distance approach to improve the efficiency of the estimates.
 We illustrate the proposed method in two applications.
-First, we show that the identification strategy can recover the true IRF of a monetary policy shock in simulated data based on a medium-sized DSGE model of US economy of Smets and Wouters (2007).
-Then we apply the method to study the impact of monetary policy shocks in six-variable VAR model using the data set of Antolín-Díaz and Rubio-Ramírez (2018).
-The variables include Real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and Federal funds rate.
-The sample period is from January 1965 to November 2007.
+First, we show that the identification strategy can recover the true IRF of a monetary policy shock in simulated data based on a medium-sized DSGE model of the US economy of Smets and Wouters (2007).
+Then we apply the method to study the impact of monetary policy shocks in a six-variable VAR model using the data set of Antolín-Díaz and Rubio-Ramirez (2018).
+The variables include Real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and the Federal funds rate.
+The sample period is January 1965 to November 2007.
 Our primary instrumental variable for conditional heteroscedasticity of the monetary policy shocks is the number of FOMC meetings per month, including telephone conferences and unscheduled meetings.
-For reduced form estimates, we use local projection with polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$.
-We report various confidence bands: point-wise, simultaneous sup-t, smoothed point-wise and smoothed simultaneous bands.
-Our main empirical findings can be summarized as follows: (i) the identification through the heteroscedasticity IV results in IRF that are identical to the classical Cholesky identification a la Christiano et al.
-(1999);(ii) smoothed LP-IRF are noticeably different from the conventional recursive VAR estimates at the long horizons; (iii) effects on real GDP are comparable with BVAR identified using the narrataive sign restriction effects (Antolín-Díaz and Rubio-Ramírez, 2018), while the effects on the monetary shocks on inflation are much weaker in the novel scheme.
-The literature on identification through hetersocedasticity has a long history (for example, Feenstra (1994); Rigobon (2003); Lewis (2021)).
-To our knowledge, we are the first to propose an IV-type procedure that assumes that the conditional heteroscedasticity is not-persistent (c.f.
-Rigobon, 2003; Lewis, 2021), or structural shocks are fully independent and non-Gaussian (c.f.
+For reduced-form estimates, we use local projection with polynomial time trends with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$.
+We report various confidence bands: point-wise, simultaneous sup-t, smoothed point-wise, and smoothed simultaneous bands.
+Our main empirical findings can be summarized as follows: (i) the identification through the heteroscedasticity IV results in IRFs that are nearly identical to the classical Cholesky identification a la Christiano et al.
+(1999); (ii) smoothed LP-IRFs are noticeably different from the conventional recursive VAR estimates at the long horizons; (iii) effects on real GDP are comparable with BVAR identified using the narrative sign restriction effects (Antolín-Díaz and Rubio-Ramirez, 2018), while the effects of the monetary shocks on inflation are much weaker in the novel scheme.
+The literature on identification through heteroscedasticity has a long history (for example, Feenstra (1994); Rigobon (2003); Lewis (2021); Brunnermeier et al.
+(2021)).
+To our knowledge, we are the first to propose an IV-type procedure that assumes that the conditional heteroscedasticity is not persistent, or structural shocks are fully independent and non-Gaussian (c.f.
 Lütkepohl and Netšunajev, 2017).
@@
 In contrast to statistical identification procedures, our approach does not require full identification of all shocks and gives an intuitive interpretation of the identified shock of interest.
-In this paper, we apply our identification method to monetary policy shocks, but the external heterscedasticity IV can be proposed in other applications.
+In this paper, we apply our identification method to monetary policy shocks, but the heteroscedasticity-based external IV can be proposed in other applications.
+[^4] There is a long-standing literature devoted to the identification of monetary policy shocks, including but not limited to Christiano et al.
+(1999), Romer and Romer (2004), Gürkaynak et al.
+(2005), Gertler and Karadi (2015), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022), Bu et al.
+(2021), Swanson and Jayawickrema (2023).
 For example, Känzig (2021) shows that OPEC meetings have high-frequency impacts on oil prices that can be used to identify oil supply shocks; Janzen et al.
-(2018) show that cotton storage levels change the heteroskedasticity of precautionary demand shocks.
-So we believe that our approach can be used more broadly.
-Our inference procedure is based on the insight that both the reduced-form impulse response and covariance estimators involved in the structural identification can be represented as a mean of vector-valued stationary time series of cross-products of the $1$-step and $h$-step ahead forecast errors.
+(2018) show that cotton storage levels change the heteroscedasticity of precautionary demand shocks.
+Therefore, we believe that our approach can be used more broadly.
+Our inference procedure is based on the insight that both the reduced-form impulse response and covariance estimators involved in the structural identification can be represented as a mean of vector-valued stationary time series of cross-products of the $1$-step-ahead and $h$-step-ahead forecast errors.
 This representation allows us to employ the dependent wild bootstrap of Shao (2010) to make draws from the joint asymptotic distribution of the estimators without explicitly computing the long-run variance-covariance matrix estimator.
@@
 Section 2 discusses the setup and the novel identification idea.
-Section 3 provides the estimation and inference algorithms with the corresponding theoretical proofs.
+Section 3 provides the estimation and inference algorithms with the corresponding theorems.
 Section 4 discusses the DSGE experiment.
-Section 5 has the empirical application to US monetary policy.
+Section 5 has an empirical application to US monetary policy.
 Section 6 concludes.
-All the proofs and additional figures are provided in Appendix.
-2 Model and the identification
+All the proofs and additional figures are provided in the appendix.
+2 Identification in SVAR models
 2.1 Motivating example
@@
 where observations $\eta_{t}\in\mathbb{R}^{n}$ and $\varepsilon_{t}$ is an i.i.d.
-zero-mean process with $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\prime}=I_{n}$ that is refereed to as structural shocks.
-[^3] The simultaneous equations model with lags were first proposed and studied in Haavelmo (1943); Mann and Wald (1943b).
+zero-mean process with $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\prime}=I_{n}$ that is referred to as structural shocks.
+[^5] The simultaneous equations model with lags was first proposed and studied in Haavelmo (1943); Mann and Wald (1943b).
 Suppose that we are interested in measuring the impact of the structural shock $\varepsilon_{1t}$ on the vector of observables $\eta_{t}$, which is given by the first column of $B=(b_{\cdot 1},\dots,b_{\cdot n})$.
-We can estimate $\Sigma_{\eta}\triangleq\mathbb{E}\eta_{t}\eta_{t}^{\prime}$ using the sample analog and the solve the equation $\Sigma_{\eta}=BB^{\prime}$ for the first column of $B$.
+We can estimate $\Sigma_{\eta}\triangleq\mathbb{E}\eta_{t}\eta_{t}^{\prime}$ using the sample analog and then solve equation $\Sigma_{\eta}=BB^{\prime}$ for the first column of $B$.
 However, there is a continuum of solutions to this equation.
 Therefore, additional assumptions or data are required to uniquely select $B$.
-The purpose of the paper is to propose a novel way of identifying $b_{\cdot 1}$ using external instruments for the conditional heteroscedasticity process of $\varepsilon_{t}$ and develop the inference procedure.
+The paper aims to propose a novel way of identifying $b_{\cdot 1}$ using external instruments for the conditional heteroscedasticity process of $\varepsilon_{t}$ and to develop the inference procedure.
 This identification problem occurs in the more general SVAR models, which we will study later in the paper.
-Before we state the novel identification proposal in the general form, let us consider a simple motivating example.
+Before we generalize the novel identification proposal, let us consider a simple motivating example.
 Example 1.
@@
 $\varepsilon_{it}=\xi_{it}\sigma_{it},$ (2)
-such that $\xi_{it}$ is i.i.d.
+such that $\xi_{it}$ is an i.i.d.
 sequence with $\mathbb{E}\xi_{t}=0$ and $\mathbb{E}\xi_{t}\xi^{\prime}_{t}=I_{2}$; the process $\sigma_{it}$ is a stationary ergodic process that is independent of $\xi_{it}$.
 The volatility process $\sigma^{2}_{1t}$ for the shock of interest may be correlated with some observable time series of $Z_{t}$.
-Suppose that this time series $Z_{t}$ only increases volatility of the first structural shock, but does not change volatility for other shocks.
-We can then use it as an instrument for identification of $b_{\cdot 1}$.
+Suppose that this time series $Z_{t}$ only increases volatility of the first structural shock but does not change volatility for other shocks.
+We can then use it as an instrument for the identification of $b_{\cdot 1}$.
 In contrast to the conventional proxy-VAR identification, the instrument $Z_{t}$ may be independent of $\xi_{1t}$.
 Indeed, suppose that
-$\mathbb{E}(Z_{t}-\mathbb{E}Z_{t})(\sigma^{2}_{it}-\mathbb{E}\sigma^{2}_{it})=\rho_{i}$ (3)
+$\mathbb{E}(Z_{t}-\mathbb{E}Z_{t})(\sigma^{2}_{it}-\mathbb{E}\sigma^{2}_{it})=\rho_{i},$ (3)
 such that $\rho_{1}\neq 0$ and $\rho_{2}=0$ and $\xi_{t}$ is independent of $Z_{t}$. Then since
-$\eta^{2}_{1t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t}).$ (4)
+$\eta^{2}_{1t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t}),$ (4)
 $\eta_{1t}\eta_{2t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{21}\varepsilon_{1t}+b_{22}\varepsilon_{2t}),$ (5)
 we get
-$\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}^{2}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}^{2}\rho_{1}$ (6)
+$\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}^{2}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}^{2}\rho_{1},$ (6)
 $\mathbb{E}\eta_{1t}\eta_{2t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}b_{21}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}b_{21}\rho_{1}.$ (7)
@@
 Thus, we can recover structural impulse responses to the first structural shock $\varepsilon_{1t}$.
-The same logic applies if $Z_{t}$ is correlated with $\xi_{t}$ as long as $\rho_{1}\neq 0$, $\rho_{2}=0$ and $\mathbb{E}\xi_{1t}\xi_{2t}(Z_{t}-\mathbb{E}Z_{t})=0$. ∎
+The same logic applies if $Z_{t}$ is correlated with $\xi_{t}$ as long as $\rho_{1}\neq 0$, $\rho_{2}=0$ and $\mathbb{E}\xi_{1t}\xi_{2t}(Z_{t}-\mathbb{E}Z_{t})=0$.
 The identification conditions in Example 1 can be generalized to
@@
 This assumption implies that the heteroscedasticity of the innovations for this time series must be correlated with the instrument $Z_{t}$.
-In the monetary shock example, one can by definition assume that the impact $b_{11}$ is non-zero for federal funds rate ($i=1$).
-The instrument can only change the conditional variance of the shock of interest and cannot result in conditional correlation of the other structural shocks.
+In the monetary shock example, one can by definition assume that the impact $b_{11}$ is nonzero for the federal funds rate ($i=1$).
+The instrument can only affect the conditional variance of the shock of interest and cannot impact the conditional correlation of the other structural shocks.
 Theorem 1.
@@
 See Appendix B.1. ∎
-A recent work Lewis (2021) showed that in the case when heterscedasticity of a structural shock is persistent, one can use internal instruments $Z_{t}=\eta^{2}_{t-1}$.
+A recent work (Lewis, 2021) showed that in the case when heteroscedasticity of a structural shock is persistent, one can use internal instruments $Z_{t}=\eta^{2}_{t-1}$.
 However, the approach can be applied more generally if appropriate external instruments are available, as in our empirical application.
@@
 $y_{t}=V\mu_{t}+A_{1}y_{t-1}+\cdots+A_{p}y_{t-p}+\eta_{t},$ (12)
-where $y_{t}$ is $n$-dimensional time series, $V\mu_{t}$ is a deterministic polynomial time trend (or time-varying drift for integrated processes) with $\mu_{t}=(1,t,t^{2},\dots,t^{k})$ for some $k$; $A_{1},\dots,A_{p},V$ are $(n\times n)$ parameter matrices, $\eta_{t}$ are serially uncorrelated forecast errors (innovations).
+where $y_{t}$ is $n$-dimensional time series, $V\mu_{t}$ is a deterministic polynomial time trend (or time-varying drift for integrated processes) with $\mu_{t}=(1,t,t^{2},\dots,t^{k})^{\prime}$ for some $k$; $A_{1},\dots,A_{p},V$ are $(n\times n)$ parameter matrices, $\eta_{t}$ are serially uncorrelated forecast errors (innovations).
 We will assume that $p$ is finite and known.
@@
   1. 1.
-    $\eta_{t}$ is a martingale difference sequence (MDS), i.e. $\forall$ $t$ $\mathbb{E}[\eta_{t}|\mathcal{F}_{t-1}]=0$
+    $\eta_{t}$ is a martingale difference sequence (MDS), i.e. $\forall$ $t$ $\mathbb{E}[\eta_{t}|\mathcal{F}_{t-1}]=0$.
   2. 2.
@@
   3. 3.
-    $\mathbb{E}||\eta_{t}||^{8+2\epsilon}<\infty$ and $\mathbb{E}|Z|^{4+\epsilon}<\infty$ for some $\epsilon>0$.
+    $\mathbb{E}||\eta_{t}||^{8+2\epsilon}<\infty$ and $\mathbb{E}|Z_{t}|^{4+\epsilon}<\infty$ for some $\epsilon>0$.
   4. 4.
@@
 Assumption 2.1 implies that innovations are white noise, which is required for the consistency of LS estimators in stationary VAR models.
-The martingale difference requirement with respect to the sigma algebra $\mathcal{F}_{t-1}=\sigma(Y_{t-1},Y_{t-2},\dots)$ is only necessary for interpretation of the impulse response coefficients that will be introduced later in this section.
+The martingale difference requirement with respect to sigma field $\mathcal{F}_{t-1}=\sigma(y_{t-1},y_{t-2},\dots)$ is only necessary for interpretation of the impulse response coefficients that will be introduced later in this section.
 Assumption 2.2 allows for many stationary models of conditional heteroscedasticity.
@@
 This assumption, for example, allows for asymmetric conditional heteroscedasticity models like the ones considered in Rabemananjara and Zakoian (1993).
-Such models are ruled out by popular residual wild bootsrap methods (Gonçalves and Kilian, 2004).
+Such models are ruled out by the popular residual wild bootstrap methods (Gonçalves and Kilian, 2004).
 We do not consider unconditional heteroscedasticity in innovations, since it would eliminate the need for structural identification.
-The 8-th moment condition imposed in our paper is sufficient for consistency of the HAC estimators, but a weaker 4-th moment condition would be sufficient for asymptotic normality of the LP estimators.
+The 8-th moment condition imposed in our paper is sufficient for the consistency of the HAC estimators, but a weaker 4-th moment condition would be sufficient for the asymptotic normality of the LP estimators.
 Assumption 2.4 is standard in papers on structural identification of the shocks.
 Assumption 2.5 is required for asymptotic normality of $\hat{\gamma}$ and would be satisfied, for example, if $\eta_{t}$ is MDS with respect to a sigma field $\mathcal{F}_{t-1}$ that also includes $Z_{t}$ and past values of $\eta_{t}$ (for example, it is possible that the number of meetings in month $t$ is fully predictable based on the past economic conditions and the past number of meetings).
-This assumption is not required for consistency of $\hat{\gamma}$ and the structural IRF.
+This assumption is not required for the consistency of $\hat{\gamma}$ and the structural IRF.
 Using the VAR model, one can define impulse responses for individual innovations $\eta_{t}$ since
-$C_{h,i,j}=\mathbb{E}(y_{t+h}|\eta_{t}=e_{j},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\eta_{t}=0,\mathcal{F}_{t-1})$ (13)
-where $e_{j}\in R^{n}$ is vector of zeros with $1$ in $j-$th position. These coefficients appear as regression coefficients in $H$-step ahead autoregression representation of (12),
+$C_{h,i,j}\triangleq\mathbb{E}(y_{t+h}|\eta_{t}=e_{j},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\eta_{t}=0,\mathcal{F}_{t-1}),$ (13)
+where $e_{j}\in R^{n}$ is a vector of zeros with $1$ in $j-$th position. These coefficients appear as regression coefficients in the $H$-step ahead autoregression representation of (12),
 $y_{t+H}$ $=V(H)\mu_{t}+\sum_{i=1}^{p}A_{i}{(H)}y_{t-i}+\eta_{t+H}(H),$ (14)
@@
 $det(I_{n}-A_{1}z-\dots-A_{p}z^{p})\neq 0$ for $|z|<1$.
-Assumption 3 allows for arbitrary number of unit roots, including complex unit roots, and cointegrating relations between $y_{t}$.
+Assumption 3 allows for an arbitrary number of unit roots, including complex unit roots, and cointegrating relations between $y_{t}$.
 It only rules out explosive roots.
-In practical terms, it means that one can apply our inference procedure to local projections estimated with or without trend in levels without seasonal adjustment.
-It can be done without any pretests for trend stationarity, abnsency of cointegration, or unit roots.
+In practical terms, it means that one can apply our inference procedure to local projections estimated with or without a trend in levels without seasonal adjustment.
+It can be done without any pretests for trend stationarity, absence of cointegration, or unit roots.
 We allow both a polynomial trend as in Sims et al.
 (1990) and complex unit roots as in Tsay and Tiao (1990).
-Under this assumption, the least squares estimator for the coefficients in equations (12) and (14) is consistent and has well-defined distribution limits as shown in Sims et al.
-(1990); Tsay and Tiao (1990) under conditionally homosecdastic MDS innovations $\eta_{t}$.
-In contrast, we allow for more general white noise innovations with mixing by using strong laws of large numbers and functional central limit theorems for strong mixing sequences (De Jong and Davidson, 2000; Rio et al., 2017).
+Under this assumption, the least squares estimator for the coefficients in equations (12) and (14) is consistent and has well-defined distribution limits, as shown in Sims et al.
+(1990); Tsay and Tiao (1990) under conditionally homoscedastic MDS innovations $\eta_{t}$.
+In contrast to the earlier studies, we allow for more general (weak) white noise innovations with mixing by using strong laws of large numbers and functional central limit theorems for strong mixing sequences (De Jong and Davidson, 2000; Rio et al., 2017).
 Using the coefficients $C_{0},\dots,C_{H}$ under Assumption 2.1, we can compute structural impulse response function of $y_{t+h}$ to $\varepsilon_{1t}$, which is given by
@@
 The coefficients $C_{1},\dots,C_{H}$ can be estimated using local projection estimators of Jordà (2005).
-However, our empirical strategy uses conditional heteroskedasticity in the structural shocks in the dataset that has unit root time series with potential cointegration and time trends.
-Furthermore, structural impulse responses depend on both estimators of $\Sigma_{\eta}$ and $C_{1},\dots,C_{H}$ and on covariances with external instruments $Z_{t}$.
+However, our empirical strategy uses conditional heteroscedasticity in the structural shocks in the dataset that has unit root time series with potential cointegration and time trends.
+Furthermore, structural impulse responses depend on both estimators of $(\Sigma_{\eta},C_{1},\dots,C_{H})$ and the covariances of external instruments $Z_{t}$.
 Existing inference methods cannot take into account all of these features together.
-Specifically, there are no proposals for joint inference of $\Sigma_{\eta}$ and $C_{0},\dots,C_{H}$ without pretests for unit roots.
+Specifically, there are no proposals for joint inference of $\Sigma_{\eta}$ and $C_{0},\dots,C_{H}$ without pre-tests for unit roots.
 [^7] Brüggemann et al.
-(2016) provide an example where commonly used residual wild-bootstrap is invalid for joint inference on $C_{i}$ and $\Sigma_{\eta}$ even in the stationary case.
+(2016) provide an example where the commonly used residual wild-bootstrap is invalid for joint inference on $C_{i}$ and $\Sigma_{\eta}$ even in the stationary case.
 Their proposed procedure, residual block bootstrap, solves the issue for stationary VARs.
-Jentsch and Lunsford (2022) then extends this residual block bootstrap approach to allow for external instrumental variables, also in the stationary case.
-Inoue and Kilian (2002); Mikusheva (2012) allow for joint inference on individual $C_{h}$ and $\Sigma_{\eta}$, but impose an i.i.d.
-assumption on the innovations.
-Consider the following example that illustrates the distinction between the single coefficient and joint inference on impulse responses in non-stationary case.
+Jentsch and Lunsford (2022) then extends this residual block bootstrap approach to allow external instrumental variables, also in the stationary case.
+Inoue and Kilian (2002); Mikusheva (2012) allow inference on individual components of $C_{h}$ in the presence of a unit root of a particular type, but impose an i.i.d.
+assumption on the innovations which is violated in our application.
+Consider the following example that illustrates the distinction between the single coefficient and joint inference on impulse responses in the nonstationary case.
 Example 2.
@@
 It is well known that the OLS estimator of $\phi_{1}$ has a non-Gaussian Dickey-Fuller (DF) distribution when $\phi_{1}=1$ (Dickey and Fuller, 1979).
-Common bootstrap procedures, such as residual bootsrap, are not valid for inference on $\phi_{1}$ and therefore functions of it when $\phi_{1}=1$.
+The common bootstrap procedures, such as residual bootstrap, are not valid for inference on $\phi_{1}$ and therefore functions of it when $\phi_{1}=1$.
 However, as proposed in Inoue and Kilian (2002), one can add a time trend or “lag-augment” the model (18),
@@
 where the restriction $\phi_{2}=0$ is not imposed.
-The addition of such an irrelevant regressor results in OLS estimates that have an asymptotic Gaussian marginal distribution of individual components but are not jointly Gaussian.
+The addition of such irrelevant regressor results in OLS estimates that have an asymptotic Gaussian marginal distribution of individual components but are not jointly Gaussian.
 Sims et al.
@@
 $y_{t}=(\phi_{1}+\phi_{2})y_{t-1}+\phi_{2}\Delta y_{t-2}+\eta_{t}.$ (20)
-Then, the OLS estimator of $\gamma=(\phi_{1}+\phi_{2})$ will be super-consistent and will have a Dickey-Fuller (DF) asymptotic distribution, while $\hat{\phi}_{2}$ will be asymptotically Gaussian.
+Then, the OLS estimator of the corresponding population parameter $\gamma=(\phi_{1}+\phi_{2})$ will be super-consistent and will have a Dickey-Fuller (DF) asymptotic distribution, while $\hat{\phi}_{2}$ will be asymptotically Gaussian.
 By the linearity property of the OLS estimators, $\hat{\phi}_{1}=\hat{\gamma}-\hat{\phi}_{2}$.
-Due to the superconsitency of $\hat{\gamma}$ for 1, the asymptotic distribution of $\hat{\phi}_{1}$ is dominated by the asymptotic Gaussian distribution of $\hat{\phi}_{2}$.
+Due to the super-consistency of $\hat{\gamma}$ for 1, the asymptotic distribution of $\hat{\phi}_{1}$ is dominated by the asymptotic Gaussian distribution of $\hat{\phi}_{2}$.
 As a result, both $\hat{\phi}_{1}$ and $\hat{\phi}_{2}$ have marginal distributions for which the residual bootstrap is valid.
@@
 $\hat{C}_{2}-C_{2}$ $=\hat{\phi}_{2}+\hat{\phi}^{2}_{1}-1=(\hat{\phi}_{1}-1)+(\hat{\gamma}-1)+(\hat{\phi}_{1}-1)^{2}=\hat{\phi}_{1}-1+O_{P}(\frac{1}{{T}})=O_{P}(\frac{1}{\sqrt{T}}),$ (22)
-have marginal asymptotic Gaussian distributions with the standard rate of consistency (since asymptotic distribution of $\hat{C}_{2}$ is dominated by the asymptotically Gaussian component $\hat{\phi}_{1}-1$).
-However the IRF vector is not jointly Gaussian, since the following difference has non-Gaussian distribution
+have marginal asymptotic Gaussian distributions with the standard rate of consistency (since the asymptotic distribution of $\hat{C}_{2}$ is dominated by the asymptotically Gaussian component $\hat{\phi}_{1}-1$).
+However, the IRF vector is not jointly Gaussian, since the following difference has a non-Gaussian distribution
 $\hat{C}_{2}-\hat{C}_{1}=(\hat{\gamma}-1)+(\hat{\phi}_{1}-1)^{2}=O_{P}(\frac{1}{T}).$ (23)
@@
 As a result, the residual bootstrap will not approximate the distribution of the difference $\hat{C}_{2}-\hat{C}_{1}$ even in the point-wise sense.
-To summarize, the residual bootstrap procedures (without a unit root pre-test and differencing transformation) cannot be directly used for simultaneous inference on IRF in a potential presence of unit root.
-Similarly, the avaliable proofs for local projection estimators of IRF study asymptotic Gaussian distribution at individual horizons, but do not prove jointly Gaussian distribution (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023).
+To summarize, the residual bootstrap procedures (without a unit root pre-test and differencing transformation) cannot be directly used for simultaneous inference on IRF in the potential presence of a unit root.
+Similarly, the available proofs for local projection estimators of IRF study asymptotic Gaussian distribution at individual horizons, but do not prove jointly Gaussian distribution (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023).
 Both approaches require a conservative Bonferroni correction for the construction of simultaneous confidence bounds on multiple horizons (Inoue et al., 2023).
 [^8] Montiel Olea and Plagborg-Møller (2019) show that sup-t bounds are less conservative than Bonferroni.
-Other articles concerned with joint inference on IRF in Bayesian VAR framework include Inoue and Kilian (2022); Arias et al.
+Other articles concerned with joint inference on IRF in the Bayesian VAR framework include Inoue and Kilian (2022); Arias et al.
 (2023).
-In the following sections, we propose an estimation algorithm based on LP, establish consistency and joint asymptotic normality of LP estimators under Assumptions 2-3.
-Finally, we propose and prove the validity of a novel dependent wild boostrap procedure under the maintained assumptions.
+In the following sections, we propose an estimation algorithm based on LP, and establish consistency and joint asymptotic normality of LP estimators under Assumptions 2-3.
+Finally, we propose and prove the validity of a novel dependent wild bootstrap procedure under the maintained assumptions.
 This inference procedure provides a way to conduct simultaneous inference on structural IRF for all horizons without pre-tests.
@@
 3.1 Estimation of IRF using LP
-Our goal is to estimate
+Our goal is to estimate the vector of parameters
 $\theta(H_{1})\triangleq vec(\Sigma_{\eta},C_{1},\dots,C_{H_{1}},\gamma)$ (24)
-and then compute functions of this parameter.
-The first step for our analysis is to estimate $C_{h}$ for $h=1,\dots,H$ and $\Sigma_{\eta}$ using the following hybrid local projection estimator.
+and then compute the functions of this parameter.
+The first step for our analysis is to estimate $C_{h}$ for $h=1,\dots,H$, and $\Sigma_{\eta}$ using the following hybrid local projection estimator.
 Algorithm 1
@@
     For each $h=1,\dots,H_{1}$ estimate (14) by least squares and store the residuals
-    $\hat{\eta}_{t+h}(h)=y_{t+h}-\hat{V}(h)\mu_{t}+\sum_{i=1}^{p}\hat{A}_{i}{(H)}y_{t-i}\text{ for }t=p,\dots,T-h.$ (25)
+    $\hat{\eta}_{t+h}(h)=y_{t+h}-\hat{V}(h)\mu_{t}-\sum_{i=1}^{p}\hat{A}_{i}{(h)}y_{t-i}\text{ for }t=p,\dots,T-h.$ (25)
   3. Step 2.
     Compute
-    $\hat{\Sigma}_{\eta}=\frac{1}{T-p}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}.$ (26)
+    $\hat{\Sigma}_{\eta}=\frac{1}{T-p}\sum_{t=p}^{T}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}.$ (26)
   4. Step 3.
@@
     $\hat{\gamma}$ $=\frac{1}{T-p}\sum_{t=p}^{T}(\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{T}),$ (28)
-    where $\bar{Z}_{T}=\frac{1}{T}\sum_{t=p}^{T}Z_{t}$.
+    where $\bar{Z}_{T}\triangleq\frac{1}{T-p}\sum_{t=p}^{T}Z_{t}$ and $\overline{\hat{\eta}_{t}\hat{\eta}_{1t}}\triangleq\frac{1}{T-p}\sum_{t=p}^{T}\hat{\eta}_{t}\hat{\eta}_{1t}$.
   6. Step 5.
@@
   7. Step 6.
-    For each $h=H_{1}+1,\dots,H_{2}$ compute recursively
-    $\hat{C}^{BR}_{h}=\hat{C}^{BR}_{h}(\hat{\theta}(H_{1}))=\sum_{\ell=1}^{i}\hat{A}^{BR}_{\ell}\hat{C}^{*}_{i-\ell},$ (30)
+    For each $h=H_{1}+1,\dots,H_{2}$ compute recursively as follows
+    $\hat{C}^{BR}_{h}=\hat{C}^{BR}_{h}(\hat{\theta}(H_{1}))=\sum_{\ell=1}^{h}\hat{A}^{BR}_{\ell}\hat{C}^{*}_{h-\ell},$ (30)
     where $\hat{C}^{*}_{h}=\hat{C}_{h}$ for $h=1,\dots,H_{1}$ and $\hat{C}^{*}_{h}=\hat{C}^{BR}_{h}$ for $h\geq H_{1}$.
-Algorithm 1 gives estimates of $C_{h}$ for the first $H_{1}$ horizons that are identical to the local projection estimator of Jordà (2005) as Theorem 2 below shows.
-Since we are interested in simultaneous inference on all relevant parameters of an SVAR, this algorithm imposes the VAR(p) structure on impulse responses of longer horizons.
+Algorithm 1 gives estimates of $C_{h}$ for the first $H_{1}$ horizons.
+These estimates are identical to the local projection estimator of Jordà (2005) as Theorem 2 below shows.
+Since we are interested in simultaneous inference on all relevant parameters of an SVAR, this algorithm imposes the VAR(p) structure on impulse responses of longer horizons, $H_{1}\leq h\leq H_{2}$.
 In doing so, we reduce the problem of inference about IRF at long horizons to the problem of simultaneous inference on a finite set of parameters $C_{1},\dots,C_{H_{1}}$.
-This is a much easier problem and can be addressed by our bootstrap procedure, which will be outlined later.
+This is a much easier problem, and our bootstrap procedure, which will be outlined later, can address it.
 Note that, in general, the backward recursion estimators of $A_{i}$ are different from the corresponding LS estimators.
-The LS estimates may be super-consitent for some linear combinations of $A$, but it comes at the expense of non-standard asymptotic distributions that are hard to use for confidence sets (Mikusheva, 2012).
-On the contrary, $\hat{A}^{BR}_{i}$ are polynomial functions of asymptotically Gaussian estimators $\hat{C}_{h}$.
-Therefore, a bootstrap procedure can be used to estimate the bias and variance of $\hat{A}^{BR}_{i}$.
-This approach allows us to bypass any need to pre-test data for presence of deterministic trends, unit roots, and cointegration relations and avoid unnecessary data transformations (which may reduce efficiency and introduce pretesting biases).
+The LS estimates may be super-consistent for some linear combinations of $A$, but come at the expense of non-standard asymptotic distributions that are hard to use for confidence sets (Mikusheva, 2012).
+In contrast, $\hat{A}^{BR}_{i}$ are polynomial functions of asymptotically Gaussian estimators $\hat{C}_{h}$.
+Therefore, a bootstrap procedure can estimate the bias and variance of $\hat{A}^{BR}_{i}$.
+This approach allows us to bypass any need to pre-test data for the presence of deterministic trends, unit roots, and cointegration relations and avoid unnecessary data transformations.
+Such pretests may introduce pretesting biases and reduce efficiency of the resulting estimators if the transformations turn out to be redundant.
 Theorem 2.
-Suppose Assumptions 2 and 3 hold. Then local projection estimators of $C_{h}$ using regressions
-$y_{t+h}=C_{h}y_{t}+\tilde{V}(h)\mu_{t-1}+\sum_{i=1}^{p}\tilde{A}_{i}(h)y_{t-i}+\xi_{t+h}(h),$ (31)
+Suppose Assumptions 2 and 3 hold. Then the local projection estimators of $C_{h}$ using regressions
+$y_{t+h}=C_{h}y_{t}+\tilde{V}(h)\mu_{t-1}+\sum_{i=1}^{p}\tilde{A}_{i}(h)y_{t-i}+\xi_{t+h}(h)$ (31)
 have representation
-$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h)$ (32)
-where $\hat{\eta}_{t}=\hat{\eta}_{t}(0)$ and $\hat{\eta}_{t+h}(h)$ are regression residuals from $h$-step ahead autoregressions defined in (25).
+$\hat{C}_{h}^{\prime}=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(h),$ (32)
+where $\hat{\eta}_{t}\triangleq\hat{\eta}_{t}(0)$ and $\hat{\eta}_{t+h}(h)$ are regression residuals from $h$-step ahead autoregressions defined in (25).
 Moreover, $\hat{C}_{h}\stackrel{{\scriptstyle p}}{{\rightarrow}}C_{h}$ for all $h$ and $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$.
@@
     For each horizon $h=0,\dots,H$, compute structural IRF using $\hat{\psi}_{h}=\hat{C}_{h}\hat{b}_{\cdot 1}$.
-  3. Step 3.
-    Compute forecast error variance decomposition for each $r=1,\dots,n$ and $H$ using
-    $\widehat{FEVD}_{r}(H)=\frac{\sum_{h=0}^{H}(e^{\prime}_{r}\hat{\psi}_{h})^{2}}{\sum_{h=0}^{H}e^{\prime}_{r}\hat{C}_{h}\hat{\Sigma}_{\eta}\hat{C}^{\prime}_{h}\iota}.$ (34)
-Algorithm 2 is consistent for estimation of structural IRF and FEVD for the shock of interest as follows from
+Algorithm 2 is consistent for the estimation of structural IRF for the shock of interest as follows from
 Theorem 3.
@@
 See Appendix Section B.3. ∎
+Remark 1.
+One can also compute forecast error variance decomposition for each $r=1,\dots,n$ and $H$ using formula
+$\widehat{FEVD}_{r}(H)=\frac{\sum_{h=0}^{H-1}(e^{\prime}_{r}\hat{\psi}_{h})^{2}}{\sum_{h=0}^{H}e^{\prime}_{r}\hat{C}_{h}\hat{\Sigma}_{\eta}\hat{C}^{\prime}_{h}\iota},$ (34)
+where $\iota$ is a vector of ones.
 3.2 HAC-inference
 In the previous section, we obtained the asymptotic Gaussian distribution and the corresponding covariance matrix for $\hat{\theta}(H_{1})$.
-This covariance matrix can be estimated using heteroscedasticity and autocorrealtion consistent (HAC) covariance estimator as the following theorem shows.
+This covariance matrix can be estimated using heteroscedasticity and autocorrelation consistent (HAC) covariance estimator as the following theorem shows.
 Theorem 4.
@@
 […]
-with $\Omega(H)\triangleq\sum^{\infty}_{\ell=-\infty}\mathbb{E}\Xi_{t}(H)\Xi^{\prime}_{t+\ell}(H)$, where
-$\Xi_{t}(H)\triangleq\begin{pmatrix}vec(\eta_{t}\eta_{t}^{\prime}-\Sigma_{\eta})\\ vec(\Sigma_{\eta}^{-1}\eta_{t}\eta_{t+1}^{\prime})\\ \cdots\\ vec(\Sigma_{\eta}^{-1}\sum_{s=0}^{H-1}\eta_{t}\eta_{t+H-s}^{\prime}C_{s}^{\prime})\\ (\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})-\gamma\end{pmatrix}.$ (36)
+with $\Omega(H_{1})\triangleq\sum^{\infty}_{\ell=-\infty}\mathbb{E}\Xi_{t}(H_{1})\Xi^{\prime}_{t+\ell}(H_{1})$, where
+$\Xi_{t}(H_{1})\triangleq\begin{pmatrix}vec(\eta_{t}\eta_{t}^{\prime}-\Sigma_{\eta})\\ vec(\Sigma_{\eta}^{-1}\eta_{t}\eta_{t+1}^{\prime})\\ \cdots\\ vec(\Sigma_{\eta}^{-1}\sum_{s=0}^{H_{1}-1}\eta_{t}\eta_{t+H_{1}-s}^{\prime}C_{s}^{\prime})\\ (\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})-\gamma\end{pmatrix},$ (36)
+and $\eta_{1t}$ is the first element of $\eta_{t}$.
 Proof.
 See Appendix Section B.4
-The variance-covariance matrix $\Omega$ can be estimated using HAC estimator along the lines in Andrews (1991). We introduce
+The variance-covariance matrix $\Omega$ can be estimated using the HAC estimator along the lines in Andrews (1991). We introduce
 $\hat{\Xi}_{t}(H)\triangleq\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec(\hat{\Sigma}_{\eta}^{-1}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec(\hat{\Sigma}_{\eta}^{-1}(\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}.$ (37)
@@
 See Appendix Section B.5
-Even though we had to modify the original proofs of Andrews (1991) to account for non-stationarity of $y_{t}$, the asymptotic theory for optimal selection of kernels and bandwidths still holds.
-This is due to the fact that the infeasible analog of $\hat{\Omega}(H)$ that uses the true $H$-step innovations $\eta_{t}(H)$ has the same properties regardless of the presence of unit roots and deterministic trends in the VAR model.
-In particular, one can use the conventional rule $B_{T}=\{0.75\sqrt[3]{T}$ that is appropriate for Bartlett kernel.
-777 Recent literature emphasized the fact that when data is very persistent, HAC estimators do not perform well (see, fore example Lazarus et al., 2021).
+Even though we had to modify the original proofs of Andrews (1991) to account for the nonstationarity of $y_{t}$, the asymptotic theory for optimal selection of kernels and bandwidths still holds.
+This is because the infeasible analog of $\hat{\Omega}(H)$ that uses the true $H$-step innovations $\eta_{t}(H)$ has the same properties regardless of the presence of unit roots and deterministic trends in the VAR model.
+In particular, one can use the conventional rule $B_{T}=\{0.75\sqrt[3]{T}\}$ that is appropriate for the Bartlett kernel.
+Recent literature emphasized the fact that when data is very persistent, HAC estimators do not perform well (see, for example Lazarus et al., 2021).
 In particular, in the near-to-unity asymptotics, one should use a very large bandwidth $B_{T}$ to control the size of $t$ tests.
-However, if the innovations are two-sided MDS, then Montiel Olea and Plagborg-Møller (2021) showed that one can use $B_{T}=0$ for inference on individual coefficients $\hat{C}_{h}$.
-Even under these stronger assumptions, joint inference on the entire vector $\theta(H_{1})$ still requires $B_{T}$ to grow with the sample size to account for potential presence of conditional hetersedasticity of unknown form.
+This is not the case under our Assumption 2.2 (weak dependence) for the innovations $\eta_{t}$.
+If, in addition, the innovations are two-sided MDS, then Montiel Olea and Plagborg-Møller (2021) showed that one can use $B_{T}=0$ for inference on individual coefficients $\hat{C}_{h}$.
+However, even under these stronger assumptions, the joint inference on the entire vector $\theta(H_{1})$ still requires $B_{T}$ to grow with the sample size to account for the potential presence of conditional heteroscedasticity of unknown form.
 3.3 Dependent wild bootstrap
-Using the Delta method, an asymptotic normal distribution can be derived for other estimators, including $\hat{A}^{BR}_{j}$, $\hat{C}^{BR}_{h}$ at long horizons $h$, and $\widehat{FEVD}(h)$.
-The wild bootstrap approach of Shao (2010) can be used as a convincing numerical implementation of the delta method.
+Using the Delta method, an asymptotic normal distribution can be derived for other estimators, including $\hat{A}^{BR}_{j}$ and $\hat{C}^{BR}_{h}$ at long horizons $h$999 Another function of interest is the FEVD.
+Gorodnichenko and Lee (2020) studied inference on FEVD using local projections and stationary VAR-bootstrap under assumption of homoscedasticity of shocks.
+They show that block bootstrap of Kilian and Kim (2011) can also be used for inference on FEVD under the assumption of stationary data.
+The dependent wild bootstrap approach can also be applied to and $\widehat{FEVD}(h)$ without requiring stationarity of the VAR model.
+The dependent wild bootstrap approach of Shao (2010) can be used as a convincing numerical implementation of the delta method.
 A particular way to perform bootstrap is given in the following algorithm.
@@
   1. Step 1.
-    Select bandwidth $B_{T}$. For example, $B_{T}=0.75\sqrt[3]{T}$
+    Select bandwidth $B_{T}$ according to Theorem 5. For example, $B_{T}=0.75\sqrt[3]{T}$.
   2. Step 2.
-    Generate i.i.d. $\zeta^{s}_{i}\sim N(0,\frac{1}{B_{T}})$ for $i=-B_{T}+1,-B_{T}+2,...,0,1,...T$
+    Generate i.i.d. $\zeta^{s}_{t}\sim N(0,\frac{1}{B_{T}})$ for $t=-B_{T}+1,-B_{T}+2,...,0,1,...T$.
   3. Step 3.
-    Recursively compute $u^{s}_{t}=\zeta^{s}_{t}+\zeta^{s}_{t-1}+...+\zeta^{s}_{t-B_{T}+1}$ , $t=1,...T$
+    Recursively compute $u^{s}_{t}=\zeta^{s}_{t}+\zeta^{s}_{t-1}+...+\zeta^{s}_{t-B_{T}+1}$ , $t=1,...T$.
   4. Step 4.
@@
   5. Step 5.
-    Compute $\hat{A}^{BR,s}_{i}=A^{BR}_{i}(\hat{\theta}^{s}(H_{1}))$, $\hat{C}^{BR,s}_{h}=C^{BR}_{h}(\hat{\theta}^{s}(H_{1}))$, $\hat{b}^{s}_{\cdot 1}(\hat{\theta}^{s}(H_{1}))$ using Algorithms 1 and 2.
+    Compute $\hat{A}^{BR,s}_{i}=A^{BR}_{i}(\hat{\theta}^{s}(H_{1}))$, $\hat{C}^{BR,s}_{h}=C^{BR}_{h}(\hat{\theta}^{s}(H_{1}))$, $\hat{b}^{s}_{\cdot 1}(\hat{\theta}^{s}(H_{1}))$ for $h\geq H_{1}$ using Algorithms 1 and 2.
 This algorithm is consistent for estimating the asymptotic distribution of $\hat{\theta}$ as the following theorem shows.
 Theorem 6.
-Suppose Assumption 1-3 holds and $B_{T}^{2}/T\to 0$. Then
+Suppose Assumption 1-3 hold and $B_{T}^{2}/T\to 0$. Then
 $\lim_{T\to\infty}\sup_{x\in R,\|a\|=1}|P(\sqrt{T}a^{\prime}(\hat{\theta}^{s}(H_{1})-\hat{\theta}(H_{1}))<x|y_{1},...,y_{T})-P(\sqrt{T}a^{\prime}(\hat{\theta}(H_{1})-\theta(H_{1}))<x)|=0.$ (40)
@@
 The validity of the dependent wild bootstrap is based on the fact that it mimics the asymptotic Gaussian distribution of $\hat{\theta}(H_{1})$ with a covariance matrix that coincides with the HAC estimator given in Theorem 5.
-This argument does not impose any parametric model on $\eta_{t}$ and $Z_{t}$ and is generally applicable under the maintained weak dependence assumption (Assumption 2.2).
-Note that the algorithm uses $\hat{\Xi}_{t}$ that depend on possibly non-stationary processes with strong dependence $Y_{t}$.
+This argument does not impose any parametric model on $\eta_{t}$ and $Z_{t}$ and is generally applicable under the maintained Assumption 2.2 (weak dependence of innovations).
+Note that the algorithm uses $\hat{\Xi}_{t}$ that depend on possibly nonstationary processes with strong dependence, $Y_{t}$.
 As a result, conventional block bootstrap procedures are not directly applicable to $\hat{\Xi}_{t}$.
@@
   1. 1.
-    Use Algoritm 3 to compute $\{\hat{\theta}^{s}\}_{s=1}^{S}$.
+    Use Algorithm 3 to compute $\{\hat{\theta}^{s}\}_{s=1}^{S}$.
   2. 2.
-    Estimate standard error for each $h=1,\dots,H$
-    $\hat{\sigma}_{h}=(q(f_{h}(\hat{\theta}^{s}),\Phi(1))-q(f_{h}(\hat{\theta}^{s}),1-\Phi(1)))/2$ (41)
+    Estimate standard error for each $h=1,\dots,H_{2}$
+    $\hat{\sigma}_{h}=(q(f_{h}(\hat{\theta}^{s}),\Phi(1))-q(f_{h}(\hat{\theta}^{s}),1-\Phi(1)))/2,$ (41)
     where $q(f_{h}(\hat{\theta}^{s}),\alpha)$ is the order statistics of $\{f_{h}(\hat{\theta}^{s})\}_{s=1}^{S}$ corresponding to the quantile $\alpha$ and $\Phi(\cdot)$ is the standard normal CDF.
@@
     For each bootstrap draw $s=1,\dots,S$ compute
-    $\xi^{s}=\sup_{h=1,\dots,H}\frac{|f_{h}(\hat{\theta}^{s})-f_{h}(\hat{\theta})|}{\hat{\sigma}_{h}}.$ (42)
+    $\xi^{s}=\sup_{h=1,\dots,H_{2}}\frac{|f_{h}(\hat{\theta}^{s})-f_{h}(\hat{\theta})|}{\hat{\sigma}_{h}}.$ (42)
   4. 4.
@@
   5. 5.
-    Compute
-    $[f_{j}(\hat{\theta})-\hat{\sigma}_{h}CV_{\alpha},f_{h}(\hat{\theta})+\hat{\sigma}_{h}CV_{\alpha}].$ (43)
-In large samples, by the delta method, the boostrap distribution of $f_{h}(\hat{\theta}^{s})$ is approximately Gaussian.
-As a result, the difference between the $q(f_{h}(\hat{\theta}^{s}),\Phi(1))$ and $q(f_{h}(\hat{\theta}^{s}),1-\Phi(1))$ quantiles is converging in probablity to 2 standard deviations of $f_{h}(\hat{\theta})$ as $T,S\to\infty$.
-Step 1 can be replaced with a conventional bootstrap standard error estimator if the bootstrap distribution tails are not fat.
+    Compute for all $h$
+    $[f_{h}(\hat{\theta})-\hat{\sigma}_{h}CV_{\alpha},f_{h}(\hat{\theta})+\hat{\sigma}_{h}CV_{\alpha}].$ (43)
+In large samples, by the delta method, the bootstrap distribution of $f_{h}(\hat{\theta}^{s})$ is approximately Gaussian.
+As a result, the difference between the $q(f_{h}(\hat{\theta}^{s}),\Phi(1))$ and $q(f_{h}(\hat{\theta}^{s}),1-\Phi(1))$ quantiles is converging in probability to 2 standard deviations of $f_{h}(\hat{\theta})$ as $T,S\to\infty$.
+Step 1 can be replaced with a conventional bootstrap standard error estimator if the sample size is sufficiently large.
+[^10] Substitution of interquartile range for standard deviation is frequently used in the bootstrap literature (e.g.
+Chernozhukov et al., 2020).
+See, for example, Remark 3.2 of Chernozhukov et al.
+(2013) which uses the insights from Kato (2011).
 Another advantage of using the s.e.
-estimator based on the quantiles is that the simultaneous bounds coincide with the $68\%$ pointwise bounds if we set the critical value $CV_{\alpha}=1$.
+estimator based on the quantiles is that the simultaneous bounds coincide with the $68\%$ point-wise bounds if we set the critical value $CV_{\alpha}=1$.
 An alternative use of Algorithm 4 is a test of misspecification.
@@
 We can construct a joint confidence set for the vector of differences of impacts under these two schemes, $\hat{b}^{Cholesky}_{.1}-\hat{b}^{HetIV}_{.1}$.
-Alogrithm 4 can then be used to compute $\sup t$ critical values for a test of the hypothesis $b^{Cholesky}_{.1}-b^{HetIV}_{.1}=0$.
+Alogrithm 4 can then be used to compute $\sup t$ critical values for a test of hypothesis $b^{Cholesky}_{.1}-b^{HetIV}_{.1}=0$.
 We illustrate both uses of Algorithm 4 in Sections 4 and 5.
-3.5 Smoothed local projections using minimum distance approach.
+3.5 Smoothed local projections using minimum distance approach
 Throughout the paper, we assume that the underlying process that generates the data follows a finite order VAR($p$) model.
-This model imposes particular relations on the matrices $C_{h}$, $h=1,\dots,H$ ($C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ with $A_{\ell}=0$ for $\ell>p$) that can be used to improve efficiency.
-The backward recursion estimator $\hat{A}^{BR}_{i}$, introduced in Algorithm 1, imposes the VAR structure on $C_{h}$ with $h=1,\dots,p$ but does not use the overidentifying information about $C_{h}$ for $h>p$.
-One can improve efficincy of the LP estimates using the following miminum distance estimator.
+This model imposes particular relations on the matrices $C_{h}$, $h=1,\dots,H$ that can be used to improve efficiency (i.e., $C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ with $A_{\ell}=0$ for $\ell>p$).
+The backward recursion estimator $\hat{A}^{BR}_{i}$, introduced in Algorithm 1, imposes the VAR structure on $C_{h}$ with $h=1,\dots,p$.
+However, it does not use the overidentifying information about $C_{h}$ for $h>p$.
+One can improve efficiency of the LP estimates using the following minimum distance estimator.
 Take any positive definite matrix $\hat{W}$ that converges to a positive definite matrix $W$ of dimension $n^{2}\times H_{1}$ and minimize
 $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})=\underset{(A_{1},\dots,A_{p})\in\mathbb{R}^{pn^{2}}}{\operatorname{argmin}}(g(A_{1},\dots,A_{p})-\hat{g}^{LP})^{\prime}\hat{W}(g(A_{1},\dots,A_{p})-\hat{g}^{LP}),$ (44)
-where the vector of local projection is denoted as $\hat{g}^{LP}=vec(\hat{C}_{1},\dots,\hat{C}_{H_{1}})$ and $g(A_{1},\dots,A_{p})=vec(C_{1}(A_{1},\dots,A_{p}),\dots,C_{H_{1}}(A_{1},\dots,A_{p}))$ is the corresponding recursively computed IRF coefficients for any values of matrices $(A_{1},\dots,A_{p})$.
-From the theory of the minimum distance estimators (see, for example, Hansen (2022)) the optimal choice of $\hat{W}$ correspond to the inverse of the long run covariance matrix, given in Theorem 4 earlier in this paper.
-In small samples, however, the HAC estimator given in Theorem 5 may result in non-invertible matrix that needs to be regularized.
-In this paper, we instead use a diagonal matrix $\hat{W}$ with vector of inverse variances of the components of $\hat{g}^{LP}$ on the diagonal.
-These variances are estimated using dependent wild bootstrap.
-Based on the estimates of minimum distance $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$, we can compute smoothed impulse responses for any horizon $H$, $(\hat{C}^{MD}_{1},\dots,\hat{C}^{MD}_{H})\triangleq C_{1}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),\dots,C_{H}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$, using the usual recursive formulas.
+where the vector of the local projection estimators is denoted as $\hat{g}^{LP}=vec(\hat{C}_{1},\dots,\hat{C}_{H_{1}})$ and $g(A_{1},\dots,A_{p})=vec(C_{1}(A_{1},\dots,A_{p}),\dots,C_{H_{1}}(A_{1},\dots,A_{p}))$ is the corresponding recursively computed IRF coefficients for any values of matrices $(A_{1},\dots,A_{p})$.
+From the theory of the minimum distance estimators (see, for example, Hansen (2022)) the optimal choice of $\hat{W}$ corresponds to the inverse of the long-run covariance matrix, given in Theorem 4 earlier in this paper.
+In small samples, however, the HAC estimator given in Theorem 5 may result in a non-invertible matrix that needs to be regularized.
+In this paper, we instead use a diagonal matrix $\hat{W}$ with the vector of inverse variances of the components of $\hat{g}^{LP}$ on the diagonal.
+These variances are estimated using the dependent wild bootstrap.
+Based on the estimates of minimum distance $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$, we can compute smoothed impulse responses for any horizon $H$,
+$(\hat{C}^{MD}_{1},\dots,\hat{C}^{MD}_{H})\triangleq C_{1}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),\dots,C_{H}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),$
+using the usual recursive formulas.
 Under the correct VAR($p$) specification, the limit of (44) coincides with the true parameters $A_{1},\dots,A_{p}$.
-Moreover, the estimator $((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$ is a smooth transformation of $\hat{\theta}$.
+Moreover, the estimator $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$ is a smooth transformation of $\hat{\theta}$.
 So, by the delta method, the dependent wild bootstrap from Theorem 6 remains valid for $\hat{C}^{MD}_{i}$.
@@
 In the presence of conditional heteroscedasticity, the smoothed LP estimators may be more efficient than the conventional recursive VAR estimators of $C_{h}$.
-The idea of smoothing local projection is not new.
+The idea of smoothing the local projection is not new.
 Barnichon and Brownlees (2019) imposed a particular flexible form on the LP estimates within a ridge-regression framework without providing the corresponding inference procedure.
-In contrast to the earlier studies, we use the VAR structure of the IRF.
+In contrast to earlier studies, we use the VAR structure of the IRF for smoothing.
 This allows us to improve efficiency under correct specification and use the dependent wild bootstrap as a robust inference procedure.
 4 Simulation experiment based on a DSGE model of US economy
-In this section, we would like to illustrate how hetersocedasticity IV identification works together with the dependent wild bootstrap using data created by a dynamic general equilibrium model.
-For the data generation process, we will use the well-known model Smets and Wouters (2007) (SW later on), which is a medium-scale New Keynesian model with a wide range of nominal and real rigidities.
+In this section, we would like to illustrate how heteroscedasticity IV identification works with the dependent wild bootstrap using data created by a dynamic general equilibrium (DSGE) model.
+For the data generation process, we will use the well-known model of Smets and Wouters (2007) (SW later on), a medium-scale New Keynesian model with a wide range of nominal and real rigidities.
 In particular, the model introduces nominal price and wage rigidity, consumption habits, investment adjustment, and capital utilization costs.
-At the same time, the Taylor rule for the interest rate assumes that the monetary authority reacts to the deviation of output from the level that would prevail in an economy with flexible prices and wages, which further increases the dimension of the problem, since this requires additionally describing the dynamics of a hypothetical economy with flexible prices.
+At the same time, the Taylor rule for the interest rate assumes that the monetary authority reacts to the deviation of output from the level that would prevail in an economy with flexible prices and wages, which further increases the dimension of the problem because this requires additionally describing the dynamics of a hypothetical economy with flexible prices.
 The state vector in the model has a dimension of about 20 variables.
-As in the original paper, we introduced 7 structural macroeconomic shocks into the model, including the monetary policy shock.
-Typically, the model is estimated using the Kalman filter with the observation vector consisting of 7 variables (the number of variables cannot be less than the number of shocks): the log difference of real GDP, real consumption, real investment and the real wage, log hours worked, the log difference of the GDP deflator, and the federal funds rate.
+As in the original paper, we introduce seven structural macroeconomic shocks into the model, including the monetary policy shock.
+Typically, the model is estimated using the Kalman filter with the observation vector consisting of seven variables (the number of variables cannot be less than the number of shocks): the log difference of real GDP, real consumption, real investment, and the real wage, log hours worked, the log difference of the GDP deflator, and the federal funds rate.
 The filter is used to estimate the linear state space model that corresponds to the solution of the log-linearized DSGE models.
-The unobserved state vector follows an SVAR model.
-Since the observation vector has a smaller dimension than the state vector (7 vs 20), the observation vector is not generally a finite-order vector autoregression.
-However, in a fairly general case, the vector of observed variables can be represented as a VARMA process, and, assuming invertibility of the MA part, a VAR of finite order can act as an approximation for the corresponding VAR($\infty$).
+The unobserved state vector follows a SVAR model.
+Since the observation vector has a smaller dimension than the state vector (7 vs.
+20), the observation vector is not generally a finite-order vector autoregression.
+However, in a fairly general case, the vector of observed variables can be represented as a VARMA process, and, assuming the invertibility of the MA part, a VAR of finite order can act as an approximation for the corresponding VAR($\infty$).
 [^11] The quality of approximation by a finite-order VAR depends on the specific structure of the model, the macroeconomics variables used as observables, and on the persistence of shocks (Ravenna, 2007; Fernandez-Villaverde et al., 2007).
-In the experiment of this section, we generate samples of various lengths from the log-linearized SW DSGE model (equivalent to a VARMA model) calibrated using the posterior mean estimates obtained in the (Smets and Wouters, 2007).
-The only exception was the standard deviation parameter of the monetary policy shock, for which we introduced two regimes described by a Markov chain.
-We consider a symmetric case where the probability of remaining in one of the regimes is 70%, and the switch to another regime is 30%.
-Then we introduce an auxiliary instrumental variable $Z_{t}$, which takes the value 0 in the first regime and 1 in the second.
-The value of the standard deviation of the monetary policy shock, $\sigma_{r,t}$, was set according to the following equation:
+In the experiment of this section, we generate samples of various lengths from the log-linearized SW DSGE model (equivalent to a VARMA model) calibrated using the posterior mean estimates obtained in (Smets and Wouters, 2007).
+The only exception was the standard deviation parameter of the monetary policy shock, for which we introduce two regimes described by a Markov chain.
+We consider a symmetric case where the probability of remaining in one of the regimes is 70%, and the probability of switching to the other regime is 30%.
+We introduce an auxiliary instrumental variable $Z_{t}$, which takes the value 0 in the first regime and 1 in the second.
+The value of the standard deviation of the monetary policy shock, $\sigma_{r,t}$, is set according to the following equation:
 $\sigma_{r,t}=\sigma_{r,0}(1+3Z_{t})$ (45)
-where $\sigma_{r,0}$ takes a value of posterior mean estimate for the standard deviation of the monetary policy shock in the Smets and Wouters (2007) paper.
-This specification also assumes that the standard deviation of the monetary shock increases by a factor of 4 in the second regime.
-Figure 1: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 160 periods.LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line)
-Figure 2: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 1600 periods.
+where $\sigma_{r,0}$ is set to a value of the posterior mean estimate for the standard deviation of the monetary policy shock in the Smets and Wouters (2007) paper.
+This specification assumes that the standard deviation of the monetary shock increases by a factor of four in the second regime.
+Figure 1: Theoretical versus empirical impulse response functions to a monetary policy shock identified through heteroscedasticity-IV, sample length is 160 periods.
 LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
-Next, using the simulated dataset of seven observable variables and the instrument $Z_{t}$ we apply our new identification method to extract the structural monetary policy shock and compute the corresponding IRFs.
-We included in the vector of observed variables in local projection model the following variables: log-deviations from the steady state level for real GDP, household consumption, investment, labor, wage, GDP deflator growth rate, and the gross nominal interest rate.
-To emphasize that the model can handle both stationary and nonstationary variables, we used GDP deflator index rather than growth rates in the econometric exercise.
-Thus, the model will use 6 stationary variables and 1 non-stationary variable.
+Figure 2: Theoretical versus empirical impulse response functions to a monetary policy shock identified through the heteroscedasticity-IV, sample length is 1600 periods.
+LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
+Next, using the simulated dataset of seven observable variables and the instrument $Z_{t}$, we apply our new identification method to extract the structural monetary policy shock and compute the corresponding IRFs.
+We include in the vector of observed variables in the local projection model the following variables: log deviations from the steady-state level for real GDP, household consumption, investment, labor, wage, GDP deflator index, and the gross nominal interest rate.121212 We use the GDP deflator index rather than growth rates in the econometric exercise in order to emphasize that the method can handle both stationary and nonstationary variables.
+Thus, the model will use six stationary variables and one nonstationary variable.
 The original calibration of the model is based on the quarterly frequency of observation.
-The model includes 4 lags, which corresponds to a lag of 1 year.
-$Z_{t}$ is used as an instrument for heteroscedasticity of monetary policy shock.
-Finally, we compare the IRFs generated from the SW model (VARMA, theoretical) to those obtained from the estimated VAR model to validate our method’s accuracy and effectiveness in capturing the dynamics prescribed by the SW framework.
+The model includes four lags, which corresponds to a lag of one year.
+$Z_{t}$ is used as an instrument for the heteroscedasticity of the monetary policy shock.
+Finally, we compare the IRFs generated from the theoretical SW model (VARMA) to those obtained from the estimated VAR model to validate our method’s accuracy and effectiveness in capturing the dynamics prescribed by the SW framework.
 Figure 1 shows the estimation results for 160 periods (40 years), and Figure 2 shows the results for 1600 periods.
 In Figure 1, it can be seen that the method can reconstruct the IRF for monetary policy shocks within the approximate VAR(4) model using a typical sample size.
-The true IRF are well within the simultaneous 68% confidence bounds.
-As we increase the sample size to 1600 periods (Figure 2), the LP estimates are getting even closer to the true values.
-This shows that LP estimators based on finite VAR(4) can reasonably well approximate the underlying VARMA model.
-Moreover, the simultaneous confidence bounds successfully covered the true IRF in all the figures, including the non-stationary GDP deflator index.
+The true IRFs are well within the simultaneous 68% confidence bounds.
+As we increase the sample size to 1600 periods (Figure 2), the LP estimates get even closer to the true values.
+This result shows that LP estimators based on finite VAR(4) can reasonably well approximate the underlying VARMA model.
+Moreover, the simultaneous confidence bounds successfully covered the true IRFs in all the figures, including the non-stationary GDP deflator index.
 5 Empirical application: identificaiton of monetary policy shock in US
-Identification of a monetary policy shock is a classical application of structural VAR models that goes back to Sims (1980).
-This application has been extensively studied with many identification schemes being proposed, including the classical Cholesky decomposition (Sims, 1980; Christiano et al., 1999), instrumental variables (for example, Romer and Romer, 1989, 2004; Gertler and Karadi, 2015), sign restrictions (for example, Uhlig, 2005; Gafarov et al., 2018; Antolín-Díaz and Rubio-Ramírez, 2018), high-frequency identification (for example, Faust et al., 2004), and regime-switching models (for example, Sims and Zha, 2006; Lütkepohl and Netšunajev, 2017) to name a few.
-[^9] For foundational context and a detailed examination of prior empirical efforts we refer reader to seminal works by Christiano et al.
+The identification of monetary policy shocks is a classic application of structural VAR models that goes back to Sims (1980).
+This application has been extensively studied with many identification schemes being proposed, including the classical Cholesky decomposition (Sims, 1980; Christiano et al., 1999), instrumental variables (for example, Romer and Romer, 1989, 2004; Gertler and Karadi, 2015), sign restrictions (for example, Uhlig, 2005; Gafarov et al., 2018; Antolín-Díaz and Rubio-Ramirez, 2018), high-frequency identification (for example, Faust et al., 2004), and regime-switching models (for example, Sims and Zha, 2006; Lütkepohl and Netšunajev, 2017), to name a few.
+[^13] For foundational context and a detailed examination of prior empirical efforts we refer the reader to seminal works by Christiano et al.
 (1999), Boivin et al.
@@
 5.1 Data and model
-To compare our identification strategy to those commonly used in the literature, we focus on the reduced-form model studied in seminal papers including, but not limited to Christiano et al.
-(1999), Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018).
-Our specification includes six variables for the U.S.: real GDP, GDP deflator, commodity price index, total reserves, nonborrowed reserves, and federal funds rate, described in Table 1.
-The sample covers the period from January 1965 to November 2007, with monthly data similar to Antolín-Díaz and Rubio-Ramírez (2018).
-[^10] The monthly series for real GDP and the GDP deflator are interpolated from quarterly data using Chow and Lin (1971) methodology.
-We use the industrial production index for real GDP interpolation and the consumer price index for the GDP deflator.
+To compare our identification strategy with those commonly used in the literature, we focus on the reduced-form model studied in seminal papers including, but not limited to Christiano et al.
+(1999), Uhlig (2005), Antolín-Díaz and Rubio-Ramirez (2018).
+Our specification includes six variables for the U.S.: real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and federal funds rate, as described in Table 1.
+The sample covers the period from January 1965 to November 2007, with monthly data similar to Antolín-Díaz and Rubio-Ramirez (2018).
+[^14] The monthly series for real GDP and the GDP deflator are interpolated from quarterly data using the Chow and Lin (1971) methodology.
+We use the industrial production index for the real GDP interpolation and the consumer price index for the GDP deflator.
 Our specification includes a constant and a polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$.
-[^11] When we compare results to those from the literature, we mimic number of lags and trends.
+[^15] When comparing our results to those from the literature, we align the number of lags and the inclusion of trends to match their specifications.
 For example, Christiano et al.
-(1999) uses constant but no time trend, number of lags 6 and 12.
-Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018) use 12 lags and no constant or deterministic trends.
+(1999) use constant but without time trend, with lag lengths of 6 and 12, while Uhlig (2005), Antolín-Díaz and Rubio-Ramirez (2018) use 12 lags without a constant or deterministic trend.
 Table 1: Description of Variables and Data Sources
-Variable Source Seasonally Mnemonics/Link Adjusted Real GDP FRED Yes https://fred.stlouisfed.org/series/GDPC1 GDP Deflator FRED Yes https://fred.stlouisfed.org/series/GDPDEF) Total Reserves FRED Yes https://fred.stlouisfed.org/series/TRARR Nonborrowed Reserves FRED Yes https://fred.stlouisfed.org/series/BOGNONBR Fed Funds Rate FRED No https://fred.stlouisfed.org/series/FEDFUNDS Commodity Price Index GFD No Global Financial Data Consumer Price Index FRED Yes https://fred.stlouisfed.org/series/CPIAUCSL Industrial Production Index FRED Yes https://fred.stlouisfed.org/series/INDPRO
+Variable Source Seasonally Mnemonics Adjusted Real GDP FRED Yes GDPC1 GDP Deflator FRED Yes GDPDEF) Total Reserves FRED Yes TRARR Nonborrowed Reserves FRED Yes BOGNONBR Federal Funds Rate FRED No FEDFUNDS Commodity Price Index GFD No Global Financial Data Consumer Price Index FRED Yes CPIAUCSL Industrial Production Index FRED Yes INDPRO
+Notes: Time series are available on the following website, https://fred.stlouisfed.org/series/.
 5.2 FOMC meetings and heteroscedasticity of monetary policy shocks
 It is well documented in the high-frequency identification literature that announcements made after FOMC meetings immediately impact financial variables.
-This change in financial indicators is associated with unforeseen changes in monetary policy, also referred to as monetary policy shocks.121212 See, for example, Cochrane and Piazzesi (2002), Faust et al.
+These changes in financial indicators are associated with unforeseen changes in monetary policy, commonly referred to as monetary policy shocks.161616 See, for example, Cochrane and Piazzesi (2002), Faust et al.
 (2004), Gürkaynak et al.
-(2005), Gertler and Karadi (2015), Ramey (2016), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022)
-We focus exclusively on the dates of FOMC meeting announcements including meetings, telephone and video conferences, unscheduled meetings, and sequential day meetings.
-Our baseline instrumental variable indicates the number of meetings occurring in a given month, including telephone conferences and unscheduled meetings.
-[^13] An alternative instrumental variable is a dummy with one’s in the months when the meetings occur and zero otherwise.
+(2005), Gertler and Karadi (2015), Ramey (2016), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022).
+We focus exclusively on the dates of FOMC meeting announcements, including meetings, telephone and video conferences, unscheduled meetings, and sequential day meetings.
+Our baseline instrumental variable indicates the number of meetings that occur in a given month, including telephone conferences and unscheduled meetings.
+[^17] An alternative instrumental variable is a dummy with ones in the months when the meetings occur and zero otherwise.
 Results are available upon request.
@@
 Louis Fed website https://fraser.stlouisfed.org/title/federal-open-market-committee-meeting-minutes-transcripts-documents-677?browse=1980s Figure 3 plots the monthly time series of our baseline instrument starting from 1965 till 2007.
-As one can see the number of the meetings in a given month varies from zero to four.
-Figure 3: FOMC meetings
-Notes: The blue bars indicate the number of FOMC meetings per month, including telephone conferences and unscheduled meetings
+As can be seen, the number of meetings in a given month varies from zero to four.
+Figure 3: Number of FOMC meetings per month from 1965 to 2007.
+Notes: The blue bars indicate the number of FOMC meetings per month, including telephone conferences and unscheduled meetings.
 Most of the papers that use FOMC meetings for identification purposes focus their analysis on the sample starting from 1994 onwards.
-This is done due to the fact the FED releases no public announcements about monetary policy decisions prior to 1994.
-Some researches, like, Jayawickrema and Swanson (2021) extend the sample backward to 1988.
-As our identification assumption remains agnostic about the direction of the change, size, we can use the sample back to 1965.
+This is because the FED released no public announcements about monetary policy decisions prior to 1994.
+Some scholars (for example, Swanson and Jayawickrema, 2023) extended the sample back to 1988.
+As our identification assumption is agnostic about the direction and size of the change, we can extend our sample back to 1965.
 5.3 Results
@@
 Figure 4 displays the impulse response functions of six endogenous variables to the structural monetary policy shock.
-We estimate those responses using Heteroskedssticity-IV LP estimator in our benchmark specification with $p=12$ without a trend.
-The dotted lines represent 68% point-wise confidence intervals, while the dot-dashed lines are simultaneous sup-t bands.
-Although simultaneous confidence bands are wider relative to point-wise, they are extremely useful, when one, for example, is interested in comparing short-run versus medium-run versus long-run effects.
-[^15] See section 5.3.4 below for comparison with an alternative simultaneous bounds based on Bonferroni inequality.
-Recent contribution regarding simultaneous confidence intervals are Jordà (2009), Inoue and Kilian (2016), Montiel Olea and Plagborg-Møller (2019), Inoue and Kilian (2022),Arias et al.
+We estimate those responses using the heteroscedasticity-IV LP estimator in our benchmark specification with $p=12$ and $k=0$ (without a trend).
+The dotted lines represent 68% point-wise confidence intervals, while the dashed lines are simultaneous sup-t bands.
+Although simultaneous confidence bands are wider relative to point-wise intervals, they are useful when, for example, one is interested in comparing short-run, medium-run, and long-run effects.
+[^19] See section 5.3.4 below for comparison with alternative simultaneous bounds based on the Bonferroni inequality.
+Recent contributions on simultaneous confidence intervals are Jordà (2009), Inoue and Kilian (2016), Montiel Olea and Plagborg-Møller (2019), Inoue and Kilian (2022), Arias et al.
 (2023).
-Figure 4: Identification of monetary policy shock in the US.
+Before we proceed to the estimation, it is worth noting that the hypothesis $\gamma_{1}=0$ is rejected in favor of the alternative $\gamma_{1}>0$ (Assumption 1) at 5% significance level.
+The corresponding test statistic is $t=1.77$, with a p-value equal to $0.039$ based on Algorithm 3.
+In other words, the variance of the Fed funds rate forecast errors is higher in the months with more FOMC meetings.
+This confirms earlier findings in the high-frequency identification literature that markets react to FOMC meeting announcements (see, for example, Faust et al., 2004).
+Figure 4: Identification of monetary policy shock in the U.S.
 IRF estimates with point-wise and simultaneous bands.
-Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
-As observed, monetary policy tightening leads to a significant drop in output, which remains significant from 10 to 60 months using point-wise confidence bands.
+The heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dash line).
+As evident from Figure 4, monetary policy tightening leads to a significant drop in output, which remains significant from 10 to 60 months using point-wise confidence bands.
 When considering joint significance, this decline is significant from approximately 15 to 35 months, a duration notably longer than previously reported results by Inoue and Kilian (2016).
-The results on inflation are similar to Wald-joint confidence intervals considered in that paper under a stationary VAR specification.
+The results on inflation are similar to those obtained using Wald-joint confidence intervals considered in that paper under a stationary VAR specification.
 In both cases, the price puzzle disappears once we consider simultaneous confidence intervals.
-It is important to keep in mind that our procedure of simultaneous confidence intervals construction allows for unit roots without pre-tests, while Inoue and Kilian (2016) requires transforming the model into a stationary one which may cause a pretesting bias.
-5.3.2 Sensitivity to number of lags and trends
+It is important to keep in mind that our procedure for constructing simultaneous confidence intervals allows for unit roots without pre-testing, while Inoue and Kilian (2016) requires transforming the model into a stationary one, which may cause a pre-testing bias.
+5.3.2 Sensitivity to the number of lags and polynomial trends
 Our inference methodology allows for a flexible trend specification.
-Appendix Figures 7-9 compare the benchmark model results (blue lines) with a more flexible models (golden lines): $(k=4,p=12)$, $(k=0,p=24)$, and $(k=4,p=24)$.
+Appendix Figures 7-9 compare the benchmark model results (blue lines) with more flexible models (golden lines): $(k=4,p=12)$, $(k=0,p=24)$, and $(k=4,p=24)$.
 The polynomial trend can account for slow and gradual changes in the mean values of stationary variables and the growth rates of unit root variables.
-Either including a trend or additional lags (or both) results in shorter-lived responses of the real GDP and weaker effects of the monetary shock on inflation and commodity prices.
-These results show that long-run effects of the monetary policy are sensitive to model specification.
-One possible explanation is that the trend (or additional lags) absorbs the long-run variation in the real GDP which is otherwise be attributed to monetary policy shocks.
+Including either a trend or additional lags (or both) results in shorter-lived responses of real GDP and weaker effects of the monetary shock on inflation and commodity prices.
+These results show that the long-run effects of monetary policy are sensitive to model specification.
+One possible explanation is that the trend (or additional lags) absorbs the long-run variation in real GDP that is otherwise attributed to monetary policy shocks.
 As a result, the more flexible specifications only detect the impact of monetary policy shocks that are neutral in the long run.
-This result is consistent with the panel regression findings at 10 year horizons in Jordà et al.
-(2020): the monetary tightening shocks can be non-neutral in the long run, while monetary loosening shocks are typically short lived.
-In the current literature, it is a convention to exclude the time trend from the monetary SVARs computed in levels (see, for example, Uhlig, 2005; Antolín-Díaz and Rubio-Ramírez, 2018).
+This result is consistent with the panel regression findings at 10-year horizons in Jordà et al.
+(2020): monetary tightening shocks can be non-neutral in the long run, while monetary loosening shocks are typically short-lived.
+In the current literature, it is conventional to exclude the time trend from monetary SVARs computed in levels (see, for example, Uhlig, 2005; Antolín-Díaz and Rubio-Ramirez, 2018).
 To facilitate the comparison, we will report only the results for the benchmark specification $(k=0,p=12)$ in the subsequent sections.
 5.3.3 Smoothed LP
-Appendix Figure 10 presents similar impulse response functions with smoothed point-wise and simultaneous bands for the benchmark specification.
-Compared to the raw LP estimates, Figure 4, the smoothed IRF results in a stronger impact of the shock on GDP at the long horizons, but a weaker impact on inflation in the long horizons.
-Notice that the confidence bounds got tighter as a result of the smoothing.
-This tightening occurs as a result of efficiency gains from imposing restrictions on the LP coefficient under the assumption of a correct specification of the VAR(12) model.
+Appendix Figure 10 presents impulse response functions with smoothed point-wise and simultaneous bands for the benchmark specification.
+Compared to the raw LP estimates in Figure 4, the smoothed IRF results in a stronger impact of the shock on GDP at longer horizons but a weaker impact on inflation at longer horizons.
+Notice that the confidence bounds become tighter as a result of the smoothing.
+This tightening occurs as a result of efficiency gains from imposing restrictions on the LP coefficients under the assumption of a correct specification of the VAR(12) model.
 The smoothed LP estimators are based on a minimum distance estimator that imposes the VAR(12) model on the coefficients $C_{h}$.
-The distance between the VAR(12) model and the raw estimates was weighed with the inverse of the asymptotic variance of the individual LP estimators.
-Such weights are not fully optimal and could be further improved by using the inverse variance-covariance matrix of LP estimators as in Theorem 5.
+The distance between the VAR(12) model and the raw estimates was weighted with the inverse of the asymptotic variance of the individual LP estimators.
+Such weights are not fully optimal and could be further improved by using the inverse variance-covariance matrix of LP estimators, as in Theorem 5.
 In practice, however, the optimal weights require some regularization in small samples.
@@
 Appendix Figure 11 shows IRFs with sup-t versus union bands.
-Similar to Montiel Olea and Plagborg-Møller (2019) we find that Bonferroni confidence bands are noticeably wider relative to sup-t confidence bands.
-5.3.5 Comparison with Cholesky
-Figure 5 compares the IRFs to a monetary policy shock identified Heteroscedasticity-IV LP to IRFs from Christiano et al.
+Similar to Montiel Olea and Plagborg-Møller (2019), we find that Bonferroni confidence bands are noticeably wider relative to sup-t confidence bands.
+5.3.5 Comparison with Cholesky-type identification
+Figure 5 compares the IRFs to a monetary policy shock identified using heteroscedasticity-IV LP to IRFs from Christiano et al.
 (1999).
@@
 As can be seen, the results of the new identification approach are very close to those of Christiano et al.
-(1999), which is a standard Cholesky identification scheme with a priori zeros impacts on real GDP, GDP deflator, and commodity price index.
-Specifically, the $\sup t$ test for the equality of the impact vectors for the two identification schemes given in Subsection 3.4 has $\sup t=1.2637$ with p-value $0.53$ (critical value at 32% is 1.8083).
-Without imposing restrictions, we obtain a similar conclusion.
-Notice that the confidence bounds for Heteroscedasticity-IV estimators account for uncertainty in the structural rotation matrix (which depends on both $\hat{\gamma}$ and $\hat{\Sigma}_{\eta}$), while the Cholesky approach only accounts for uncertainty in $\hat{\Sigma}_{\eta}$.
-As a result, the Cholesky IRF has misleadingly tighter confidence bounds that are only correct under the a priori zero restrictions on real GDP, GDP deflator and commodity prices implied by the identificaiton scheme.
+(1999), which is a standard Cholesky identification scheme with a priori zero impacts on real GDP, GDP deflator, and commodity price index.
+Specifically, the $\sup t$ test for the equality of the impact vectors for the two identification schemes given in Subsection 3.4 has $\sup t=1.2637$ with a p-value $0.53$ (critical value at 32% level is 1.8083).
+Without imposing the apriori zero restrictions, we obtain a similar conclusion.
+Notice that the confidence bounds for the heteroscedasticity-IV estimators account for uncertainty in the structural rotation matrix (which depends on both $\hat{\gamma}$ and $\hat{\Sigma}_{\eta}$), while the Cholesky approach only accounts for uncertainty in $\hat{\Sigma}_{\eta}$.
+As a result, the Cholesky IRF has misleadingly tighter confidence bounds that are only correct under the a priori zero restrictions on real GDP, GDP deflator, and commodity prices implied by the identification scheme.
 The classical paper Christiano et al.
-(1999) used recursive formulas for the computation of IRF.
-Appendix Figure 12 compares the results of smoothed LP Cholesky with the conventional VAR estimations under homoscedastic residuals.
-One can see that smoothed LP point estimates are only similar to the conventional VAR estimates at the short horizons.
+(1999) uses recursive formulas for the computation of IRF.
+Appendix Figure 12 compares the results of the smoothed LP Cholesky method with conventional VAR estimations under homoscedastic residuals.
+One can see that smoothed LP point estimates are similar to the conventional VAR estimates only at the short horizons.
 At the long horizons, the plots for real GDP, GDP deflator, and the commodity price index are substantially different.
-What is more surprising is that the smoothed LP point-wise confidence bounds are comparable and even tighter than the conventional VAR bootstrap bounds.
-It highlights the fact that the additional robustness of the proposed procedure does not automatically imply wider confidence bounds than its conventional alternative.
-Figure 5: Identification of monetary policy shock in the US.
-Comparison with Christiano et al.
+What is more surprising is that the smoothed LP point-wise confidence bounds are comparable to, and even tighter than the conventional VAR bootstrap bounds.
+This highlights the fact that the additional robustness of the proposed procedure does not automatically imply wider confidence bounds than its conventional alternative.
+Figure 5: Identification of the monetary policy shock in the U.S.: Comparison with Christiano et al.
 (1999).
-Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and Cholesky-LP estimator (golden dot-dash) with corresponding 68% point-wise bands (blue dots and golden dashes).
+Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and Cholesky-LP estimator (golden dashed line) with corresponding 68% point-wise bands (blue dots and golden dashes).
 5.3.6 Comparison with narrative sign restrictions
-In the previous subsection, we found out that the zero restrictions on impacts on real GDP, GDP deflator, and the commodity price index cannot be rejected simultaneously.
-However, these restrictions were questioned by the literature.
-Uhlig (2005) replaced these zero restrictions with sign restrictions on impulse responses on short horizons.
-Such an approach by construction eliminates the price puzzle.
-In a more recent study by Antolín-Díaz and Rubio-Ramírez (2018) the authors further imposed restrictions on signs of monetary policy shock on particular dates to tighten the bounds.
+In the previous subsection, we found that the zero restrictions on impacts on real GDP, GDP deflator, and the commodity price index cannot be rejected simultaneously.
+However, these restrictions have been questioned in the literature.
+In particular, Uhlig (2005) replaced these zero restrictions with sign restrictions on impulse responses at short horizons.
+Such an approach eliminates the price puzzle by construction.
+In a more recent study, Antolín-Díaz and Rubio-Ramirez (2018) further imposed additional restrictions on the signs of the monetary policy shock at particular dates to tighten the bounds on IRF.
 Both papers use Bayesian Gaussian homoscedastic SVAR models with uninformative priors over the structural rotation matrix for inference.
@@
 This approach results in wider confidence bounds since it is designed to cover the worst-case values of the unidentified structural rotations as opposed to the prior distribution weighted average.
-Figure 6: Identification of monetary policy shock in the US.
-Comparison with narrative sign restrictions (Antolín-Díaz and Rubio-Ramírez, 2018).
-Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and narrative sign restricted BVAR estimator (golden dot-dash) with corresponding 68% point-wise confidence and credibility bands (blue dots and golden dashes)
+Figure 6: Identification of monetary policy shock in the U.S.
+Comparison with narrative sign restrictions (Antolín-Díaz and Rubio-Ramirez, 2018).
+Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and narrative sign restricted BVAR estimator (golden dash) with corresponding 68% point-wise confidence and credibility bands (blue dots and golden dashes).
 […]
-There is an interesting discussion in the literature and some researchers argue that there is no price puzzle and this is a reality.
-For example, Mishkin (2007) argues that an increase in short-term interest rate increases the cost of building new houses and reduces housing activity.
-Through the supply cost channel that may cause an increase in prices in the short run.
-This example is from the construction sector, although similar discussions are shared in the agricultural sector.
+There is an interesting discussion in the literature where some researchers argue that the price puzzle may reflects the reality.
+For example, Mishkin (2007) argues that an increase in the short-term interest rate raises the cost of building new houses and reduces housing activity.
+Through the supply cost channel this may cause an increase in prices in the short run.
+This example pertains to the construction sector.
 With monetary policy tightening, the cost of mortgages increases, which impacts production costs in the affected sectors.
 6 Conclusion
-This paper makes several contributions to the literature on SVARs with potential unit roots and co-integration.
-First, we propose a novel heteroscedasticity IV for monetary policy shocks that requires weaker assumtions than the conventional external IV approaches.
-As a result, we are able to construct much longer time series of IV.
-Second, we show how one can perform point-wise and simultaneous inference on IRF with such extenral IV using LP.
-We specifically generalize the HAC estimator of LP account for the non-statioanry regressors and show equivalence of dependent wild bootstrap with the analytical HAC inference.
-Third, we use the joint draws of the LP to construct minimum distance smoothed LP IRF estimator and the corresponding confidence bounds.
-We then illustrate the usefullness of these procedures in two applications: a simulated data based on a realistic DSGE model and a monetary SVAR for the US.
-We find that the heterscedasticiy IV based on the number of FOMC meetings in a given month results in IRF that are very close to the classical Christiano et al.
-(1999) bounds and are different from a comparable BVAR with (narrative) sign restrictions that rule out price puzzle.
+This paper makes several contributions to the literature on SVARs with potential unit roots and cointegration.
+First, we propose a novel heteroscedasticity-baesd IV for monetary policy shocks that requires weaker assumptions than the conventional external IV approaches.
+As a result, we are able to construct much longer time series of instruments.
+Second, we show how one can perform point-wise and simultaneous inference on IRF with such heteroscedasticity-based external IV using the LP estimators.
+We specifically generalize the HAC estimator of LP account for the non-statioanry regressors and show equivalence of dependent wild bootstrap and the analytical HAC inference.
+Third, we use the joint draws of the LP estimates to construct minimum distance smoothed LP estimator and the corresponding confidence bounds.
+We then illustrate the usefulness of these procedures in two applications: a simulated data based on a realistic DSGE model and a monetary SVAR for the U.S.
+We find that the heteroscedasticity IV based on the number of FOMC meetings in a given month results in IRF that are very close to the classical Christiano et al.
+(1999) bounds and are different from those obtained using a comparable BVAR with narrative sign restrictions that rule out the price puzzle.
 In particular, our results suggest a much weaker long-run impact on inflation than the BVAR would imply.
@@
 A.1 Sensitivity to model specification
-Figure 7: Identification of monetary policy shock in the US.
+Figure 7: Identification of the monetary policy shock in the U.S.
 IRF estimates with a flexible trend ($k=4$ and $p=12$).
-Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=12$ (golden dot-dash line) with corresponding 68% point-wise bands
-Figure 8: Identification of monetary policy shock in the US.
-IRF estimates with a additional lags ($k=0$ and $p=24$).
-Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=0,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands
-Figure 9: Identification of monetary policy shock in the US.
-IRF estimates with both flexible trend and additional lags ($k=4$ and $p=24$).
-Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands
+Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=12$ (golden dashed line) with corresponding 68% point-wise bands.
+Figure 8: Identification of the monetary policy shock in the U.S.
+IRF estimates with additional lags ($k=0$ and $p=24$).
+Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=0,p=24$ (golden dashed line) with corresponding 68% point-wise bands.
+Figure 9: Identification of the monetary policy shock in the U.S.
+IRF estimates with both a flexible trend and additional lags ($k=4$ and $p=24$).
+Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=24$ (golden dashed line) with corresponding 68% point-wise bands.
 A.2 Smoothed LP
-Figure 10: Identification of monetary policy shock in the US.
+Figure 10: Identification of the monetary policy shock in the U.S.
 Smoothed IRF estimates with point-wise and simultaneous bands.
-Heteroscedasticity-IV smoothed LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
+Heteroscedasticity-IV smoothed LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dash line).
 A.3 Bonferroni bounds
-Figure 11: Identification of monetary policy shock in the US.
+Figure 11: Identification of the monetary policy shock in the U.S.
 Sup-t vs Bonferroni simultaneous bands.
-Heteroscedasticity-IV LP with $k=0,p=12$ estimator (blue solid line), Bonferroni bounds (gold dots) and simultaneous sup-t bounds (dot-dash line)
+Heteroscedasticity-IV LP with $k=0,p=12$ estimator (blue solid line), Bonferroni bounds (gold dots) and simultaneous sup-t bounds (dashed line).
 A.4 Comparison with conventional VAR approch
-Figure 12: Identification of monetary policy shock in the US.
+Figure 12: Identification of the monetary policy shock in the U.S.
 LP vs VAR using Christiano et al.
 (1999).
-Cholesky identification with LP estimator with $k=0,p=12$ (blue line) and homoscedastic VAR estimator (violet solid) with corresponding 68% point-wise bands ( gold dashes and violet dots)
+Cholesky identification with LP estimator with $k=0,p=12$ (blue line) and homoscedastic VAR estimator (violet solid) with corresponding 68% point-wise bands ( gold dashes and violet dots).
 Appendix B Proofs of the main results
@@
 In particular, it means that the first column of $B$, $b_{\cdot 1}$, satisfies
-$b_{\cdot 1}^{\prime}\Sigma^{-1}_{\eta}b_{\cdot 1}=1$ (48)
-Step 2. The reduced form shocks have the following representation for any $k=1,\dots,n$
-$\eta_{1t}\eta_{kt}$ $=b_{k1}b_{11}\varepsilon^{2}_{1t}+\sum_{i\neq j}b_{ki}b_{1j}\varepsilon_{it}\varepsilon_{jt}+\sum_{i\neq 1}b_{ki}b_{1i}\varepsilon^{2}_{it}$ (49)
+$b_{\cdot 1}^{\prime}\Sigma^{-1}_{\eta}b_{\cdot 1}=1.$ (48)
+Step 2. The reduced form shocks have the following representation for any $k=1,\dots,n,$
+$\eta_{1t}\eta_{kt}$ $=b_{k1}b_{11}\varepsilon^{2}_{1t}+\sum_{i\neq j}b_{ki}b_{1j}\varepsilon_{it}\varepsilon_{jt}+\sum_{i\neq 1}b_{ki}b_{1i}\varepsilon^{2}_{it}.$ (49)
 Then by Assumption 1,
 $\gamma$ $=b_{\cdot 1}b_{11}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})+\sum_{i\neq j}b_{\cdot i}b_{1j}\mathbb{E}\varepsilon_{it}\varepsilon_{jt}(Z_{t}-\mathbb{E}Z_{t})+\sum_{i\neq 1}b_{\cdot i}b_{1i}\mathbb{E}\varepsilon^{2}_{it}(Z_{t}-\mathbb{E}Z_{t})$ (51)
-$=b_{\cdot 1}b_{11}\mathbb{E}(\varepsilon^{2}_{1t}-1)(Z_{t}-\mathbb{E}Z_{t})=b_{\cdot 1}b_{11}\rho$ (52)
+$=b_{\cdot 1}b_{11}\mathbb{E}(\varepsilon^{2}_{1t}-1)(Z_{t}-\mathbb{E}Z_{t})=b_{\cdot 1}b_{11}\rho_{1}.$ (52)
 By Step 1,
-$\sqrt{\gamma^{\prime}\Sigma^{-1}_{\eta}\gamma}=|b_{11}\rho|$ (53)
+$\sqrt{\gamma^{\prime}\Sigma^{-1}_{\eta}\gamma}=|b_{11}\rho_{1}|,$ (53)
 and
-$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma\operatorname{sign}(b_{11}\rho).$ (54)
+$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma\operatorname{sign}(b_{11}\rho_{1}).$ (54)
 The statement of the Theorem follows once we assume specific sign normalization for one of the components of $b_{\cdot 1}$. ∎
@@
 Lemma B.1.
-Suppose that Assumptions 2 and 3 hold. Consider LS residuals $\hat{\eta}_{t}(H)$ from $H-$step forward autorgression,
+Suppose that Assumptions 2 and 3 hold. Consider LS residuals $\hat{\eta}_{t}(H)$ from $H-$step forward autoregression,
 $y_{t+H}$ $=V(H)\mu_{t}+\sum_{i=1}^{p}A_{i}{(H)}y_{t-i}+\eta_{t+H}(H),$ (55)
@@
 Then, we show that the LS estimators for individual processes have well-defined asymptotic distributions.
-Finally, we use invariance of LS estiates with respect to linear transformations to estimate the residual in (56).
-Step 1. Notice that $\mu_{t}=(1,t,\dots,t^{k})^{\prime}$ satisfies a first order deterministic difference equation
-$\mu_{t}=Q_{\mu}\mu_{t-1}$ (57)
+Finally, we use the invariance of LS estimates with respect to linear transformations to estimate the difference in (56).
+Step 1. Notice that $\mu_{t}=(1,t,\dots,t^{k})^{\prime}$ satisfies a first-order deterministic difference equation
+$\mu_{t}=Q_{\mu}\mu_{t-1},$ (57)
 where $Q_{\mu}$ is a lower triangular matrix
@@
 and $U_{t}=vec(\eta_{t},0,\dots,0)$.
-Matrix $Q_{y}$ can be represented in Jordan normal form  [^18] See, for example, Gantmakher (2000).
-$Q_{y}=P_{y}J_{y}P_{y}^{-1}.$ (63)
-Then, following ideas in Tsay and Tiao (1990) and Sims et al. (1990), we get VAR(1) representation
+Matrix $Q_{y}$ can be represented in Jordan normal form
+$Q_{y}=P_{y}J_{y}P_{y}^{-1},$ (63)
+where $P_{y}$ is an invertible matrix and $J_{y}$ is a block diagonal matrix.
+[^22] See, for example, Gantmakher (2000).
+Then, following ideas in Tsay and Tiao (1990) and Sims et al.
+(1990), we get VAR(1) representation
 $\begin{pmatrix}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}=\begin{bmatrix}J_{y}&0_{np\times k}\\ 0_{k\times np}&Q_{\mu}\end{bmatrix}\begin{pmatrix}P_{y}^{-1}Y^{\circ}_{t-1}\\ \mu_{t-1}\end{pmatrix}+\begin{pmatrix}\tilde{U}_{t}\\ 0\end{pmatrix},$ (64)
@@
 Suppose, without loss of generality, that the blocks $J_{y,i}$ are arranged in decreasing order according to the absolute values of the corresponding $\lambda_{i}$.
-If unit root is present, then start with $\lambda_{1}=1$.
+If a unit root is present, then start with $\lambda_{1}=1$.
 By Assumption 3, there are no roots larger than 1 in absolute value.
 Now let us find a row permutation matrix $P_{x}$ that rearranges rows of $P_{y}^{-1}Y^{\circ}_{t-1}$ to get
-$\tilde{X}_{t}\triangleq\begin{pmatrix}P_{x}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}\triangleq F(L)\nu_{t}$ (66)
+$\tilde{X}_{t}\triangleq\begin{pmatrix}P_{x}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}\triangleq F(L)\nu_{t},$ (66)
 where
 $F(L)=\begin{bmatrix}F_{11}(L)&0&0&0&\cdots&0&0\\ 0&F_{22}&0&0&\cdots&0&0\\ 0&F_{32}&F_{33}&0&\cdots&0&0\\ 0&F_{42}&F_{43}&F_{44}&\cdots&0&0\\ \vdots&\vdots&&&&\vdots&\vdots\\ 0&F_{G,2}&\cdots&\cdots&\cdots&F_{G,G}&0\\ 0&0&\cdots&\cdots&\cdots&0&I_{k}\end{bmatrix}$ (67)
-and $\nu_{t}=(\eta^{\prime}_{t},(\xi^{1}_{t}(\lambda_{1}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{1}))^{\prime},(\xi^{1}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{1}_{t}(\lambda_{b}))^{\prime},\dots,(\xi^{g_{b}}_{t}(\lambda_{1}))^{\prime},1,t,\dots,t^{k})^{\prime}$ with $\xi^{j}_{t}(\lambda)\triangleq\lambda\sum_{s=1}^{t}\xi^{j-1}_{t}(\lambda)$, $\xi^{0}_{t}(\lambda)\triangleq\eta_{t}\lambda^{t}$, $i=1,\dots,b$ $\lambda_{i}$ are unit roots of (12) and $g_{i}$ are their corresponding maximum geometric multiplicities.
-The blocks $F_{i,j}$ corresponding to two different unit roots $\lambda_{i}$ and $\lambda_{j}$ are 0.
+and $\nu_{t}=(\eta^{\prime}_{t},(\xi^{1}_{t}(\lambda_{1}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{1}))^{\prime},(\xi^{1}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{1}_{t}(\lambda_{b}))^{\prime},\dots,(\xi^{g_{b}}_{t}(\lambda_{1}))^{\prime},1,t,\dots,t^{k})^{\prime}$ with $\xi^{j}_{t}(\lambda)\triangleq\lambda\sum_{s=1}^{t}\xi^{j-1}_{s}(\lambda)$, $\xi^{0}_{t}(\lambda)\triangleq\eta_{t}\lambda^{t}$, $i=1,\dots,b$ $\lambda_{i}$ are unit roots of (12) and $g_{i}$ are their corresponding maximum geometric multiplicities.
+The blocks $F_{i,j}$ corresponding to two different unit-roots $\lambda_{i}$ and $\lambda_{j}$ are 0.
 The lag matrix $F_{11}(L)$ is obtained by inversion of the stationary rows of (64) and therefore $\sum^{\infty}_{j=0}j^{g}|F_{11j}|<\infty$ and $\sum_{j=0}^{\infty}F_{11j}F_{11j}^{\prime}$ is not singular.
 Following Sims et al.
-(1990); Tsay and Tiao (1990), one can arrange the properly chosen powers of $T^{1/2}$ in a diagonal matrix $\Upsilon_{T}$ such that $\sum_{t=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{t-1}$ have a bounded variance (for example, for stationary components it is $T^{1/2}$, for components containing $\xi^{j}_{t}(\lambda_{2}))$ it is $T^{(1+j)/2}$ etc.).
+(1990) and Tsay and Tiao (1990), one can arrange the properly chosen powers of $T^{1/2}$ in a diagonal matrix $\Upsilon_{T}$ such that $\sum_{t=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{t-1}$ have a bounded variance (for example, for stationary components it is $T^{1/2}$, for components containing $\xi^{j}_{t}(\lambda_{2}))$ it is $T^{(1+j)/2}$ etc.).
 Step 2. We would like to show now that for all $h=0,1,2,\dots,H$, there exist a diagonal matrix $\Upsilon_{T}$ such that
@@
 Consider $h=0$ first.
-The proof for deterministic, stationary and ordinary unit root components follows the proofs of Lemmas 1 and 2 in Sims et al.
+The proof for deterministic, stationary, and ordinary unit root components follows the proofs of Lemmas 1 and 2 in Sims et al.
 (1990) with a modification required to account for conditionally heteroscedastic shocks $\eta_{t}$.
@@
 (1990) requires replacement of Functional CLT for sums of $\xi^{1}_{t}$ in part (a) and Theorem 2.4(ii) from Chan and Wei (1988) in part (c) for case $p=1$.
-Under Assumption 2, Theorem 3.1 (De Jong and Davidson, 2000) applied to $\xi^{1}_{t}$ gives the required FCLT under mixing processes assumption.
-(De Jong and Davidson, 2000, Theorem 4.1) for any pair of components of $\xi^{1}_{t}$ replaces Theorem 2.4(ii) of Chan and Wei (1988) for mixing case (white noise assumption on $\eta_{t}$ ensures that we do not need recenter the sums).
-For complex unit roots, the limiting distribution is given in Theorem 4.4 of Tsay and Tiao (1990), which also requires replacing FCLT for martingale differences with mixing white noise (De Jong and Davidson, 2000, Theorem 4.1) and stong law of large numbers used in from Chan and Wei (1988) for MDS with Corollary 3.2 in Rio et al.
+Under Assumption 2, Theorem 3.1 (De Jong and Davidson, 2000) applied to $\xi^{1}_{t}$ gives the required FCLT under the mixing processes assumption.
+Theorem 4.1 of De Jong and Davidson (2000) for any pair of components of $\xi^{1}_{t}$ replaces Theorem 2.4(ii) of Chan and Wei (1988) for the mixing sequence case (the MDS assumption on $\eta_{t}$ ensures that we do not need to recenter the sums).
+For complex unit roots, the limiting distribution is given in Theorem 4.4 of Tsay and Tiao (1990), which also requires substituting FCLT for martingale differences with mixing white noise (De Jong and Davidson, 2000, Theorem 4.1) and strong law of large numbers used in from Chan and Wei (1988) for MDS with Corollary 3.2 in Rio et al.
 (2017).
-For our purposes, we do not need explicit form of $\phi$ and $\tilde{V}$.
+For our purposes, we do not need the explicit form of $\phi$ and $\tilde{V}$.
 For $h>0$, the existence of the limiting distributions follows the proof of Theorem 4.5 of Tsay and Tiao (1990).
@@
 $y_{t+h}=\Theta(h)X_{t-1}+\eta_{t+h}(h),$ (70)
-where $\Theta(h)\triangleq(A_{1}(h),A_{2}(h),\dots,A_{p}(h),V(h))$ and $X_{t-1}\triangleq vec(y_{t-1},y_{t-2},\dots,y_{t-p},1,(t-1),\dots(t-1)^{k})$.
+where
+$\Theta(h)\triangleq(A_{1}(h),A_{2}(h),\dots,A_{p}(h),V(h))$
+and
+$X_{t-1}\triangleq vec(y_{t-1},y_{t-2},\dots,y_{t-p},y_{t-p-1},1,(t-1),\dots(t-1)^{k}).$
 Each period $t=p,\dots,(T-h)$ we have
-$\hat{\eta}_{t+h}(h)-\eta_{t+h}(h)=(\Theta(h)-\hat{\Theta}(h))X_{t-1}=-X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h).$ (71)
+$\hat{\eta}_{t+h}(h)-\eta_{t+h}(h)=(\Theta(h)-\hat{\Theta}(h))X_{t-1}=-X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h).$ (71)
 By Step 1, there exist a linear transformation $P$ such that $X_{t}=P^{-1}\tilde{X}_{t}$. So we can get a representation
-$X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h)$ $=\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime}(P^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime})^{-1}\sum_{r=p}^{T-h}(P^{-1})\tilde{X}_{r-1}\eta_{r+h}(h),$ (72)
-$=\tilde{X}_{t-1}^{\prime}(\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\eta_{r+h}(h).$ (73)
+$X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h)$ (72)
+$=$ $\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime}(P^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}(P^{-1})^{\prime})^{-1}\sum_{r=p}^{T-h}(P^{-1})\tilde{X}_{r-1}\eta_{r+h}(h)$ (73)
+$=$ $\tilde{X}_{t-1}^{\prime}(\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\eta_{r+h}(h).$ (74)
 Taking weighted sums over $t$ using matrix $\Upsilon_{T}$ from Step 2, we get representations
@@
 $\hat{C}_{h}^{\prime}$ $=(\sum_{t=p}^{T-h}\hat{\eta}_{t}\hat{\eta}_{t}^{\prime})^{-1}\sum_{t=p}^{T-h}\hat{\eta}_{t}(\hat{\eta}_{t+h}(h))^{\prime}=\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\sum_{s=0}^{h}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))$ (82)
-$=C_{h}^{\prime}+\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\frac{1}{T}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))+O_{p}(T^{-1})$ (82)
-Since $\eta_{t}$ is stationary and mixing and thus sequence $\eta_{t}\eta_{t+h}$ is ergodic with finite fourth moments, so we can apply the ergodic theorem to get the desired result.
+$=C_{h}^{\prime}+\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\frac{1}{T}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))+O_{p}(T^{-1}).$ (83)
+Since $\eta_{t}$ is stationary and mixing and thus sequence $\eta_{t}\eta_{t+h}$ is ergodic with finite fourth moments, we can apply the ergodic theorem to get the desired result.
 The same proof of consistency applies to $\hat{\Sigma}_{\eta}$. ∎
@@
 $\hat{\eta}_{it}\hat{\eta}_{1t}(Z_{t}-\bar{Z}_{T})=$ $\eta_{it}\eta_{1t}(Z_{t}-\bar{Z}_{T})+(\hat{\eta}_{it}-\eta_{it})\eta_{1t}(Z_{t}-\bar{Z}_{T})$ (84)
-$+\eta_{it}(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})+(\hat{\eta}_{it}-\eta_{it})(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})$ (84)
+$+\eta_{it}(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})+(\hat{\eta}_{it}-\eta_{it})(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T}).$ (85)
 First note that by Step 3 in proof of Lemma B.1,
-$\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})\eta_{jt}=O_{p}(1)$ (85)
-Next, by Assumption 2.5, $Z_{t}\eta_{t}$ is zero meand and is uncorrelated with $\tilde{X}_{t-1}^{\prime}$.
+$\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})\eta_{jt}=O_{p}(1).$ (86)
+Next, by Assumption 2.5, $Z_{t}\eta_{t}$ is zero mean and is uncorrelated with $\tilde{X}_{t-1}^{\prime}$.
 Using the same argument as in Step 2 of the proof of Lemma B.1, $\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}$ is a tight random sequence by De Jong and Davidson (2000, Theorem 4.1).
 So
-$(\sum_{t=p}^{T}(\hat{\eta}_{t}-\eta_{t})\eta^{\prime}_{t}Z_{t})^{\prime}=\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r}=O_{p}(1)$ (86)
-The last term in (84) can be bounded by Cauchy-Schwarz theorem,
+$(\sum_{t=p}^{T}(\hat{\eta}_{t}-\eta_{t})\eta^{\prime}_{t}Z_{t})^{\prime}=\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r}=O_{p}(1).$ (87)
+The last term in (85) can be bounded by the Cauchy-Schwarz theorem,
 $|\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})(\hat{\eta}_{1t}-\eta_{1t})(Z_{t}-\bar{Z}_{T})|\leq\sqrt{\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})^{2}\sum_{t=p}^{T}(\hat{\eta}_{1t}-\eta_{1t})^{2}(Z_{t}-\bar{Z}_{T})^{2}}=O_{p}(T^{1/4}),$ (88)
@@
 By Assumption 2, the time series $\eta_{t}\eta_{1t}$ and $Z_{t}$ satisfy CLT for mixing processes. Therefore,
-$\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\bar{Z}_{T})$ $=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(Z_{t}-\mathbb{E}Z_{t})\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})$ (92)
-$=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(1)$ (93)
+$\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\bar{Z}_{T})=$ $\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})$ (93)
+$+\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(Z_{t}-\mathbb{E}Z_{t})\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})$ (94)
+$=$ $\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(1).$ (95)
 To summarize,
@@
 Proof of Theorem 4.
-By Assumption 2, $\Xi^{H}_{t}$ is a strong mixing process of size at least $-2(1+\epsilon)/\epsilon$ with zero mean and finite $2+\epsilon$ moments (condition DMR is satisfied, see p.
+By Assumption 2, $\Xi_{t}(H)$ is a strong mixing process of size at least $-2(1+\epsilon)/\epsilon$ with zero mean and finite $2+\epsilon$ moments (condition DMR is satisfied, see p.
 12 in Rio et al.
@@
 By Theorem 4.2 in Rio et al.
-(2017) and the Cramer-Wold theorem, there exist the long run variance matrix $\Omega$ for $\Xi^{H}_{t}$ and
+(2017) and the Cramer-Wold theorem, there exists the long run variance matrix $\Omega$ for $\Xi_{t}(H)$ and
 $\frac{1}{\sqrt{T-p-H}}\sum_{t=p}^{T-H}\Xi_{t}(H)\to N(0,\Omega).$ (97)
@@
 Proof of Theorem 5.
-Step 1.To show consitency of the whole matrix $\hat{\Omega}=\hat{\Omega}(H_{1})$, we need to establish consistency of HAC estimators for all combinations $a^{\prime}\Xi_{t}$.
+Step 1.To show consistency of the whole matrix $\hat{\Omega}=\hat{\Omega}(H_{1})$, we need to establish consistency of HAC estimators for all combinations $a^{\prime}\Xi_{t}$.
 Then the HAC estimator for particular $a$ is defined as
-$\hat{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\hat{\Xi}_{t})(\hat{\Xi}_{t+\ell}^{\prime}a)$ (99)
+$\hat{\Omega}(a)\triangleq\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\hat{\Xi}_{t})(\hat{\Xi}_{t+\ell}^{\prime}a).$ (101)
 This estimator has an infeasible analog
-$\tilde{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\Xi_{t})(\Xi_{t+\ell}^{\prime}a)$ (100)
+$\tilde{\Omega}(a)\triangleq\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\Xi_{t})(\Xi_{t+\ell}^{\prime}a).$ (102)
 Since $\Xi_{t}=\Xi_{t}(H_{1})$ is strong mixing with $4+\epsilon$ moment, Theorem 1(a) in Andrews (1991) implies that $\tilde{\Omega}(a)$ is consistent for $\Omega(a)=a^{\prime}\Omega a$.
@@
 What remains to show is $\hat{\Omega}(a)-\tilde{\Omega}(a)\stackrel{{\scriptstyle p}}{{\rightarrow}}0$.
-Since by Theorem 2, $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$, we just need to establish HAC consitency for vectors
-$\Psi_{t}=\begin{pmatrix}\hat{\psi}^{0}_{t}\\ \hat{\psi}^{1}_{t}\\ \cdots\\ \hat{\psi}^{H_{1}}_{t}\\ \hat{\psi}^{H_{1}+1}_{t}\end{pmatrix}=\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}$ (101)
+Since by Theorem 2, $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$, we just need to establish HAC consistency for vectors
+$\Psi_{t}=\begin{pmatrix}\hat{\psi}^{0}_{t}\\ \hat{\psi}^{1}_{t}\\ \cdots\\ \hat{\psi}^{H_{1}}_{t}\\ \hat{\psi}^{H_{1}+1}_{t}\end{pmatrix}=\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}.$ (103)
 The last component, $\hat{\xi}^{H_{1}+1}_{t}$, has a different structure from the other components since it is multiplied by $(Z_{t}-\bar{Z}_{t})$.
 From the proof of Theorem 3, we know that $\max_{t=1,\dots T}|Z_{t}-\bar{Z}_{t}|=O_{p}(T^{1/4})$.
-So, the conistency of the HAC components, including the linear combinations of $\hat{\psi}^{H_{1}+1}_{t}$ can be proven in a similar fashion as the other components.
+So, the consistency of the HAC components, including the linear combinations of $\hat{\psi}^{H_{1}+1}_{t}$ can be proven similarly to the other components.
 So, to save space, we only consider vector $a=vec(a_{0},a_{1},\dots,a_{H_{1}},0_{n})$ that puts 0 on the last $n$ components and all $a_{i}\in R^{n^{2}}$.
 $\hat{\Psi}_{t}^{\prime}aa^{\prime}\hat{\Psi}_{t+\ell}-\Psi_{t}^{\prime}aa^{\prime}\Psi_{t+\ell}=$ $\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}\hat{\psi}^{s}_{t+\ell}$ (104)
-$+\sum_{h,s=0}^{H}(\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})+\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})$ (103)
+$+\sum_{h,s=0}^{H}(\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})$ (105)
+$+\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell}).$ (106)
 The first term after summation over $t$ becomes
@@
 $\leq\sqrt{b^{\prime}(\sum_{t=p}^{T-\ell}\Upsilon_{T}^{-1}\tilde{X}_{t-1}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1})b}\sqrt{\sum_{t=p}^{T-\ell}(a_{js}^{\prime}(\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)-\mathbb{E}\eta_{t+\ell}\eta^{\prime}_{t+\ell+s}(s)e_{j}e^{\prime}_{i}\eta_{t})^{2}}.$ (116)
-The r.h.s. is $O_{p}(\sqrt{T})$ by proof of Step 2 of Lemma B.1 and Markov inequality implied by existence of $6+\epsilon$ moments of the stationary mixing sequence $\eta_{t}$.
-The sums of remaining two terms in (103) are also $O_{p}(\sqrt{T})$, which can be shown using the same techniques.
+The r.h.s. is $O_{p}(\sqrt{T})$ by proof of Step 2 of Lemma B.1 and Markov inequality implied by the existence of $6+\epsilon$ moments of the stationary mixing sequence $\eta_{t}$.
+The sums of the remaining two terms in (106) are also $O_{p}(\sqrt{T})$, which can be shown using the same techniques.
 Step 3. Putting all bounds together gives us
@@
 $\hat{Q}_{T}$ $=\frac{1}{(T-p-H)}\sum_{t,l=p}^{T-H}\hat{\Xi}_{t}(H_{1})\hat{\Xi}^{\prime}_{l}(H_{1})\mathbb{E}u^{s}_{t}u^{s}_{l}=\frac{1}{(T-p-H)^{2}}\sum_{t,l=p}^{T-H}\hat{\Xi}_{t}(H_{1})\hat{\Xi}^{\prime}_{|t-l|}(H_{1})K(|t-l|/B_{T})$ (118)
-$=\sum_{\ell=-B_{T}}^{B_{T}}\frac{K(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})-\sum_{\ell=-B_{T}}^{B_{T}}\frac{\ell K(\ell/B_{T})}{T(T-\ell)}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})=\hat{J}_{T}+O_{P}(\frac{B_{T}}{T})$ (116)
+$=\sum_{\ell=-B_{T}}^{B_{T}}\frac{K(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})-\sum_{\ell=-B_{T}}^{B_{T}}\frac{\ell K(\ell/B_{T})}{T(T-\ell)}\sum_{t=p}^{T-\ell}\hat{\Xi}_{t}(H_{1})\hat{\Xi}_{t+\ell}^{\prime}(H_{1})=\hat{J}_{T}+O_{P}(\frac{B_{T}}{T}).$ (119)
 Since $\frac{B_{T}}{T}\to 0$ , by Theorem 5 we get $\hat{Q}_{T}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Omega$. So by Theorem 4,